### v2

# SEMF: Supervised Expectation-Maximization Framework for Predicting Intervals

Abstract
This work introduces the Supervised Expectation-Maximization Framework (SEMF), a versatile and model-agnostic framework that generates prediction intervals for datasets with complete or missing data. SEMF extends the Expectation-Maximization (EM) algorithm, traditionally used in unsupervised learning, to a supervised context, enabling it to extract latent representations for uncertainty estimation. The framework demonstrates robustness through extensive empirical evaluation across 11 tabular datasets, achieving—in some cases—narrower normalized prediction intervals and higher coverage than traditional quantile regression methods [^1] Code: https://github.com/Unco3892/semf-paper. Furthermore, SEMF integrates seamlessly with existing machine learning algorithms, such as gradient-boosted trees and neural networks, exemplifying its usefulness for real-world applications. The experimental results highlight SEMF’s potential to advance state-of-the-art techniques in uncertainty quantification.
1 Introduction
In the evolving field of machine learning (ML), the quest for models capable of predicting outcomes while quantifying the uncertainty of their predictions is critical. Estimating prediction uncertainty is particularly vital in high-stakes domains such as healthcare [1], finance [2], and autonomous systems [3], where prediction-based decisions can have significant consequences. Traditional approaches have primarily focused on point estimates, providing limited insight into the reliability of these predictions. This limitation underscores the necessity for frameworks that generate precise point predictions and robust prediction intervals. These intervals offer a range within which the true outcome is expected to lie with a specified probability, enabling a deeper understanding of prediction uncertainty. Consequently, research has increasingly focused on methodologies that extend beyond point estimation to include uncertainty quantification, thereby enhancing decision-making in applications that rely on predictive modeling [4].
In this paper, we introduce the Supervised Expectation-Maximization Framework (SEMF) based on the Expectation-Maximization (EM) algorithm [5]. Traditionally recognized as a clustering technique, EM is used for supervised learning in SEMF, allowing for both point estimates and prediction intervals using any ML model (model-agnostic). SEMF generates representations for latent or missing modalities, making it particularly relevant for incomplete multi-modal data. This paper details the methodology behind the framework and proposes a training algorithm based on Monte Carlo (MC) sampling, also used in variational inference for Variational Auto-Encoders (VAEs) [6, 7]. Comparisons with other models across several datasets showcase its contribution to the existing uncertainty quantification and predictive modeling literature [8].
The remainder of this paper is organized as follows: Section 2 details the theory and the methodology of SEMF. Section 3 reviews related works in latent representation learning, uncertainty estimation, and handling of missing data. Section 4describes the experimental setup, including datasets and evaluation metrics. Section 5 discusses the results, demonstrating the efficacy of SEMF. Lastly, Section 6 concludes the paper, and Section 7 outlines the limitations and potential research directions.
2 Method
This section presents the founding principles of SEMF from its parameters, training, and inference procedure with, at its core, the EM algorithm. This algorithm, first introduced by [5], is an unsupervised method for handling latent variables and incomplete data. Invented to maximize the model likelihood, it builds a sequence of parameters that guarantee an increase in the log-likelihood [9] by iterating between the Expectation (E) and the Maximization (M) steps. In the E-step, one computes
[…]
where $p^{\prime}$ stands for the current estimates, $\log p(x,z)$ is the log-likelihood of the complete observation $(x,z)$, and $z$ is a latent variable. The M-step maximizes this $Q$-function: $p^{\prime}\leftarrow\arg\max_{p}Q(p|p^{\prime})$ The sequence is repeated until convergence.
The EM algorithm has already been used for supervised learning tasks using specific models [10, 11, 12, 13], and, to the best of our knowledge, mainly focusing on point prediction through the use of Gaussian Mixture Models (GMMs). Our work differs by modifying the EM algorithm and MC sampling to generate prediction intervals with any ML model and, in theory, under any distribution. This results in SEMF, a framework consisting of an encoder-decoder architecture and a missing data simulator.
2.1 Problem Scenario
Let $x=(x_{1},x_{2},\ldots,x_{K})$ denote $K$ inputs and the output be $y$. For simplicity, we limit $y$ to be numerical, although it could be categorical without loss of generality. Component $x_{k}$ is a source: a modality, a single or group of variables, or an unstructured input such as an image or text. For clarity, we limit to $K=2$, where $x_{1}$ and $x_{2}$ are single variables. We assume that only $x_{1}$ may contain missing values, either at random or partially at random.
[…]
where $p^{\prime}$ is our current estimate. Eq. 2 can be estimated by MC sampling. Since sampling from $p^{\prime}(z|y,x)$ can be inefficient, we rather rely on the decomposition $p^{\prime}(z|y,x)=p^{\prime}(y|z)p^{\prime}(z|x)/p^{\prime}(y|x)$. Thus, we sample $z_{r}$ from $p^{\prime}(z|x)$, $r=1,\ldots,R$, and, setting $w_{r}=p^{\prime}(y|z_{r})/\sum_{t}p^{\prime}(y|z_{t})$, approximate the right-hand side term of Eq. 2
[…]
If $x_{1}$ is missing, a similar development leads to
[…]
where $x_{1,r}$ and $z_{r}$ are respectively sampled from $p^{\prime}(x_{1}|x_{2})$ and $p^{\prime}(z|x_{1,r},x_{2})$.
2.2 Objective Function
Adapting Eq. 3 and Eq. 4 for the observed data $\{(y_{i},x_{i})\}_{i=1}^{N}$, the overall loss function, $\mathcal{L}$, is
$\mathcal{L}(\phi,\theta,\xi)=-\sum_{i=1}^{N}\sum_{r=1}^{R}\left\{\log p_{\phi}(z_{i,r}|x_{i,r})+\log p_{\theta}(y_{i}|z_{i,r})+\mathds{1}_{\{i\in I_{m}\}}\log p_{\xi}(x_{1,i,r}|x_{2,i})\right\}w_{i,r},$ (5)
where $I_{m}$ is the set of those $i$’s such that $x_{1,i}$ is missing. The models of $p(y|z)$, $p(z|x)$, and $p(x_{1}|x_{2})$ inherit parameters $\theta$, $\phi$, and $\xi$, respectively. Also, $x_{1,i,r}$ is sampled from $p_{\xi^{\prime}}(x_{1}|x_{2,i})$, if $x_{1,i}$ is missing, and $z_{1,i,r}$ and $z_{2,i,r}$ are sampled from $p_{\phi^{\prime}}(z_{1}|x_{1,i,r})$ and $p_{\phi^{\prime}}(z_{2}|x_{2,i})$. Furthermore, for compactness of notation, $x_{i,r}$ is $(x_{1,i},x_{2,i})$ if $x_{1,i}$ is observed, and $(x_{1,i,r},x_{2,i})$ if $x_{1,i,r}$ is missing. Finally, the weights are
$w_{i,r}=\frac{p_{\theta^{\prime}}(y_{i}|z_{i,r})}{\sum_{t=1}^{R}p_{\theta^{\prime}}(y_{i}|z_{i,t})}.$ (6)
Eq. 5 shows that ${\cal L}$ is a sum of losses associated with the encoder model, $p_{\phi}$, for each source, the decoder model, $p_{\theta}$, from the latent variables to the output, and, if applicable, the model handling missing data, $p_{\xi}$. At each M-step, ${\cal L}$ is minimized with respect to $\theta$, $\phi$, and $\xi$. Then, $\theta^{\prime}$, $\phi^{\prime}$, and $\xi^{\prime}$ are updated, as well as the weights and the sampling. Then the process is iterated until convergence.
2.2.1 Example: $\mathcal{L}(\phi,\theta,\xi)$ under normality
Similar to [7], we develop further ${\cal L}$ under normality assumptions for the encoder, $p_{\phi}$, and the decoder, $p_{\theta}$. These simple cases are illustrative, though any other distributions could be adopted, including non-continuous or non-numerical outcomes.
Encoder $p_{\phi}(z|x)$.
Let $m_{k}$ be the length of the latent variable $z_{k}$, $k=1,2$. We assume a normal model for $Z_{k}$ given $X_{k}=x_{k}$,
$Z_{k}|X_{k}=x_{k}\sim{\cal N}_{m_{k}}(g_{\phi_{k}}(x_{k}),\sigma_{k}^{2}J_{m_{k}}),$ (7)
where $J_{m_{k}}$ is the $m_{k}\times m_{k}$ identity matrix. In particular,
$\log p_{\phi}(z_{k}|x_{k})$ $=$ $-\frac{m_{k}}{2}\log 2\pi-\frac{1}{2}\log\sigma_{k}^{2}-\frac{1}{2\sigma_{k}^{2}}\sum_{j=1}^{m_{k}}\{z_{k,j}-g_{\phi_{k},j}(x_{k})\}^{2},\quad k=1,2.$ (8)
The mean $g_{\phi_{k}}(x_{k})$ can be any model, such as a neural network, with output of length $m_{k}$, $k=1,2$. The scale $\sigma_{k}$ can be fixed, computed via the weighted residuals, or learned through a separate set of models. It controls the amount of noise introduced in the latent dimension and is pivotal in determining the prediction interval width for $p(y|z)$. In this paper, $\sigma_{k}$ is fixed for simplicity.
Decoder $p_{\theta}(y|z)$.
We assume a normal model for $Y$ given $Z=z$,
$Y|Z=z\sim\mathcal{N}(f_{\theta}(z),\sigma^{2}).$ (9)
This results in a log-likelihood contribution,
$\log p_{\theta}(y|z)=-\frac{1}{2}\log 2\pi-\frac{1}{2}\log\sigma^{2}-\frac{1}{2\sigma^{2}}\{y-f_{\theta}(z)\}^{2}.$ (10)
Again, the mean $f_{\theta}(z)$ can be any model, such as a neural network.
Model for missing data $p_{\xi}(x_{1}|x_{2})$.
We use an empirical model for $X_{1}$ given $X_{2}=x_{2}$ where $p_{\xi}(x_{1}|x_{2})$ put masses only on those non-missing $x_{1}$’s in the training set. Let $I_{nm}$ be the set of indices such that $x_{1,j}$, $j\in I_{nm}$, are all the non-missing $x_{1}$ in the training set. Additionally, for a given $j\in I_{nm}$, $x_{1}[j]$ is the observed $x_{1}$ corresponding to $j$. For a given $x_{2}$, $p_{\xi}(x_{2})$ is a vector of length $|I_{nm}|$, the cardinality of $I_{nm}$, with components
$p_{\xi}(j|x_{2})=\frac{\exp\{h_{\xi,j}(x_{2})\}}{\sum_{t\in I_{nm}}\exp\{h_{\xi,t}(x_{2})\}},\quad j\in I_{nm},$ (11)
where $h_{\xi}$ is a vector of length $|I_{nm}|$, typically a neural network with input $x_{2}$ and output on $I_{nm}$. Now, the probability of $X_{1}=x_{1}$ given $X_{2}=x_{2}$ is
$p_{\xi}(x_{1}|x_{2})=\sum_{j\in I_{nm}}p_{\xi}(j|x_{2})\cdot\mathds{1}\{x_{1}=x_{1}^{(nm)}[j]\}.$ (12)
Summary.
Overall, the M-step is
$\phi_{k}^{*}=\arg\min_{\phi_{k}}\sum_{i,r}w_{i,r}\sum_{j=1}^{m_{k}}\{z_{k,i,r,j}-g_{\phi_{k},j}(x_{k,i,r})\}^{2},\quad k=1,2,$ (13)
$\theta^{*}=\arg\min_{\theta}\sum_{i,r}w_{i,r}\{y_{i}-f_{\theta}(z_{i,r})\}^{2},$ (14)
$(\sigma^{*})^{2}=\frac{1}{N}\sum_{i,r}w_{i,r}\{y_{i}-f_{\theta^{*}}(z_{i,r})\}^{2},$ (15)
$\xi^{*}=\arg\max_{\xi}\sum_{t\in D_{x}}w_{t}\log p_{\xi}(j_{t}|x_{2,t}).$ (16)
When $x_{1,i}$ is missing, the sampling is enriched by simulated $x_{1,i,r}$. Eq. 16 selects $j^{*}$ from $I_{nm}$ based on $p_{\xi^{\prime}}(j|x_{2})$ according to Eq. 12 and Eq. 11. The above $D_{x}$ is a subset of the data for $\xi$ is only learned on the missing data part. After discussing the training steps in Subsection 2.3, the details are given in Algorithm 1 (Appendix A). We also note that the parameters above can be learned in parallel.
2.3 Training
For efficiency purpose, the training set, $\{1,\ldots,N\}$, is segmented into batches $\{b_{1},\ldots,b_{L}\}$ on which the index $i$ runs (and thus the denominator of Eq. 15 must be adapted accordingly). The process iterates for each batch until the maximum number of steps is reached or an early stopping criterion is satisfied. The framework requires tuning hyper-parameters such as the number of MC samples $R$, the number of latent nodes $m_{k}$, and the standard deviation $\sigma_{k}$ of $Z_{k}$. Monitoring the point prediction on a hold-out validation is important to combat overfitting and terminate the training early with a patience hyper-parameter. Moreover, due to the generative nature of SEMF, the variation resulting from the initial random seed is measured in Subsection 4.2. Additionally. the model-specific hyper-parameters ($p_{\phi}$, $p_{\theta}$ and $p_{\xi}$) are also discussed in the same Subsection.
2.4 Inference
The encoder-decoder structure of SEMF entails the simulations of $z_{r}$ during inference, as depicted in Figure 1. In theory, any inference can be performed for $\hat{y}$, for instance the mean value $\hat{y}=\frac{1}{R}\sum_{r=1}^{R}f_{\theta}(z_{r})$, where $z_{r}\sim p_{\phi}(z|x)$ (see Algorithm 2 in Appendix A). For prediction intervals, a double simulation scheme is used,
$z_{r}\sim p(z|x),\quad\hat{y}_{r,s}\sim p_{\theta}(y|z_{r}),\quad r,s=1,\ldots,R.$ (17)
Prediction interval at a given level of certainty $\alpha$ follows as
$PI=\text{quantile}\left(\{\hat{y}_{r,s}\};\frac{\alpha}{2},1-\frac{\alpha}{2}\right).$ (18)
Remark.
We denote the $R$ for inference as $R_{\mathrm{infer}}$.
Figure 1: Inference procedure with the SEMF’s learnable parameters $\phi_{k},\theta$ and $\xi$. Here, we illustrate the number of inputs $k$ as $k=1,2$, assuming that $x_{1}$ may contain missing values
3 Related Work
3.1 Latent Representation Learning
Latent representation learning involves modeling hidden variables from observed data for various ML tasks, most notably Auto-Encoders (AEs) and VAEs. AEs are neural networks reconstructing inputs by learning an intermediate latent representation. The leftmost part of AE is the encoder $g_{\phi}(\cdot)$, which embeds the input $x$ into a latent variable $z$. It is passed to the rightmost part of AE, the decoder $f_{\theta}(\cdot)$, for the reconstruction $\hat{x}=f_{\theta}(z)$ of $x$ through $z$. The parameters $\phi$ and $\theta$ are trained by minimizing the reconstruction loss. Unlike AEs, which focus on reconstruction, VAEs use variational methods to fit distributions of latent variables and of the output [7]. Given a sample $x$ with a latent variable of $z$, VAEs learn the marginal likelihood, $p_{\theta}(x)$. Direct maximization of this likelihood is difficult [14]. Thus, alternative methods exist, such as maximizing the evidence lower bound (ELBO), which provides guarantees on the log-likelihood [15].
For supervised and semi-supervised tasks, latent representation learning can include task-specific predictions [16]. More specifically, models such as AEs follow the classical encoder-decoder objective while training a predictor $h_{\psi}(z)$ through an additional layer or model to estimate the output $y$. This dual objective helps in learning more task-relevant embeddings [17, 18]. Semi-supervised VAEs are similar, with the distinction that they couple the reconstruction loss of the unlabeled data with a variational approximation of latent variables. This is effective even with sparse labels [19, 20]. Similarly, the EM algorithm adapts well to minimal supervision [21] and using labeled and unlabeled data in semi-supervised settings for both single and multiple modalities [12, 22].
3.2 Prediction Intervals
Crucial for estimating uncertainty, prediction intervals in regression are often derived using methods such as Bayesian approaches [23, 24, 25], ensemble techniques [26, 27], or quantile regression [28, 29]. Additionally, these methods can be complemented with conformal prediction, a framework for calibrating any point predictor to produce prediction intervals [30, 31], making it highly relevant for enhancing reliability in applications requiring rigorous uncertainty quantification.
A key component of quantile regression is the pinball loss function, which effectively balances the residuals to capture the desired quantiles even for non-parametric models [32], making it ideal for asymmetric distributions where tail behavior is of critical importance [29]. This loss function is pivotal not only for single model scenarios but also enhances ensemble methods by refining their quantile estimate [33]. Conformal prediction further extends the applicability of these intervals by providing a layer of calibration that adjusts intervals obtained from any predictive model, ensuring they cover the true value with a pre-specified probability [34].
3.3 Missing Data
Managing missing values is a pivotal aspect when dealing with real-world data. Naive methods such as discarding instances or mean/median-imputation may be infeasible or carry the risk of changing the data distribution [35, 36] introducing bias if the missing values are not missing at random (MAR) or completely at random (MCAR). The chosen technique for handling missing data should adhere to the dataset’s characteristics and mechanisms behind the missing data [37].
More advanced approaches, like the Iterative Imputer from scikit-learn [38], an implementation of Multiple Imputation by Chained Equations (MICE) [39], expand on the simple imputations by iteratively modeling each feature with missing values as a function of other features [40, 41]. Due to its complexity, MICE is best compared with alternatives such as K-means clustering and artificial neural networks (ANNs). K-means assigns missing values based on cluster centroids [42] as opposed to ANNs, which learn complex, non-linear relationships between variables [43]. ANNs effectively predict or reconstruct missing values and thus are particularly useful in datasets where relationships between variables are intricate and not easily captured by straightforward imputation methods. Accordingly, ANNs have demonstrated superior performance over MICE and GMM (with EM for missing values) in scenarios where a large proportion of the data is missing [44].
4 Experimental Setup
4.1 Datasets
We systematically curate a subset of datasets from the OpenML-CTR23 [45] benchmark suite to evaluate and carry out our experiments. Initially comprising 35 datasets, we apply an exclusion criteria to refine this collection to 11 datasets. The details and overview are in Appendix B. We remove duplicated rows from all the datasets and carry out the standardization (scaling) of all predictors, including the outcome, which we transformed to have zero means and unit variances. The features of these datasets are then treated as separate inputs to SEMF. In the second stage of our experiments, we artificially introduce 50% missing values in our datasets for any predictor except for the first feature of a randomly chosen row, which emulates MAR or MCAR data. In all our datasets, 70% of the data is used to train all models, 15% as a hold-out validation set to monitor SEMF’s performance, and 15% to evaluate the models. To combat overfitting, baseline models that benefit from early stopping are allocated another 15% from the training data. Lastly, it is essential to note that all data in SEMF are processed batch-wise, without employing mini-batch training, to ensure consistency and stability in the training process.
4.2 Models
Our baseline consists of both point and quantile regression eXtreme Gradient Boosting (XGBoost) [46], Extremely Randomized Trees (ET) [47], and neural networks [48], depicted in Table 1. To ensure consistency in our experimental setup, we align the families and hyper-parameters of $p_{\phi}$ and $p_{\theta}$ with our baseline models. For example, in the case of XGBoost in SEMF, we use $K$ XGBoost’s $g_{\phi_{k}}(x_{k})$, one for each input $x_{k}$, $k=1,\ldots,K$, and one XGBoost for $f_{\theta}(z)$ with the same hyper-parameters. We refer to the SEMF’s adoption of these models as MultiXGBs, MultiETs, and MultiMLPs. When establishing prediction intervals, we set our uncertainty tolerance at 5% for both the baseline and SEMF (Eq. 18). The missing data simulator, $p_{\xi}$, is constructed using a shallow neural network, which employs the $\mathrm{SELU}$ activation function [49]. It is experimented with two distinct node counts: 50 and 100.
To constrain the breadth of our parameter exploration, the simulator for the missing model adopts the optimal set of hyper-parameters identified from analyses involving complete datasets. We target (larger) $\sigma_{k}$ values that introduce more noise and produce better intervals than point predictions. The optimal models are then trained and tested with five different seeds, and the results are averaged. The point prediction, $\hat{y}$, uses the mean inferred values. We then study the performance of SEMF against mean and median imputation techniques, five nearest neighbors, and MICE from [38]. The imputers are used within the point and interval baseline models explained in the following subsection to form the missing value baseline. Appendix C contains more details on the hyper-parameters for each SEMF model and dataset.
Table 1: SEMF models, baselines, and hyper-parameters.
SEMF Point Prediction Baseline Interval Prediction Baseline
MultiXGBs XGBoost [50]
Trees: 100, Maximum depth: 5, Early stopping steps: 10
Quantile XGBoost
Same as point prediction baseline, XGBoost
MultiETs Extremely Randomized Trees [38]
Trees: 100, Maximum depth: 10
Quantile Extremely Randomized Trees [51]
Same as point prediction baseline, Extremely Randomized Trees
MultiMLPs Deep Neural Network
Hidden layers: 2, Nodes per layer: 100, Activation functions: ReLU, Epochs: 1000 or 5000, Learning rate: 0.001, Batch training, Early stopping steps: 100
Simultaneous Quantile Regression [48]
Same as point prediction baseline, Deep Neural Network
4.3 Metrics
The evaluation of point predictions employs Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and R-squared ($\text{R}^{2}$). For prediction intervals, the chosen metrics, following [52, 53], are the Prediction Interval Coverage Probability (PICP), Mean Prediction Interval Width (MPIW), and Normalized Mean Prediction Interval Width (NMPIW) as described in Subsection D.1. This paper also evaluates SEMF on our new metric, termed Coverage-Width Ratio (CWR),
$\text{CWR}=\frac{\text{PICP}}{\text{NMPIW}},$ (19)
which evaluates the coverage probability ratio to the prediction interval’s width. CWR provides a refined understanding of the balance between an interval’s accuracy (coverage) and precision (width). Though a larger value of this metric is better in higher confidence levels, the marginal increase in NMPIW is likely higher than that of PICP, resulting in decreasing CWR.
In our case, measuring the performance of SEMF over the baseline models is far more critical than reviewing absolute metrics in isolation. For any metric above, except for ($\text{R}^{2}$), this is computed as
$\text{Metric}_{\Delta}(\%)=\left(\frac{\text{Metric}_{\text{SEMF}}-\text{Metric}_{\text{Baseline}}}{\text{Metric}_{\text{Baseline}}}\right)\times 100,$ (20)
on which we base our decisions for selecting the best hyper-parameters as explained in Subsection D.2.
5 Results
We trained and tested 330 models corresponding to the three model types—MultiXGBs, MultiETs, and MultiMLPs—across 11 datasets with both complete and 50% missing data, with five seeds for each combination. Table 2 and Table 3 provide the mean and standard deviation for our metrics aggregated over the five seeds. All the results are in Appendix E. Below, we discuss the performances with complete data and with incomplete data.
5.1 Complete Data
The results of models with complete data are in Table 2. Overall, MultiXGBs and MultiMLPs performed generally well in producing intervals compared to their baselines as shown by positive $\Delta$CWR and $\Delta$PICP, notably, for instance, on cpu_activity, naval_propulsion_plant, and energy_efficiency. Interestingly, MultiMLPs also attained good relative performance improvements for point prediction, which can indicate that the chosen $\sigma_{k}$ was too low for generating performant prediction intervals. MultiETs attained mixed results as, despite significant improvements in $\Delta$CWR, it often failed to beat the baseline on $\Delta$PICP.
Table 2: Test results for all models with complete data at 95% quantiles aggregated over five seeds. For each metric, the mean and standard deviation of the performance across the seeds are separated by ±. Performance over the baseline is highlighted in bold.
Interval Predictions Point Predictions
Relative Absolute Relative Absolute
Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP MPIW NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
MultiXGBs
space_ga -1%±6% 6%±2% -7%±6% 0.89±0.01 1.80±0.12 0.20±0.01 -9%±1% -10%±2% 0.60±0.01
cpu_activity 25%±7% 9%±1% 12%±5% 0.89±0.01 0.37±0.01 0.07±0.00 21%±1% -1%±3% 0.98±0.00
naval_propulsion_plant 156%±15% 8%±1% 58%±2% 0.96±0.00 0.39±0.01 $0$14%±4% -12%±2% 0.99±0.00
miami_housing 61%±6% -7%±2% 43%±3% 0.82±0.01 0.45±0.01 0.06±0.00 4%±5% 3%±3% 0.91±0.01
kin8nm 10%±3% 5%±2% 4%±4% 0.90±0.01 1.94±0.07 0.38±0.01 -20%±2% -22%±2% 0.63±0.01
concrete_compressive_strength 13%±5% 26%±6% -11%±8% 0.91±0.01 1.25±0.03 0.25±0.01 -26%±7% -40%±13% 0.85±0.01
cars 18%±9% 15%±4% 2%±9% 0.89±0.02 0.70±0.05 0.12±0.01 -3%±3% -1%±4% 0.95±0.00
energy_efficiency 247%±99% 22%±7% 60%±20% 0.88±0.02 0.16±0.04 0.05±0.01 -15%±25% -16%±19% 1.00±0.00
california_housing 31%±5% 0%±1% 23%±4% 0.88±0.01 1.17±0.02 $0$1%±1% -1%±1% 0.81±0.00
airfoil_self_noise 61%±45% 4%±10% 31%±18% 0.81±0.07 0.89±0.22 0.19±0.05 -64%±36% -73%±46% 0.86±0.05
QSAR_fish_toxicity 10%±4% 11%±4% -1%±3% 0.76±0.01 1.46±0.03 0.23±0.00 3%±4% 1%±5% 0.55±0.02
MultiETs
space_ga 9%±4% -2%±1% 9%±4% 0.91±0.01 2.04±0.07 0.22±0.01 -16%±2% -18%±3% 0.54±0.02
cpu_activity 27%±3% -8%±0% 27%±2% 0.90±0.01 0.47±0.01 $0$14%±2% -20%±2% 0.98±0.00
naval_propulsion_plant 219%±36% -9%±0% 71%±3% 0.91±0.00 0.66±0.07 0.19±0.02 -320%±45% -409%±40% 0.96±0.01
miami_housing 69%±4% -13%±0% 48%±1% 0.86±0.00 0.65±0.02 $0$10%±2% -20%±2% 0.90±0.00
kin8nm 10%±2% -9%±0% 17%±2% 0.88±0.01 2.27±0.05 0.44±0.01 -35%±2% -38%±2% 0.48±0.01
concrete_compressive_strength 9%±10% -12%±2% 19%±9% 0.81±0.02 1.18±0.11 0.24±0.02 -67%±7% -94%±10% 0.76±0.02
cars 11%±7% 16%±2% -5%±7% 0.83±0.02 0.53±0.04 0.09±0.01 3%±4% 1%±2% 0.95±0.00
energy_efficiency -4%±6% 0%±8% -5%±15% 0.65±0.04 0.08±0.01 0.02±0.00 3%±4% 0%±2% 1.00±0.00
california_housing 31%±2% -7%±0% 29%±1% 0.90±0.00 1.70±0.03 0.41±0.01 -15%±1% -21%±1% 0.71±0.01
airfoil_self_noise 18%±28% -15%±5% 25%±18% 0.83±0.05 1.22±0.31 0.26±0.07 -117%±50% -141%±57% 0.80±0.09
QSAR_fish_toxicity 12%±2% -9%±1% 19%±2% 0.82±0.01 1.81±0.02 $0$6%±1% -11%±1% 0.53±0.01
MultiMLPs
space_ga 8%±1% 0%±3% 7%±3% 0.81±0.02 1.25±0.02 0.14±0.00 0%±1% -1%±1% 0.75±0.01
cpu_activity 8%±3% -9%±4% 15%±6% 0.72±0.02 0.24±0.01 0.05±0.00 7%±2% 5%±2% 0.98±0.00
naval_propulsion_plant 21%±20% 0%±3% 16%±15% 0.92±0.01 0.22±0.03 0.07±0.01 -45%±28% -36%±23% 1.00±0.00
miami_housing 9%±4% 2%±3% 6%±6% 0.81±0.02 0.41±0.03 $0$3%±2% 4%±2% 0.91±0.00
kin8nm 2%±8% 8%±6% -6%±12% 0.81±0.02 0.68±0.03 0.13±0.01 7%±3% 7%±2% 0.93±0.00
concrete_compressive_strength 98%±15% -23%±8% 61%±5% 0.56±0.04 0.28±0.01 0.06±0.00 12%±6% 15%±4% 0.91±0.01
cars 0%±15% 6%±13% -9%±33% 0.76±0.09 0.40±0.10 0.07±0.02 -3%±3% 0%±3% 0.95±0.00
energy_efficiency 82%±20% 25%±26% 31%±12% 0.62±0.05 0.06±0.00 0.02±0.00 32%±4% 33%±3% 1.00±0.00
california_housing -14%±3% 8%±1% -26%±4% 0.89±0.01 1.31±0.02 0.31±0.01 7%±0% 9%±1% 0.82±0.00
airfoil_self_noise 74%±46% 0%±7% 39%±17% 0.75±0.02 0.35±0.01 0.08±0.00 18%±4% 16%±5% 0.97±0.00
QSAR_fish_toxicity -6%±4% 11%±4% -19%±5% 0.76±0.04 1.49±0.06 0.23±0.01 4%±5% 7%±5% 0.55±0.04
5.2 Missing Data
The results of models with 50% missing data are in Table 3. Note that relative metrics for SEMF are compared with the best result from any baseline imputer on that metric, regardless of how the imputer performed on the other metrics. Overall, the results are less consistent than with complete data. MultiXGBs maintained strong performance on some datasets, such as naval_propulsion_plant, but declined on others, like cpu_activity. MultiETs exhibited narrow predictions, improved absolute PICP in some cases, such as energy_efficiency, and improved relative point predictions. Conversely, MultiMLPs, while offering narrow prediction intervals, had higher PICP, indicating decreased predictive power and increased model uncertainty.
Table 3: Test results for all models with missing data at 95% quantiles aggregated over five seeds. For each metric, the mean and standard deviation of the performance across the seeds are separated by ±. Performance over the baseline is highlighted in bold.
Interval Predictions Point Predictions
Relative Absolute Relative Absolute
Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP MPIW NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
MultiXGBs
space_ga -5%±6% 4%±2% -12%±7% 0.86±0.02 2.05±0.16 0.22±0.02 -7%±5% -9%±3% 0.45±0.07
cpu_activity -5%±21% 12%±9% -37%±52% 0.89±0.07 0.76±0.24 0.14±0.05 -1%±24% -17%±14% 0.83±0.08
naval_propulsion_plant 105%±33% -16%±7% 56%±6% 0.70±0.07 0.62±0.06 0.18±0.02 -14%±22% -29%±39% 0.77±0.06
miami_housing 16%±14% -5%±6% 10%±18% 0.82±0.07 0.77±0.15 0.10±0.02 -14%±12% -18%±11% 0.72±0.11
kin8nm -5%±2% 3%±2% -14%±5% 0.88±0.01 2.46±0.07 0.48±0.01 -3%±3% -4%±5% 0.40±0.02
concrete_compressive_strength 2%±3% 8%±4% -10%±4% 0.77±0.04 1.47±0.03 0.30±0.01 -16%±14% -20%±16% 0.58±0.09
cars -15%±15% 26%±10% -63%±29% 0.85±0.04 1.58±0.18 0.27±0.03 -30%±25% -35%±27% 0.66±0.17
energy_efficiency 10%±29% 11%±6% -24%±48% 0.92±0.03 0.77±0.19 0.22±0.05 -186%±296% -153%±210% 0.91±0.09
california_housing 26%±8% -2%±3% 21%±7% 0.88±0.03 1.67±0.12 0.40±0.03 -4%±4% -5%±5% 0.69±0.05
airfoil_self_noise -2%±21% -3%±4% -4%±19% 0.73±0.05 1.59±0.33 0.34±0.07 -37%±34% -43%±40% 0.41±0.13
QSAR_fish_toxicity -2%±3% 0%±5% -6%±8% 0.73±0.03 1.70±0.13 0.26±0.02 -4%±5% -2%±4% 0.36±0.05
MultiETs
space_ga 4%±4% -5%±1% 7%±4% 0.88±0.02 2.23±0.10 0.24±0.01 -12%±4% -14%±5% 0.40±0.06
cpu_activity -16%±12% -5%±2% -20%±21% 0.93±0.01 1.07±0.20 0.20±0.04 -30%±26% -54%±22% 0.79±0.11
naval_propulsion_plant 58%±5% -17%±5% 46%±3% 0.82±0.05 1.38±0.10 0.41±0.03 -36%±23% -93%±53% 0.71±0.06
miami_housing 20%±16% -10%±5% 21%±14% 0.87±0.05 1.15±0.26 0.14±0.03 -27%±12% -37%±12% 0.67±0.12
kin8nm 5%±3% -8%±2% 12%±4% 0.87±0.02 2.55±0.10 0.50±0.02 -11%±3% -13%±3% 0.32±0.03
concrete_compressive_strength 1%±6% -14%±4% 13%±2% 0.80±0.02 1.83±0.06 0.37±0.01 -35%±23% -50%±28% 0.49±0.07
cars -23%±7% -8%±6% -35%±16% 0.83±0.04 1.41±0.15 0.24±0.02 -33%±20% -31%±20% 0.66±0.10
energy_efficiency -15%±18% 0%±10% -36%±47% 0.85±0.06 0.51±0.05 0.15±0.01 -73%±82% -84%±72% 0.94±0.04
california_housing 21%±7% -7%±2% 22%±5% 0.91±0.02 2.19±0.13 0.52±0.03 -14%±3% -17%±4% 0.61±0.06
airfoil_self_noise -7%±20% -20%±2% 8%±18% 0.76±0.03 1.96±0.39 0.42±0.08 -65%±60% -93%±78% 0.33±0.12
QSAR_fish_toxicity 10%±8% -9%±3% 16%±4% 0.82±0.04 2.12±0.20 0.33±0.03 -8%±9% -10%±10% 0.39±0.10
MultiMLPs
space_ga -11%±6% -15%±2% 1%±7% 0.68±0.04 1.44±0.09 0.16±0.01 -40%±26% -34%±17% 0.21±0.25
cpu_activity -50%±18% 22%±16% -193%±132% 0.86±0.11 0.97±0.42 0.18±0.08 -15%±21% -23%±19% 0.75±0.14
naval_propulsion_plant -52%±21% -14%±27% -102%±62% 0.63±0.20 1.17±0.28 0.34±0.08 -108%±74% -183%±218% -0.03±0.53
miami_housing -33%±17% 13%±11% -92%±78% 0.85±0.06 0.99±0.32 0.12±0.04 -25%±9% -28%±12% 0.67±0.11
kin8nm -11%±6% -5%±2% -8%±6% 0.77±0.01 1.76±0.09 0.35±0.02 -16%±9% -15%±10% 0.37±0.04
concrete_compressive_strength -15%±14% -13%±31% -10%±54% 0.60±0.25 1.11±0.59 0.22±0.12 -12%±6% -12%±7% 0.54±0.06
cars -51%±5% 43%±8% -211%±38% 0.93±0.02 1.74±0.22 0.29±0.04 -28%±14% -26%±15% 0.66±0.11
energy_efficiency -13%±20% 9%±17% -30%±28% 0.66±0.15 0.33±0.15 0.09±0.04 1%±10% -1%±18% 0.95±0.03
california_housing -17%±7% 4%±4% -28%±16% 0.88±0.03 1.84±0.17 0.44±0.04 -5%±3% -3%±4% 0.61±0.06
airfoil_self_noise 2%±18% -2%±6% -1%±23% 0.75±0.08 1.15±0.41 0.25±0.09 -15%±9% -9%±9% 0.55±0.20
QSAR_fish_toxicity -17%±5% 6%±7% -30%±12% 0.76±0.04 1.82±0.18 0.28±0.03 -13%±7% -14%±9% 0.34±0.07
5.3 Discussion
The results indicate that specific models, particularly MultiXGBs, perform strongly with complete data. However, their robustness diminishes with increasing amounts of missing data. Further, we observe that MultiETs can produce narrow prediction intervals at the expense of reduced coverage. In contrast, despite the experiment design prioritizing interval estimation over point accuracy, MultiMLPs also deliver good point predictions. Precise choices, such as processing the columns separately, using the same hyper-parameters for both the complete and incomplete data, and fixing the same number of latent nodes per input ($m_{k}$), might have further constrained the models’ ability to tailor their performance to varying dataset complexities.
The stability of SEMF across different seeds, as indicated by the absolute PICP and NMPIW, contrasts with the significant variability observed in baseline models, suggesting that SEMF offers a more reliable performance framework. The method’s sampling operation, akin to cross-validation, helps combat overfitting, ensuring robust results despite variations in the data. Nonetheless, most models do not achieve the desired 95% prediction intervals. This shortfall could be addressed by integrating conformal prediction techniques with the intervals generated by SEMF, enhancing the reliability and coverage of prediction intervals across different datasets and missing data scenarios [34].
6 Conclusion
This paper introduces the Supervised Expectation-Maximization Framework (SEMF), a novel model-agnostic approach for generating prediction intervals in datasets with missing values. SEMF draws from the EM algorithm for supervised learning to devise latent representations that produce better prediction intervals than quantile regression. Due to SEMF’s iterative simulation technique, training and inference can be done with complete and incomplete data. A comprehensive set of 330 experimental runs on 11 datasets with three different model types showed that SEMF is superior to quantile regression, particularly for models such as XGBoost, which intrinsically lacks latent representations. This research underscores the potential of SEMF in various application domains and opens new avenues for further exploration for supervised latent representation learning and uncertainty estimation.
7 Limitations & Future Work
The primary limitation of this study was its reliance on the normality assumption, which may not fully capture the potential of SEMF across diverse data distributions. Although in Appendix F we demonstrate that the framework can learn non-normal patterns, further investigation and exploration of SEMF under other distributions, such as uniform and log-normal, are needed. The computational complexity of the approach presents another significant challenge, as the current implementation can be optimized for large-scale applications. Additionally, while the CWR metric is useful, it implicitly assumes that a 1% drop in PICP equates to a 1% reduction in NMPIW, thus assuming a uniform distribution. Evaluating CWR under various distributional assumptions would provide a more comprehensive assessment of its implications. Finally, applying the same hyper-parameters across all datasets without specific tuning for incomplete data likely contributes to the observed decline in accuracy and robustness.
Future work presents several intriguing avenues for exploration. A promising direction is the application of SEMF in multi-modal data settings, where the distinct $p_{\phi}$ components of the framework could be adapted to process diverse data types—from images and text to tabular datasets—enabling a more nuanced and powerful approach to integrating heterogeneous data sources. This capability positions the framework as a versatile tool for addressing missing data challenges across various domains and can also help expand it to discrete and multiple outputs. Another valuable area for development is the exploration of methods to capture and leverage dependencies among input features, which could improve the model’s predictive performance and provide deeper insights into the underlying data structure. These advancements can enhance the broader appeal of end-to-end approaches like SEMF in the ML community.
Appendix A SEMF algorithm
Algorithm 1 SEMF Training: two input sources where $x_{1}$ can be missing
1:$y,x_{1},x_{2},R$
2:$\theta,\phi_{1},\phi_{2},\xi$
3:Initialize $\theta,\phi_{1},\phi_{2},\xi$
4:Initialize $D_{y},D_{z_{1}},D_{z_{2}},D_{x},D_{i}$ to $\emptyset$
5:Split $I=\{1,\ldots,N\}$ into $L$ batches $\{b_{1},\ldots,b_{L}\}$
6:for $\ell=1$, …, $L$ do
7:   for all $i$ in $b_{\ell}$ do
8:      if $x_{1,i}$ is absent then
9:         for $r=1$, …, $R$ do
10:            Simulate $[j_{i,r},x_{1,i,r}]\sim p_{\xi}(\cdot|x_{2,i})$
11:            Simulate $z_{1,i,r}\sim p_{\phi_{1}}(\cdot|x_{1,i,r})$
12:            Simulate $z_{2,i,r}\sim p_{\phi_{2}}(\cdot|x_{2,i})$
13:            Set $z_{i,r}=[z_{1,i,r},z_{2,i,r}]$
14:         end for
15:      else
16:         for $r=1$, …, $R$ do
17:            Simulate $z_{1,i,r}\sim p_{\phi_{1}}(\cdot|x_{1,i})$
18:            Simulate $z_{2,i,r}\sim p_{\phi_{2}}(\cdot|x_{2,i})$
19:            Set $z_{i,r}=[z_{1,i,r},z_{2,i,r}]$
20:         end for
21:      end if
22:      for $r=1$, …, $R$ do
23:         Compute
$w_{i,r}=\frac{p_{\theta}(y_{i}|z_{i,r})}{\sum_{t=1}^{R}p_{\theta}(y_{i}|z_{i,t})}$
24:         Update $D_{y}\leftarrow D_{y}\cup[y_{i}|z_{i,r}|w_{i,r}]$
25:         Update $D_{z_{2}}\leftarrow D_{z_{2}}\cup[z_{2,i,r}|x_{2,i}|w_{i,r}]$
26:         if $x_{1,i}$ is absent then
27:            Update $D_{z_{1}}\leftarrow D_{z_{1}}\cup[z_{1,i,r}|x_{1,i,r}|w_{i,r}]$
28:            Update $D_{x}\leftarrow D_{x}\cup[j_{i,r}|x_{2,i}|w_{i,r}]$
29:         else
30:            Update $D_{z_{1}}\leftarrow D_{z_{1}}\cup[z_{1,i,r}|x_{1,i}|w_{i,r}]$
31:         end if
32:      end for
33:      Update $\theta\leftarrow Q_{y}(\theta,D_{y})$
34:      Update $\phi_{1}\leftarrow Q_{1}(\phi_{1},D_{z_{1}})$
35:      Update $\phi_{2}\leftarrow Q_{2}(\phi_{2},D_{z_{2}})$
36:      Update $\xi\leftarrow Q_{x}(\xi,D_{x})$
37:   end for
38:end for
39:Check convergence; Go to step 4 if not
Algorithm 2 SEMF Inference
1:$\theta^{*},\phi_{1}^{*},\phi_{2}^{*},\xi^{*},x_{1},x_{2},R$
2:$z_{i,r}$
3:for $r=1,\ldots,R$ do
4:   if $x_{1,i}$ is absent then
5:      Simulate $[j_{i,r},x_{1,i,r}]\sim p_{\xi^{*}}(\cdot|x_{2,i})$
6:      Simulate $z_{1,i,r}\sim p_{\phi_{1}^{*}}(\cdot|x_{1,i,r})$
7:      Simulate $z_{2,i,r}\sim p_{\phi_{2}^{*}}(\cdot|x_{2,i})$
8:      Set $z_{i,r}=[z_{1,i,r},z_{2,i,r}]$
9:   else
10:      Simulate $z_{1,i,r}\sim p_{\phi_{1}^{*}}(\cdot|x_{1,i})$
11:      Simulate $z_{2,i,r}\sim p_{\phi_{2}^{*}}(\cdot|x_{2,i})$
12:      Set $z_{i,r}=[z_{1,i,r},z_{2,i,r}]$
13:   end if
14:end for
Appendix B Datasets for tabular benchmark
OpenML-CTR23 [45] datasets are selected in the following manner. The first criterion is to exclude datasets exceeding 30,000 instances or 30 features to maintain computational tractability. Moreover, we exclude the moneyball data [54] to control for missing values and any datasets with non-numeric features, such as those with temporal or ordinal data not encoded numerically. We then categorize the datasets based on size: small for those with less than ten features, medium for 10 to 19 features, and large for 20 to 29 features. We apply a similar size classification based on the number of instances, considering datasets with more than 10,000 instances as large. To avoid computational constraints, we exclude datasets that were large in both features and instances, ensuring a varied yet manageable set for our experiments. This leads us to the final list of 11 datasets listed in Table 4.
Table 4: Summary of benchmark tabular datasets retained from [45]
Dataset Name N Samples N Features OpenML Data ID Y [Min:Max] Source
space_ga 3,107 7 45402 [-3.06:0.1] [55]
cpu_activity 8,192 22 44978 [0:99] [56]
naval_propulsion_plant 11,934 15 44969 [0.95:1.0] [57]
miami_housing 13,932 16 44983 [72,000:2,650,000] [58]
kin8nm 8,192 9 44980 [0.04:1.46] [59]
concrete_compressive_strength 1,030 9 44959 [2.33:82.6] [60]
cars 804 18 44994 [8,639:70,756] [61]
energy_efficiency 768 9 44960 [6.01:43.1] [62]
california_housing 20,640 9 44977 [14,999:500,001] [63]
airfoil_self_noise 1,503 6 44957 [103.38:140.98] [64]
QSAR_fish_toxicity 908 7 44970 [0.053:9.612] [65]
Appendix C Optimal set of hyper-parameters
The hyper-parameter tuning for SEMF is implemented and monitored using Weights & Biases [66]. A random search is done in the hyper-parameter space for a maximum of 500 iterations on all 11 datasets, focusing on tuning the models only on the complete datasets. Key hyper-parameters are varied across a predefined set to balance accuracy and computational efficiency. The following grid is used for hyper-parameter tuning: the number of importance sampling operations $R\in\{5,10,25,50,100\}$ (100 is omitted for MultiMLPS), nodes per latent dimension $m_{k}\in\{1,5,10,20,30\}$, and standard deviations $\sigma_{m_{k}}\in\{0.001,0.01,0.1,1.0\}$. Early stopping steps (patience) are set to five or ten, and $R_{\text{infer}}$ is explored at [30, 50, 70]. The option to run the models in parallel must be consistently enabled. Table 5 shows the optimal set of hyper-parameters. This table includes common hyper-parameters for complete and 50% datasets and another part showing $\xi_{\mathrm{nodes}}$, which is tuned manually and only relevant to the missing data.
Table 5: Hyper-parameters for MultiXGBs, MultiETs, and MultiMLPs used for both complete and missing data.
Dataset Complete and Missing Missing
R $m_{k}$ $\sigma_{k}$ Patience $R_{\mathrm{infer}}$ $\xi_{\mathrm{nodes}}$
MultiXGBs
space_ga 10 30 1.0 5 70 100
cpu_activity 5 30 1.0 5 70 50
naval_propulsion_plant 5 30 0.01 5 50 100
miami_housing 5 10 0.1 5 50 50
kin8nm 5 30 1.0 10 70 100
concrete_compressive_strength 25 30 1.0 5 70 100
cars 50 10 1.0 10 70 100
energy_efficiency 5 1 0.01 10 70 100
california_housing 5 10 0.1 10 50 100
airfoil_self_noise 25 1 0.01 10 70 100
QSAR_fish_toxicity 50 30 1.0 5 70 100
MultiETs
space_ga 10 30 1.0 10 70 100
cpu_activity 5 30 1.0 10 70 50
naval_propulsion_plant 5 30 0.01 10 50 100
miami_housing 10 10 0.1 10 50 50
kin8nm 5 30 1.0 10 70 100
concrete_compressive_strength 25 30 1.0 10 70 100
cars 100 5 0.1 10 100 100
energy_efficiency 5 1 0.01 10 70 100
california_housing 5 10 0.1 5 50 100
airfoil_self_noise 25 1 0.01 10 70 100
QSAR_fish_toxicity 50 30 1.0 10 70 100
MultiMLPs
space_ga 25 10 0.001 10 50 100
cpu_activity 5 20 0.001 5 50 50
naval_propulsion_plant 5 20 0.001 5 50 100
miami_housing 5 20 0.01 5 50 50
kin8nm 5 20 0.001 5 50 100
concrete_compressive_strength 5 30 0.001 10 50 100
cars 5 30 0.1 5 50 100
energy_efficiency 50 30 0.1 10 50 100
california_housing 5 20 0.01 5 50 100
airfoil_self_noise 25 10 0.01 10 50 100
QSAR_fish_toxicity 50 30 1.0 10 70 100
MultiXGBs and MultiMLPs benefit from early stopping to reduce computation time in complete and incomplete cases. Similarly, the baseline models for these instances use the same hyper-parameters for early stopping. Further, the number of epochs in the case of MultiMLPs is set as 1000, except for energy_efficiency and QSAR_fish_toxicity, where this is changed to 5000. Any model-specific hyperparameter we did not specify in this paper remains at the implementation’s default value (e.g., the number of leaves in XGBoost [50]). Along with the supplementary code, we provide two additional CSV files: one for the optimal hyperparameters of SEMF models and the other for the results and hyperparameters of all 330 runs.
Additional conditions are applied only for experiments with missing data. Datasets california_housing, cpu_activity, miami_housing, and naval_propulsion_plant—have a patience of five to expedite the training process. Additionally, for california_housing and cpu_activity, the $R_{\text{infer}}$ value is set to 30, while for all the other datasets, it is set to 50. We do this to ensure efficient computation, speed, and memory usage (especially for the GPU).
For training MultiXGBs and MultiETs, the computations are performed in parallel using CPU cores (Intel® Core i9-13900KF). For MultiMLPs, they are done on a GPU (NVIDIA® GeForce RTX 4090). The GPU is also consistently used for the missing data simulator and training $p_{\xi}$. All the computations are done on a machine with 32 GB of memory. The code provides further details on hardware and reproducibility.
Appendix D Metrics for prediction intervals
D.1 Common metrics
The most common metrics for evaluating prediction intervals [52, 53] are:
  • Prediction Interval Coverage Probability (PICP): This metric assesses the proportion of times the true value of the target variable falls within the constructed prediction intervals. For a set of test examples ${(x_{1},y_{1}),\ldots,(x_{N},y_{N})}$, a given level of confidence $\alpha$, and their corresponding prediction intervals ${I_{1},\ldots,I_{N}}$, the PICP is calculated as:
    $\text{PICP}=\frac{1}{N}\sum_{i=1}^{N}\mathds{1}(y_{i}\in[L_{i},U_{i}]),$ (21)
    where $U_{i}$ and $L_{i}$ are the upper and lower bounds of the predicted values for the $i$-th instance. $y_{i}$ is the actual value of the $i$-th test example, and $\mathds{1}$ is the indicator function, which equals 1 if $y_{i}$ is in the interval $[L_{i},U_{i}]$ and 0 otherwise. $0\leq\text{{PICP}}\leq 1$ where PICP closer to 1 and higher than the confidence level $\alpha$ is favored.
  • Mean Prediction Interval Width (MPIW): The average width is computed as
    $\text{MPIW}=\frac{1}{N}\sum_{i=1}^{N}(U_{i}-L_{i}),$ (22)
    which shows the sharpness or uncertainty, where $0\leq\text{{MPIW}}<\infty$ and MPIW close to 0 is preferred.
  • Normalized Mean Prediction Interval Width (NMPIW): Since MPIW varies by dataset, it can be normalized by the range of the target variable
    $\text{NMPIW}=\frac{\text{MPIW}}{\max(y)-\min(y)}$ (23)
    where $\max(y)$ and $\min(y)$ are the maximum and minimum values of the target variable, respectively. The interpretation remains the same as MPIW.
D.2 Impact of relative metrics for modeling
As our primary focus is on interval prediction, configurations demonstrating the most significant improvements in $\Delta$CWR and $\Delta$PICP are prioritized when selecting the optimal hyper-parameters. Furthermore, both $\Delta$PICP and $\Delta$CWR must be positive, indicating that we must at least have the same reliability of the baseline (PICP) with better or same interval ratios (CWR). In instances where no configuration meets the initial improvement criteria for both metrics, we relax the requirement for positive $\Delta$PICP to accept values greater than -5% and subsequently -10%, allowing us to consider configurations where SEMF significantly improves CWR, even if the PICP improvement is less marked but remains within an acceptable range for drawing comparisons.
Appendix E Full results
Table 6: Test results for MultiXGBs with complete data at 95% quantiles for seeds 0, 10, 20, 30, 40, with rows ordered by seed (ascending). Performance over the baseline is highlighted in bold.
Interval Predictions Point Predictions
Relative Absolute Relative Absolute
Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP MPIW NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
space_ga -1% 6% -7% 0.88 1.85 0.20 -10% -13% 0.60
space_ga -4% 5% -9% 0.89 1.85 0.20 -8% -10% 0.59
space_ga 1% 9% -8% 0.90 1.76 0.19 -10% -10% 0.62
space_ga -7% 5% -12% 0.90 1.92 0.21 -8% -10% 0.59
space_ga 8% 3% 4% 0.87 1.61 0.17 -9% -8% 0.60
cpu_activity 27% 7% 16% 0.88 0.37 0.07 21% 0% 0.98
cpu_activity 33% 11% 17% 0.89 0.37 0.07 23% -2% 0.98
cpu_activity 30% 10% 15% 0.88 0.36 0.07 20% -5% 0.98
cpu_activity 16% 9% 7% 0.89 0.38 0.07 20% -3% 0.98
cpu_activity 18% 9% 8% 0.90 0.38 0.07 22% 2% 0.98
naval_propulsion_plant 176% 8% 61% 0.96 0.39 0.12 -20% -14% 0.99
naval_propulsion_plant 151% 7% 58% 0.96 0.40 0.12 -9% -15% 0.99
naval_propulsion_plant 136% 8% 54% 0.96 0.40 0.12 -15% -12% 0.99
naval_propulsion_plant 156% 7% 58% 0.95 0.38 0.11 -12% -9% 0.99
naval_propulsion_plant 161% 9% 58% 0.96 0.38 0.11 -15% -11% 0.99
miami_housing 50% -5% 37% 0.82 0.46 0.06 3% -1% 0.91
miami_housing 65% -7% 44% 0.83 0.45 0.06 4% 1% 0.91
miami_housing 62% -10% 44% 0.81 0.44 0.06 7% 7% 0.90
miami_housing 62% -7% 43% 0.82 0.44 0.05 9% 5% 0.91
miami_housing 65% -9% 45% 0.81 0.44 0.05 -4% 0% 0.90
kin8nm 12% 4% 8% 0.89 1.88 0.37 -17% -20% 0.64
kin8nm 11% 5% 6% 0.89 1.91 0.37 -21% -22% 0.63
kin8nm 11% 5% 6% 0.90 1.92 0.38 -19% -21% 0.64
kin8nm 5% 7% -3% 0.91 2.06 0.40 -20% -23% 0.62
kin8nm 8% 3% 5% 0.88 1.91 0.37 -21% -24% 0.62
concrete_compressive_strength 7% 34% -26% 0.91 1.26 0.26 -34% -52% 0.86
concrete_compressive_strength 13% 22% -7% 0.89 1.25 0.25 -29% -52% 0.84
concrete_compressive_strength 13% 18% -5% 0.91 1.20 0.24 -21% -34% 0.86
concrete_compressive_strength 20% 31% -9% 0.93 1.27 0.26 -16% -22% 0.86
concrete_compressive_strength 14% 26% -10% 0.90 1.25 0.25 -28% -39% 0.83
cars 35% 14% 16% 0.89 0.67 0.11 -5% -1% 0.95
cars 16% 10% 5% 0.85 0.67 0.11 -4% -4% 0.95
cars 13% 20% -6% 0.91 0.78 0.13 -2% 0% 0.95
cars 12% 16% -4% 0.88 0.68 0.11 2% 5% 0.96
cars 16% 18% -2% 0.91 0.71 0.12 -7% -4% 0.95
energy_efficiency 255% 24% 64% 0.88 0.14 0.04 -50% -29% 1.00
energy_efficiency 320% 19% 71% 0.91 0.14 0.04 -3% -13% 1.00
energy_efficiency 257% 17% 67% 0.90 0.16 0.05 -21% -31% 1.00
energy_efficiency 322% 16% 73% 0.85 0.12 0.04 17% 16% 1.00
energy_efficiency 78% 34% 26% 0.89 0.22 0.06 -17% -22% 1.00
california_housing 24% 1% 18% 0.88 1.19 0.28 -2% -3% 0.81
california_housing 28% 1% 21% 0.88 1.16 0.28 0% -1% 0.82
california_housing 33% 0% 25% 0.89 1.19 0.28 0% 0% 0.81
california_housing 30% 1% 22% 0.87 1.16 0.28 -1% -1% 0.82
california_housing 38% -2% 29% 0.87 1.15 0.27 -1% -2% 0.81
airfoil_self_noise 68% 6% 37% 0.83 0.80 0.17 -41% -43% 0.89
airfoil_self_noise 134% 5% 55% 0.81 0.59 0.13 -18% -18% 0.93
airfoil_self_noise 13% -3% 15% 0.72 1.10 0.23 -109% -138% 0.78
airfoil_self_noise 38% 20% 13% 0.91 1.12 0.24 -73% -76% 0.85
airfoil_self_noise 51% -6% 38% 0.76 0.84 0.18 -82% -93% 0.86
QSAR_fish_toxicity 11% 17% -5% 0.77 1.48 0.23 3% 3% 0.53
QSAR_fish_toxicity 11% 9% 3% 0.75 1.42 0.22 8% 7% 0.58
QSAR_fish_toxicity 4% 8% -5% 0.77 1.48 0.23 4% 3% 0.57
QSAR_fish_toxicity 10% 9% 1% 0.77 1.45 0.23 -2% -1% 0.54
QSAR_fish_toxicity 13% 15% -1% 0.74 1.48 0.23 0% -6% 0.52
Table 7: Test results for MultiETs with complete data at 95% quantiles for seeds 0, 10, 20, 30, 40, with rows ordered by seed (ascending). Performance over the baseline is highlighted in bold.
Interval Predictions Point Predictions
Relative Absolute Relative Absolute
Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP MPIW NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
space_ga 3% 0% 3% 0.92 2.15 0.23 -19% -24% 0.51
space_ga 8% -1% 9% 0.91 2.04 0.22 -15% -18% 0.54
space_ga 10% -1% 10% 0.91 2.04 0.22 -16% -17% 0.54
space_ga 11% -2% 11% 0.91 2.01 0.22 -14% -17% 0.55
space_ga 12% -3% 13% 0.90 1.96 0.21 -14% -15% 0.55
cpu_activity 24% -8% 26% 0.91 0.48 0.09 -15% -22% 0.98
cpu_activity 24% -8% 25% 0.91 0.49 0.09 -17% -22% 0.97
cpu_activity 30% -9% 30% 0.90 0.45 0.09 -13% -18% 0.98
cpu_activity 29% -8% 28% 0.90 0.47 0.09 -13% -18% 0.98
cpu_activity 27% -8% 27% 0.91 0.48 0.09 -14% -20% 0.98
naval_propulsion_plant 265% -9% 75% 0.91 0.58 0.17 -268% -371% 0.97
naval_propulsion_plant 205% -9% 70% 0.90 0.69 0.20 -332% -419% 0.96
naval_propulsion_plant 178% -9% 67% 0.91 0.74 0.22 -382% -464% 0.95
naval_propulsion_plant 198% -9% 69% 0.91 0.69 0.20 -336% -422% 0.96
naval_propulsion_plant 248% -10% 74% 0.90 0.58 0.17 -283% -368% 0.97
miami_housing 70% -12% 48% 0.86 0.66 0.08 -12% -22% 0.90
miami_housing 63% -12% 46% 0.86 0.68 0.08 -9% -21% 0.90
miami_housing 68% -13% 48% 0.86 0.65 0.08 -9% -20% 0.90
miami_housing 71% -13% 49% 0.86 0.63 0.08 -7% -18% 0.90
miami_housing 72% -12% 49% 0.86 0.65 0.08 -12% -21% 0.89
kin8nm 12% -9% 19% 0.88 2.20 0.43 -32% -36% 0.50
kin8nm 9% -9% 17% 0.88 2.28 0.45 -36% -40% 0.46
kin8nm 11% -10% 18% 0.88 2.24 0.44 -34% -37% 0.49
kin8nm 9% -9% 16% 0.89 2.29 0.45 -34% -36% 0.49
kin8nm 7% -9% 15% 0.88 2.33 0.46 -37% -40% 0.47
concrete_compressive_strength 25% -15% 32% 0.79 1.02 0.21 -59% -85% 0.78
concrete_compressive_strength 9% -12% 19% 0.81 1.19 0.24 -71% -96% 0.76
concrete_compressive_strength 4% -11% 14% 0.83 1.22 0.25 -62% -92% 0.77
concrete_compressive_strength -2% -9% 8% 0.83 1.31 0.27 -77% -110% 0.74
concrete_compressive_strength 11% -12% 20% 0.81 1.14 0.23 -64% -88% 0.77
cars 3% 15% -11% 0.83 0.54 0.09 3% 2% 0.95
cars 16% 18% -2% 0.85 0.52 0.09 3% 0% 0.95
cars 21% 13% 7% 0.80 0.47 0.08 4% -1% 0.95
cars 6% 17% -10% 0.84 0.56 0.10 -2% 0% 0.94
cars 10% 17% -7% 0.85 0.54 0.09 8% 5% 0.95
energy_efficiency -3% -3% 0% 0.63 0.07 0.02 3% 0% 1.00
energy_efficiency -3% 1% -5% 0.66 0.08 0.02 7% 3% 1.00
energy_efficiency -14% 9% -29% 0.70 0.09 0.03 0% 0% 1.00
energy_efficiency 2% -11% 14% 0.59 0.07 0.02 -3% -2% 1.00
energy_efficiency -5% 1% -5% 0.65 0.08 0.02 6% 0% 1.00
california_housing 33% -7% 30% 0.91 1.71 0.41 -14% -19% 0.72
california_housing 28% -7% 27% 0.91 1.74 0.41 -16% -22% 0.70
california_housing 34% -7% 31% 0.90 1.67 0.40 -14% -20% 0.72
california_housing 29% -7% 28% 0.91 1.72 0.41 -16% -22% 0.71
california_housing 31% -8% 30% 0.90 1.69 0.40 -15% -20% 0.71
airfoil_self_noise 8% -21% 27% 0.77 1.20 0.26 -168% -189% 0.74
airfoil_self_noise -16% -14% -3% 0.84 1.68 0.36 -170% -214% 0.67
airfoil_self_noise 20% -8% 24% 0.89 1.22 0.26 -95% -105% 0.84
airfoil_self_noise 62% -18% 49% 0.81 0.82 0.18 -57% -81% 0.89
airfoil_self_noise 18% -13% 26% 0.86 1.19 0.25 -93% -118% 0.85
QSAR_fish_toxicity 11% -10% 19% 0.81 1.82 0.28 -6% -13% 0.53
QSAR_fish_toxicity 13% -10% 21% 0.82 1.79 0.28 -5% -11% 0.54
QSAR_fish_toxicity 14% -10% 21% 0.82 1.82 0.28 -8% -12% 0.51
QSAR_fish_toxicity 11% -8% 17% 0.82 1.79 0.28 -5% -11% 0.53
QSAR_fish_toxicity 9% -8% 16% 0.84 1.84 0.29 -6% -10% 0.52
Table 8: Test results for MultiMLPs with complete data at 95% quantiles for seeds 0, 10, 20, 30, 40, with rows ordered by seed (ascending). Performance over the baseline is highlighted in bold.
Interval Predictions Point Predictions
Relative Absolute Relative Absolute
Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP MPIW NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
space_ga 7% -4% 10% 0.79 1.22 0.13 -1% -1% 0.75
space_ga 8% 2% 5% 0.82 1.25 0.14 0% 0% 0.76
space_ga 7% 4% 3% 0.83 1.27 0.14 -3% -3% 0.74
space_ga 8% -1% 8% 0.79 1.24 0.14 0% -1% 0.74
space_ga 9% -2% 10% 0.81 1.27 0.14 1% 0% 0.76
cpu_activity 8% -11% 16% 0.72 0.25 0.05 6% 5% 0.98
cpu_activity 14% -15% 25% 0.68 0.23 0.04 8% 5% 0.98
cpu_activity 8% -8% 15% 0.73 0.24 0.04 8% 7% 0.98
cpu_activity 5% -3% 8% 0.74 0.25 0.05 6% 5% 0.98
cpu_activity 6% -8% 12% 0.70 0.24 0.04 4% 2% 0.98
naval_propulsion_plant 50% 0% 34% 0.92 0.19 0.06 -44% -48% 1.00
naval_propulsion_plant 20% 0% 17% 0.93 0.20 0.06 -35% -13% 1.00
naval_propulsion_plant 3% 1% 3% 0.92 0.26 0.08 -51% -45% 0.99
naval_propulsion_plant 3% 3% 0% 0.94 0.23 0.07 -86% -63% 1.00
naval_propulsion_plant 29% -6% 27% 0.91 0.23 0.07 -9% -11% 1.00
miami_housing 13% -2% 13% 0.78 0.37 0.05 -2% 7% 0.91
miami_housing 10% 3% 6% 0.81 0.40 0.05 -6% 1% 0.91
miami_housing 14% 1% 10% 0.82 0.42 0.05 -2% 3% 0.92
miami_housing 8% 1% 7% 0.82 0.43 0.05 -1% 4% 0.91
miami_housing 3% 6% -4% 0.83 0.44 0.06 -4% 6% 0.91
kin8nm -5% 9% -15% 0.81 0.69 0.13 6% 5% 0.93
kin8nm 15% 1% 12% 0.78 0.62 0.12 7% 7% 0.93
kin8nm 2% 4% -2% 0.82 0.67 0.13 10% 10% 0.94
kin8nm 2% 14% -12% 0.83 0.69 0.14 10% 6% 0.94
kin8nm -3% 12% -15% 0.83 0.71 0.14 4% 5% 0.93
concrete_compressive_strength 89% -14% 54% 0.62 0.30 0.06 4% 10% 0.90
concrete_compressive_strength 82% -21% 57% 0.55 0.29 0.06 9% 15% 0.91
concrete_compressive_strength 120% -17% 62% 0.60 0.28 0.06 18% 19% 0.92
concrete_compressive_strength 100% -33% 66% 0.52 0.27 0.06 9% 12% 0.91
concrete_compressive_strength 101% -30% 65% 0.52 0.28 0.06 18% 20% 0.91
cars 10% 3% 7% 0.77 0.38 0.06 -4% 5% 0.95
cars -23% 28% -65% 0.88 0.54 0.09 1% 1% 0.96
cars 16% -4% 17% 0.65 0.29 0.05 -6% -2% 0.95
cars 2% -6% 8% 0.68 0.34 0.06 -2% -1% 0.95
cars -5% 8% -13% 0.81 0.45 0.08 -5% -2% 0.95
energy_efficiency 93% 59% 17% 0.70 0.07 0.02 37% 38% 1.00
energy_efficiency 83% 27% 29% 0.61 0.06 0.02 28% 32% 1.00
energy_efficiency 109% 28% 39% 0.64 0.06 0.02 35% 32% 1.00
energy_efficiency 65% 26% 24% 0.58 0.07 0.02 31% 31% 1.00
energy_efficiency 60% -15% 47% 0.58 0.06 0.02 31% 32% 1.00
california_housing -13% 9% -25% 0.90 1.33 0.32 7% 8% 0.81
california_housing -15% 6% -24% 0.89 1.27 0.30 7% 9% 0.82
california_housing -11% 9% -23% 0.90 1.29 0.31 8% 11% 0.82
california_housing -14% 8% -26% 0.89 1.31 0.31 7% 8% 0.81
california_housing -18% 9% -33% 0.90 1.32 0.32 7% 9% 0.82
airfoil_self_noise 140% -10% 62% 0.74 0.34 0.07 18% 18% 0.97
airfoil_self_noise 42% 9% 23% 0.73 0.34 0.07 11% 9% 0.97
airfoil_self_noise 43% -1% 31% 0.77 0.36 0.08 22% 22% 0.97
airfoil_self_noise 39% 3% 26% 0.73 0.35 0.07 17% 12% 0.97
airfoil_self_noise 105% -2% 52% 0.77 0.38 0.08 21% 19% 0.97
QSAR_fish_toxicity -6% 8% -14% 0.72 1.40 0.22 -4% 6% 0.52
QSAR_fish_toxicity 0% 17% -17% 0.82 1.50 0.23 6% 11% 0.54
QSAR_fish_toxicity -5% 8% -13% 0.77 1.57 0.24 0% -1% 0.57
QSAR_fish_toxicity -8% 13% -23% 0.78 1.51 0.23 10% 9% 0.61
QSAR_fish_toxicity -12% 10% -25% 0.73 1.49 0.23 5% 11% 0.51
Table 9: Test results for MultiXGBs with 50% missing data at 95% quantiles for seeds 0, 10, 20, 30, 40, with rows ordered by seed (ascending). Performance over the baseline is highlighted in bold.
Interval Predictions Point Predictions
Relative Absolute Relative Absolute
Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP MPIW NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
space_ga -12% 5% -20% 0.87 2.25 0.24 -11% -13% 0.35
space_ga -9% 6% -18% 0.87 2.15 0.23 -12% -11% 0.44
space_ga 3% 6% -5% 0.85 1.90 0.21 -1% -6% 0.46
space_ga -8% 4% -14% 0.89 2.06 0.22 -5% -11% 0.54
space_ga 0% 1% -4% 0.82 1.88 0.20 -6% -6% 0.45
cpu_activity 2% 2% -1% 0.79 0.58 0.11 -30% -30% 0.73
cpu_activity 9% 7% -12% 0.87 0.61 0.12 -7% -23% 0.94
cpu_activity -41% 27% -127% 0.96 1.17 0.22 -1% -26% 0.81
cpu_activity -8% 14% -34% 0.94 0.78 0.15 36% 5% 0.88
cpu_activity 12% 8% -12% 0.91 0.66 0.12 -2% -11% 0.81
naval_propulsion_plant 82% -9% 47% 0.76 0.64 0.19 -30% -42% 0.83
naval_propulsion_plant 121% -13% 59% 0.76 0.65 0.19 5% -6% 0.84
naval_propulsion_plant 127% -13% 58% 0.73 0.66 0.20 6% 6% 0.76
naval_propulsion_plant 134% -21% 63% 0.64 0.52 0.15 -7% -11% 0.70
naval_propulsion_plant 58% -25% 53% 0.62 0.64 0.19 -45% -90% 0.73
miami_housing 16% -6% 13% 0.84 0.86 0.11 -4% -4% 0.72
miami_housing -9% 4% -22% 0.90 0.94 0.12 -14% -20% 0.79
miami_housing 25% -11% 17% 0.72 0.60 0.07 -35% -35% 0.54
miami_housing 23% -4% 21% 0.85 0.83 0.10 -7% -12% 0.80
miami_housing 25% -9% 21% 0.78 0.63 0.08 -10% -19% 0.76
kin8nm -3% 2% -12% 0.88 2.41 0.47 0% 1% 0.40
kin8nm -6% 3% -14% 0.89 2.47 0.48 -6% -8% 0.39
kin8nm -4% 2% -10% 0.88 2.48 0.48 -5% -6% 0.37
kin8nm -9% 7% -23% 0.90 2.57 0.50 -5% -7% 0.44
kin8nm -3% 3% -10% 0.87 2.38 0.47 1% 2% 0.41
concrete_compressive_strength -2% 0% -6% 0.71 1.51 0.30 -19% -22% 0.48
concrete_compressive_strength 2% 11% -15% 0.81 1.44 0.29 -4% -3% 0.69
concrete_compressive_strength 4% 11% -12% 0.76 1.44 0.29 -15% -24% 0.50
concrete_compressive_strength 1% 8% -9% 0.77 1.47 0.30 -4% -9% 0.64
concrete_compressive_strength 7% 8% -7% 0.80 1.49 0.30 -37% -44% 0.57
cars -3% 24% -34% 0.85 1.34 0.23 -9% -16% 0.75
cars -35% 9% -98% 0.82 1.78 0.30 -68% -80% 0.55
cars -6% 29% -52% 0.88 1.54 0.26 -8% -14% 0.76
cars -29% 32% -90% 0.81 1.73 0.29 -28% -39% 0.40
cars -5% 34% -40% 0.92 1.49 0.25 -37% -25% 0.81
energy_efficiency 20% 15% -13% 0.93 0.71 0.21 -37% -39% 0.94
energy_efficiency 35% 3% 23% 0.88 0.58 0.17 -15% -24% 0.97
energy_efficiency 37% 7% 19% 0.92 0.63 0.18 -16% -44% 0.96
energy_efficiency -21% 18% -81% 0.95 1.01 0.29 -155% -137% 0.93
energy_efficiency -22% 14% -66% 0.91 0.92 0.27 -705% -520% 0.76
california_housing 30% -8% 28% 0.82 1.52 0.36 -4% -7% 0.61
california_housing 26% -2% 21% 0.89 1.61 0.39 0% 1% 0.73
california_housing 34% -1% 26% 0.90 1.66 0.40 -11% -13% 0.73
california_housing 13% 0% 11% 0.90 1.85 0.44 -3% -3% 0.65
california_housing 27% 0% 21% 0.91 1.70 0.41 -1% -3% 0.73
airfoil_self_noise -8% 4% -14% 0.77 1.70 0.36 -14% -8% 0.52
airfoil_self_noise -18% -2% -20% 0.80 1.77 0.38 -94% -110% 0.50
airfoil_self_noise -12% -7% -11% 0.70 1.86 0.40 -43% -48% 0.19
airfoil_self_noise 34% -5% 29% 0.69 1.02 0.22 -19% -29% 0.38
airfoil_self_noise -6% -6% -5% 0.71 1.58 0.34 -14% -22% 0.43
QSAR_fish_toxicity -3% -4% -1% 0.68 1.56 0.24 -4% -5% 0.38
QSAR_fish_toxicity -1% -4% 2% 0.75 1.83 0.28 -9% -3% 0.35
QSAR_fish_toxicity 3% 9% -8% 0.75 1.66 0.26 4% 4% 0.43
QSAR_fish_toxicity -6% 1% -19% 0.72 1.61 0.25 -7% -3% 0.35
QSAR_fish_toxicity -1% -3% -2% 0.74 1.84 0.29 -5% -3% 0.30
Table 10: Test results for MultiETs with 50% missing data at 95% quantiles for seeds 0, 10, 20, 30, 40, with rows ordered by seed (ascending). Performance over the baseline is highlighted in bold.
Interval Predictions Point Predictions
Relative Absolute Relative Absolute
Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP MPIW NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
space_ga 7% -6% 12% 0.86 2.17 0.23 -17% -18% 0.30
space_ga 3% -5% 8% 0.87 2.14 0.23 -14% -15% 0.41
space_ga 2% -5% 5% 0.88 2.23 0.24 -9% -10% 0.41
space_ga 7% -5% 9% 0.86 2.21 0.24 -7% -7% 0.40
space_ga -1% -3% 0% 0.91 2.39 0.26 -14% -18% 0.47
cpu_activity -8% -5% -8% 0.93 0.90 0.17 -36% -46% 0.93
cpu_activity -8% -6% -4% 0.92 0.89 0.17 2% -23% 0.87
cpu_activity -30% -3% -43% 0.94 1.33 0.25 -36% -80% 0.73
cpu_activity -27% -3% -41% 0.94 1.21 0.23 -13% -49% 0.74
cpu_activity -6% -6% -3% 0.91 1.00 0.19 -65% -70% 0.67
naval_propulsion_plant 55% -22% 48% 0.77 1.29 0.38 -18% -50% 0.64
naval_propulsion_plant 55% -10% 42% 0.89 1.52 0.45 -66% -158% 0.79
naval_propulsion_plant 54% -15% 44% 0.84 1.44 0.42 -29% -73% 0.74
naval_propulsion_plant 67% -16% 50% 0.83 1.29 0.38 -56% -142% 0.74
naval_propulsion_plant 57% -20% 48% 0.78 1.38 0.41 -14% -42% 0.66
miami_housing 20% -6% 22% 0.92 1.33 0.17 -13% -26% 0.78
miami_housing 32% -16% 31% 0.79 0.87 0.11 -41% -50% 0.48
miami_housing -4% -5% -1% 0.92 1.39 0.17 -38% -51% 0.67
miami_housing 37% -14% 36% 0.84 0.88 0.11 -21% -34% 0.79
miami_housing 16% -7% 16% 0.89 1.30 0.16 -20% -26% 0.65
kin8nm 5% -9% 13% 0.86 2.48 0.49 -10% -12% 0.32
kin8nm 4% -8% 9% 0.87 2.58 0.51 -10% -12% 0.33
kin8nm 5% -8% 12% 0.87 2.62 0.51 -13% -15% 0.30
kin8nm 8% -11% 17% 0.84 2.41 0.47 -8% -8% 0.30
kin8nm 1% -5% 6% 0.90 2.66 0.52 -14% -18% 0.37
concrete_compressive_strength 4% -11% 15% 0.79 1.88 0.38 -24% -36% 0.46
concrete_compressive_strength -3% -17% 12% 0.78 1.74 0.35 -26% -36% 0.42
concrete_compressive_strength 4% -12% 13% 0.81 1.83 0.37 -30% -47% 0.56
concrete_compressive_strength 7% -11% 15% 0.83 1.81 0.37 -19% -31% 0.58
concrete_compressive_strength -7% -18% 9% 0.80 1.87 0.38 -76% -99% 0.45
cars -18% 1% -39% 0.89 1.29 0.22 -16% -15% 0.81
cars -18% -11% -11% 0.84 1.39 0.23 -23% -30% 0.72
cars -33% -6% -54% 0.84 1.63 0.28 -59% -59% 0.60
cars -28% -12% -41% 0.80 1.27 0.21 -18% -10% 0.59
cars -19% -14% -29% 0.79 1.48 0.25 -49% -42% 0.59
energy_efficiency -11% 13% -35% 0.91 0.50 0.14 -7% -13% 0.98
energy_efficiency 0% -7% 0% 0.83 0.49 0.14 -8% -20% 0.95
energy_efficiency -9% -14% 0% 0.76 0.47 0.14 -112% -137% 0.93
energy_efficiency -11% 4% -30% 0.87 0.59 0.17 -38% -76% 0.96
energy_efficiency -46% 2% -114% 0.89 0.52 0.15 -198% -176% 0.86
california_housing 30% -7% 28% 0.92 2.09 0.50 -11% -13% 0.64
california_housing 28% -9% 28% 0.87 2.02 0.48 -14% -18% 0.50
california_housing 13% -6% 17% 0.92 2.24 0.53 -19% -23% 0.64
california_housing 19% -5% 20% 0.93 2.23 0.53 -12% -16% 0.64
california_housing 16% -6% 18% 0.93 2.35 0.56 -11% -16% 0.60
airfoil_self_noise -7% -18% 10% 0.77 2.02 0.43 -27% -51% 0.36
airfoil_self_noise -26% -18% -11% 0.80 2.21 0.47 -171% -229% 0.30
airfoil_self_noise 25% -20% 36% 0.74 1.28 0.27 -28% -38% 0.41
airfoil_self_noise -22% -24% -2% 0.73 2.26 0.48 -54% -86% 0.13
airfoil_self_noise -7% -18% 10% 0.79 2.02 0.43 -44% -58% 0.43
QSAR_fish_toxicity 11% -8% 15% 0.85 2.37 0.37 -11% -10% 0.40
QSAR_fish_toxicity 13% -8% 18% 0.82 2.02 0.31 -1% -5% 0.47
QSAR_fish_toxicity 19% -7% 22% 0.84 1.90 0.29 -3% -3% 0.48
QSAR_fish_toxicity -3% -15% 12% 0.76 2.02 0.32 -24% -27% 0.25
QSAR_fish_toxicity 8% -9% 14% 0.83 2.29 0.36 -3% -5% 0.36
Table 11: Test results for MultiMLPs with 50% missing data at 95% quantiles for seeds 0, 10, 20, 30, 40, with rows ordered by seed (ascending). Performance over the baseline is highlighted in bold.
Interval Predictions Point Predictions
Relative Absolute Relative Absolute
Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP MPIW NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
space_ga -12% -16% 4% 0.68 1.50 0.16 -41% -40% 0.15
space_ga -21% -17% -11% 0.64 1.41 0.15 -85% -61% -0.21
space_ga -5% -14% 6% 0.65 1.30 0.14 -21% -20% 0.37
space_ga -9% -13% 4% 0.70 1.50 0.16 -28% -25% 0.35
space_ga -9% -13% 1% 0.73 1.49 0.16 -25% -26% 0.38
cpu_activity -38% 14% -89% 0.87 0.75 0.14 14% 1% 0.92
cpu_activity -60% 32% -240% 0.94 1.08 0.20 -20% -17% 0.78
cpu_activity -67% 36% -355% 0.95 1.50 0.28 -30% -34% 0.72
cpu_activity -24% -2% -28% 0.68 0.39 0.07 -3% -16% 0.78
cpu_activity -63% 28% -251% 0.85 1.13 0.21 -37% -49% 0.54
naval_propulsion_plant -41% 0% -70% 0.63 0.85 0.25 -62% -68% -0.21
naval_propulsion_plant -53% -13% -87% 0.82 1.60 0.47 -141% -163% 0.50
naval_propulsion_plant -48% -3% -85% 0.72 1.23 0.36 -99% -94% 0.16
naval_propulsion_plant -88% -61% -210% 0.29 1.10 0.33 -215% -562% -0.85
naval_propulsion_plant -33% 4% -55% 0.70 1.04 0.31 -24% -27% 0.27
miami_housing -48% 22% -148% 0.93 1.33 0.17 -21% -20% 0.74
miami_housing -54% 27% -202% 0.89 1.33 0.17 -41% -48% 0.53
miami_housing -19% 11% -40% 0.84 0.75 0.09 -18% -19% 0.75
miami_housing -17% 6% -32% 0.82 0.69 0.09 -23% -22% 0.77
miami_housing -28% 1% -39% 0.76 0.82 0.10 -24% -32% 0.57
kin8nm -7% -5% -3% 0.75 1.67 0.33 -12% -10% 0.39
kin8nm -11% -6% -8% 0.77 1.80 0.35 -20% -22% 0.36
kin8nm -19% -4% -19% 0.77 1.88 0.37 -20% -19% 0.31
kin8nm -4% -2% -5% 0.77 1.68 0.33 -3% 0% 0.43
kin8nm -15% -9% -8% 0.78 1.79 0.35 -26% -24% 0.36
concrete_compressive_strength -15% 5% -26% 0.74 1.53 0.31 -9% -6% 0.51
concrete_compressive_strength -5% -38% 34% 0.40 0.57 0.12 -10% -10% 0.46
concrete_compressive_strength -32% 21% -80% 0.84 1.69 0.34 -5% -9% 0.60
concrete_compressive_strength 2% -53% 54% 0.27 0.39 0.08 -20% -23% 0.52
concrete_compressive_strength -25% 0% -33% 0.76 1.37 0.28 -15% -13% 0.59
cars -50% 38% -176% 0.93 1.62 0.27 -36% -38% 0.75
cars -44% 49% -169% 0.93 1.55 0.26 -3% -5% 0.75
cars -53% 32% -232% 0.94 2.04 0.34 -34% -33% 0.70
cars -49% 44% -223% 0.89 1.59 0.27 -32% -15% 0.48
cars -57% 51% -257% 0.93 1.90 0.32 -32% -37% 0.63
energy_efficiency -35% 4% -68% 0.81 0.43 0.13 -8% 1% 0.97
energy_efficiency -12% 25% -47% 0.60 0.27 0.08 8% 2% 0.92
energy_efficiency -19% -14% -5% 0.57 0.31 0.09 -1% -32% 0.93
energy_efficiency -18% 7% -29% 0.83 0.50 0.14 -8% 6% 0.96
energy_efficiency 20% 26% -3% 0.51 0.12 0.03 14% 18% 0.98
california_housing -16% 1% -20% 0.88 1.76 0.42 -7% -3% 0.63
california_housing -14% 4% -25% 0.83 1.66 0.40 -4% -3% 0.54
california_housing -9% 0% -10% 0.89 1.79 0.43 -1% 0% 0.68
california_housing -20% 5% -30% 0.90 1.91 0.45 -2% -1% 0.65
california_housing -27% 10% -52% 0.89 2.09 0.50 -8% -9% 0.55
airfoil_self_noise 3% -1% 0% 0.76 1.22 0.26 -3% 1% 0.54
airfoil_self_noise 27% -7% 26% 0.77 0.84 0.18 -20% -11% 0.84
airfoil_self_noise 8% -7% 14% 0.62 0.63 0.14 -22% -13% 0.31
airfoil_self_noise -9% 1% -12% 0.76 1.44 0.31 -23% -20% 0.44
airfoil_self_noise -20% 6% -33% 0.83 1.62 0.35 -9% -1% 0.62
QSAR_fish_toxicity -17% -1% -22% 0.75 1.89 0.29 -21% -28% 0.28
QSAR_fish_toxicity -24% 9% -45% 0.79 1.97 0.31 -8% -9% 0.44
QSAR_fish_toxicity -19% 0% -23% 0.70 1.56 0.24 -6% -8% 0.39
QSAR_fish_toxicity -17% 16% -41% 0.76 1.72 0.27 -21% -17% 0.28
QSAR_fish_toxicity -9% 8% -18% 0.81 1.98 0.31 -10% -9% 0.32
Appendix F Example of learning with non-normality
To illustrate how SEMF adapts to non-normal outcomes, we provide an example from the naval_propulsion_plant dataset [57]. Figure 2 shows the distribution of the ground-truth $y$ variable which in this case is gt_compressor_decay_state_coefficient. The values are uniformly distributed, and we only standardize the values without changing the shape of the distribution.
Figure 2: Distribution of the standardized outcome ($y$) variable for the naval_propulsion_plant dataset which shows that $y$ is uniformly distributed prior to any training.
After training our SEMF model under the normality assumption with the ideal hyper-parameters (and a seed of 0), sampling from a normal distribution for the $z$ dimension, we infer on some randomly sampled test instances that provide us with the prediction intervals in Figure 3. The ‘SEMF intervals’ can be compared with XGBoost quantile regression, constituting our baseline. This figure shows that SEMF’s predicted intervals are better than the baseline. This plot alone does not tell us much about the predicted output distribution. Therefore, we provide Figure 4. The last plot shows that for a handful of the instances, the predicted values can take any shape and are not necessarily normal.
Figure 3: Predicted intervals and true values on 100 randomly selected test samples for SEMF MultiXGBs and XGBoost quantile regression with complete data at 95% quantiles (according to Eq 13).
Figure 4: Violin plot for the first ten instances of the test set where both the ground truth and 50 values inferred by SEMF (according to Eq 12) are generated for each instance. We have added jitter here, so the points do not perfectly align along the x-axis.

### v3

@@
 Abstract
-This work introduces the Supervised Expectation-Maximization Framework (SEMF), a versatile and model-agnostic framework that generates prediction intervals for datasets with complete or missing data.
-SEMF extends the Expectation-Maximization (EM) algorithm, traditionally used in unsupervised learning, to a supervised context, enabling it to extract latent representations for uncertainty estimation.
-The framework demonstrates robustness through extensive empirical evaluation across 11 tabular datasets, achieving—in some cases—narrower normalized prediction intervals and higher coverage than traditional quantile regression methods [^1] Code: https://github.com/Unco3892/semf-paper.
-Furthermore, SEMF integrates seamlessly with existing machine learning algorithms, such as gradient-boosted trees and neural networks, exemplifying its usefulness for real-world applications.
-The experimental results highlight SEMF’s potential to advance state-of-the-art techniques in uncertainty quantification.
+This work introduces the Supervised Expectation-Maximization Framework (SEMF), a versatile and model-agnostic approach for generating prediction intervals in datasets with complete or missing data.
+SEMF extends the Expectation-Maximization algorithm, traditionally used in unsupervised learning, to a supervised context, leveraging latent variable modeling for uncertainty estimation.
+Extensive empirical evaluations across 11 tabular datasets show that SEMF often achieves narrower normalized prediction intervals and higher coverage rates than traditional quantile regression methods [^1] Code: https://github.com/Unco3892/semf-paper.
+Furthermore, SEMF can be integrated with machine learning models like gradient-boosted trees and neural networks, highlighting its practical applicability.
+The results indicate that SEMF enhances uncertainty quantification, particularly in scenarios with complete data.
 1 Introduction
-In the evolving field of machine learning (ML), the quest for models capable of predicting outcomes while quantifying the uncertainty of their predictions is critical.
-Estimating prediction uncertainty is particularly vital in high-stakes domains such as healthcare [1], finance [2], and autonomous systems [3], where prediction-based decisions can have significant consequences.
-Traditional approaches have primarily focused on point estimates, providing limited insight into the reliability of these predictions.
-This limitation underscores the necessity for frameworks that generate precise point predictions and robust prediction intervals.
-These intervals offer a range within which the true outcome is expected to lie with a specified probability, enabling a deeper understanding of prediction uncertainty.
-Consequently, research has increasingly focused on methodologies that extend beyond point estimation to include uncertainty quantification, thereby enhancing decision-making in applications that rely on predictive modeling [4].
+In the evolving field of machine learning (ML), the quest for models able to predict outcomes while quantifying the uncertainty of their predictions is critical.
+The ability to estimate prediction uncertainty is particularly vital in high-stakes domains such as healthcare [1], finance [2], and autonomous systems [3], where prediction-based decisions have important consequences.
+Traditional approaches have primarily focused on point estimates, with little to no insight into prediction reliability.
+This limitation underscores the need for frameworks that can generate both precise point predictions and robust prediction intervals.
+Such intervals provide a range within which the true outcome is expected to lie with a fixed probability, offering a finer understanding of prediction uncertainty.
+This need has spurred research into methodologies that extend beyond point estimation to include uncertainty quantification, thereby enabling more informed decision-making in applications reliant on predictive modeling [4].
 In this paper, we introduce the Supervised Expectation-Maximization Framework (SEMF) based on the Expectation-Maximization (EM) algorithm [5].
 Traditionally recognized as a clustering technique, EM is used for supervised learning in SEMF, allowing for both point estimates and prediction intervals using any ML model (model-agnostic).
-SEMF generates representations for latent or missing modalities, making it particularly relevant for incomplete multi-modal data.
+SEMF generates representations for latent or missing modalities, which can be relevant for incomplete data and holds potential for multi-modal data applications, though multi-modal settings are left for future exploration.
 This paper details the methodology behind the framework and proposes a training algorithm based on Monte Carlo (MC) sampling, also used in variational inference for Variational Auto-Encoders (VAEs) [6, 7].
-Comparisons with other models across several datasets showcase its contribution to the existing uncertainty quantification and predictive modeling literature [8].
+SEMF differs from prominent supervised EM approaches such as [8], which focus on point prediction using Gaussian Mixture Models (GMMs).
+Additionally, our method operates in a frequentist paradigm, directly maximizing the likelihood function through iterative EM steps without integrating over posterior distributions.
+Although SEMF can be extended to a Bayesian framework as its likelihood component, this extension lies beyond the scope of this paper.
 The remainder of this paper is organized as follows: Section 2 details the theory and the methodology of SEMF.
@@
 The M-step maximizes this $Q$-function: $p^{\prime}\leftarrow\arg\max_{p}Q(p|p^{\prime})$ The sequence is repeated until convergence.
-The EM algorithm has already been used for supervised learning tasks using specific models [10, 11, 12, 13], and, to the best of our knowledge, mainly focusing on point prediction through the use of Gaussian Mixture Models (GMMs).
-Our work differs by modifying the EM algorithm and MC sampling to generate prediction intervals with any ML model and, in theory, under any distribution.
-This results in SEMF, a framework consisting of an encoder-decoder architecture and a missing data simulator.
 2.1 Problem Scenario
@@
 11.
-The above $D_{x}$ is a subset of the data for $\xi$ is only learned on the missing data part.
-After discussing the training steps in Subsection 2.3, the details are given in Algorithm 1 (Appendix A).
-We also note that the parameters above can be learned in parallel.
+The $D_{x}$ above is a subset of the data for $\xi$, learned on the missing data part.
+We also note that learning the parameters above is parallelizable.
 2.3 Training
@@
 The process iterates for each batch until the maximum number of steps is reached or an early stopping criterion is satisfied.
+The full details are given in Algorithm 1 (Appendix A).
 The framework requires tuning hyper-parameters such as the number of MC samples $R$, the number of latent nodes $m_{k}$, and the standard deviation $\sigma_{k}$ of $Z_{k}$.
@@
 Latent representation learning involves modeling hidden variables from observed data for various ML tasks, most notably Auto-Encoders (AEs) and VAEs.
-AEs are neural networks reconstructing inputs by learning an intermediate latent representation.
-The leftmost part of AE is the encoder $g_{\phi}(\cdot)$, which embeds the input $x$ into a latent variable $z$.
-It is passed to the rightmost part of AE, the decoder $f_{\theta}(\cdot)$, for the reconstruction $\hat{x}=f_{\theta}(z)$ of $x$ through $z$.
-The parameters $\phi$ and $\theta$ are trained by minimizing the reconstruction loss.
-Unlike AEs, which focus on reconstruction, VAEs use variational methods to fit distributions of latent variables and of the output [7].
-Given a sample $x$ with a latent variable of $z$, VAEs learn the marginal likelihood, $p_{\theta}(x)$.
-Direct maximization of this likelihood is difficult [14].
+AEs are neural networks that reconstruct inputs by learning an intermediate latent representation.
+The encoder $g_{\phi}(\cdot)$ in an AE embeds the input $x$ into a latent variable $z$, which then passes through the decoder $f_{\theta}(\cdot)$, reconstructing $x$ as $\hat{x}=f_{\theta}(z)$.
+The training process optimizes the parameters $\phi$ and $\theta$ by minimizing the reconstruction loss.
+Unlike AEs, which focus on reconstruction, VAEs use variational methods to fit distributions of latent variables and the output [7].
+Given a sample $x$ with a latent variable $z$, VAEs model the marginal likelihood $p_{\theta}(x)$.
+However, directly maximizing this likelihood is difficult [10].
 Thus, alternative methods exist, such as maximizing the evidence lower bound (ELBO), which provides guarantees on the log-likelihood [11].
@@
 This is effective even with sparse labels [15, 16].
+The EM algorithm has already been used for supervised learning tasks using specific models [8, 17, 18], where the goal has been point prediction with GMMs.
 Similarly, the EM algorithm adapts well to minimal supervision [19] and using labeled and unlabeled data in semi-supervised settings for both single and multiple modalities [20, 21].
+Our work differs by modifying using MC sampling to generate prediction intervals with any ML model and, in theory, under any distribution.
 3.2 Prediction Intervals
-Crucial for estimating uncertainty, prediction intervals in regression are often derived using methods such as Bayesian approaches [23, 24, 25], ensemble techniques [26, 27], or quantile regression [28, 29].
+Crucial for estimating uncertainty, prediction intervals in regression are often derived using methods such as Bayesian approaches [22, 23, 24], ensemble techniques [25, 26, 27], or quantile regression [28, 29].
 Additionally, these methods can be complemented with conformal prediction, a framework for calibrating any point predictor to produce prediction intervals [30, 31], making it highly relevant for enhancing reliability in applications requiring rigorous uncertainty quantification.
@@
 Managing missing values is a pivotal aspect when dealing with real-world data.
-Naive methods such as discarding instances or mean/median-imputation may be infeasible or carry the risk of changing the data distribution [35, 36] introducing bias if the missing values are not missing at random (MAR) or completely at random (MCAR).
+Naive methods such as discarding instances or mean/median-imputation may be infeasible or carry the risk of changing the data distribution [35, 36].
 The chosen technique for handling missing data should adhere to the dataset’s characteristics and mechanisms behind the missing data [37].
@@
 The features of these datasets are then treated as separate inputs to SEMF.
-In the second stage of our experiments, we artificially introduce 50% missing values in our datasets for any predictor except for the first feature of a randomly chosen row, which emulates MAR or MCAR data.
+In the second stage of our experiments, we artificially introduce 50% missing values in our datasets for any predictor except for the first feature of a randomly chosen row, which emulates missing completely at random (MCAR) data.
 In all our datasets, 70% of the data is used to train all models, 15% as a hold-out validation set to monitor SEMF’s performance, and 15% to evaluate the models.
@@
 4.2 Models
-Our baseline consists of both point and quantile regression eXtreme Gradient Boosting (XGBoost) [46], Extremely Randomized Trees (ET) [47], and neural networks [48], depicted in Table 1.
+Our baseline consists of both point and quantile regression eXtreme Gradient Boosting (XGBoost) [46], Extremely Randomized Trees (ET) [47], and neural networks [48], all summarized and depicted in Table 1.
 To ensure consistency in our experimental setup, we align the families and hyper-parameters of $p_{\phi}$ and $p_{\theta}$ with our baseline models.
-For example, in the case of XGBoost in SEMF, we use $K$ XGBoost’s $g_{\phi_{k}}(x_{k})$, one for each input $x_{k}$, $k=1,\ldots,K$, and one XGBoost for $f_{\theta}(z)$ with the same hyper-parameters.
+For example, in the case of XGBoost in SEMF, we use $K$ XGBoosts, $g_{\phi_{k}}(x_{k})$, one for each input $x_{k}$, $k=1,\ldots,K$, and one XGBoost for $f_{\theta}(z)$ with the same hyper-parameters.
 We refer to the SEMF’s adoption of these models as MultiXGBs, MultiETs, and MultiMLPs.
-When establishing prediction intervals, we set our uncertainty tolerance at 5% for both the baseline and SEMF (Eq.
+When establishing prediction intervals, we conformalize our prediction intervals according to [34] at an uncertainty tolerance of 5% for both the baseline and SEMF (Eq.
 18).
@@
 MultiXGBs XGBoost [50]
-Trees: 100, Maximum depth: 5, Early stopping steps: 10
+Trees: 100, Maximum depth: 6, Early stopping steps: 10
 Quantile XGBoost
@@
 The evaluation of point predictions employs Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and R-squared ($\text{R}^{2}$).
-For prediction intervals, the chosen metrics, following [52, 53], are the Prediction Interval Coverage Probability (PICP), Mean Prediction Interval Width (MPIW), and Normalized Mean Prediction Interval Width (NMPIW) as described in Subsection D.1.
+For prediction intervals, the chosen metrics, following [52] and [53], are the Prediction Interval Coverage Probability (PICP), and Normalized Mean Prediction Interval Width (NMPIW) as described in Subsection D.1.
 This paper also evaluates SEMF on our new metric, termed Coverage-Width Ratio (CWR),
@@
 5 Results
-We trained and tested 330 models corresponding to the three model types—MultiXGBs, MultiETs, and MultiMLPs—across 11 datasets with both complete and 50% missing data, with five seeds for each combination.
-Table 2 and Table 3 provide the mean and standard deviation for our metrics aggregated over the five seeds.
-All the results are in Appendix E.
-Below, we discuss the performances with complete data and with incomplete data.
+We trained and tested 330 models corresponding to the three model types—MultiXGBs, MultiETs, and MultiMLPs—across 11 datasets with both complete and 50% missing data, using five seeds for each combination.
+Table 2 and Table 3 present the mean and standard deviation for our metrics aggregated over the five seeds.
+Appendix E includes the results from each individual run.
+For comparison, we also present the non-conformalized prediction intervals in Appendix F, though we solely discuss the results for conformalized intervals on both complete and incomplete data.
 5.1 Complete Data
 The results of models with complete data are in Table 2.
-Overall, MultiXGBs and MultiMLPs performed generally well in producing intervals compared to their baselines as shown by positive $\Delta$CWR and $\Delta$PICP, notably, for instance, on cpu_activity, naval_propulsion_plant, and energy_efficiency.
+Overall, MultiXGBs and MultiMLPs performed generally well in producing intervals compared to their baselines, as shown by positive $\Delta$CWR and $\Delta$NMPIW while achieving similar $\Delta$PICP to the baselines.
+Notably, all models attained suitable intervals on naval_propulsion_plant and energy_efficiency.
 Interestingly, MultiMLPs also attained good relative performance improvements for point prediction, which can indicate that the chosen $\sigma_{k}$ was too low for generating performant prediction intervals.
-MultiETs attained mixed results as, despite significant improvements in $\Delta$CWR, it often failed to beat the baseline on $\Delta$PICP.
+MultiETs attained mixed results, with significant improvements in $\Delta$CWR for other datasets such as naval_propulsion_plant, but its overall performance remains less conclusive.
 Table 2: Test results for all models with complete data at 95% quantiles aggregated over five seeds.
-For each metric, the mean and standard deviation of the performance across the seeds are separated by ±.
+For each metric, the mean and standard deviation of the performance across the seeds are separated by $\pm$.
 Performance over the baseline is highlighted in bold.
@@
 Relative Absolute Relative Absolute
-Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP MPIW NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
+Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
 MultiXGBs
-space_ga -1%±6% 6%±2% -7%±6% 0.89±0.01 1.80±0.12 0.20±0.01 -9%±1% -10%±2% 0.60±0.01
-cpu_activity 25%±7% 9%±1% 12%±5% 0.89±0.01 0.37±0.01 0.07±0.00 21%±1% -1%±3% 0.98±0.00
-naval_propulsion_plant 156%±15% 8%±1% 58%±2% 0.96±0.00 0.39±0.01 $0$14%±4% -12%±2% 0.99±0.00
-miami_housing 61%±6% -7%±2% 43%±3% 0.82±0.01 0.45±0.01 0.06±0.00 4%±5% 3%±3% 0.91±0.01
-kin8nm 10%±3% 5%±2% 4%±4% 0.90±0.01 1.94±0.07 0.38±0.01 -20%±2% -22%±2% 0.63±0.01
-concrete_compressive_strength 13%±5% 26%±6% -11%±8% 0.91±0.01 1.25±0.03 0.25±0.01 -26%±7% -40%±13% 0.85±0.01
-cars 18%±9% 15%±4% 2%±9% 0.89±0.02 0.70±0.05 0.12±0.01 -3%±3% -1%±4% 0.95±0.00
-energy_efficiency 247%±99% 22%±7% 60%±20% 0.88±0.02 0.16±0.04 0.05±0.01 -15%±25% -16%±19% 1.00±0.00
-california_housing 31%±5% 0%±1% 23%±4% 0.88±0.01 1.17±0.02 $0$1%±1% -1%±1% 0.81±0.00
-airfoil_self_noise 61%±45% 4%±10% 31%±18% 0.81±0.07 0.89±0.22 0.19±0.05 -64%±36% -73%±46% 0.86±0.05
-QSAR_fish_toxicity 10%±4% 11%±4% -1%±3% 0.76±0.01 1.46±0.03 0.23±0.00 3%±4% 1%±5% 0.55±0.02
+space_ga 6%$\pm$3-1%$\pm$07%$\pm$20.94$\pm$0.01 0.26$\pm$0.01 -9%$\pm$1-10%$\pm$20.60$\pm$0.01
+cpu_activity 16%$\pm$61%$\pm$112%$\pm$50.94$\pm$0.01 0.09$\pm$0.00 21%$\pm$1-1%$\pm$20.98$\pm$0.00
+naval_propulsion_plant 172%$\pm$140%$\pm$163%$\pm$20.95$\pm$0.01 $0$14%$\pm$4-12%$\pm$20.99$\pm$0.00
+miami_housing -7%$\pm$30%$\pm$0-8%$\pm$30.95$\pm$0.00 0.13$\pm$0.00 4%$\pm$43%$\pm$30.91$\pm$0.01
+kin8nm 6%$\pm$21%$\pm$05%$\pm$20.94$\pm$0.01 0.45$\pm$0.01 -20%$\pm$1-22%$\pm$10.63$\pm$0.01
+concrete_compressive_strength 41%$\pm$12-3%$\pm$231%$\pm$70.94$\pm$0.02 0.31$\pm$0.01 -26%$\pm$6-40%$\pm$110.85$\pm$0.01
+cars 40%$\pm$18-3%$\pm$130%$\pm$80.91$\pm$0.01 0.13$\pm$0.01 -3%$\pm$3-1%$\pm$30.95$\pm$0.00
+energy_efficiency 222%$\pm$45-4%$\pm$370%$\pm$30.92$\pm$0.02 0.05$\pm$0.01 -15%$\pm$22-16%$\pm$171.00$\pm$0.00
+california_housing -1%$\pm$30%$\pm$0-2%$\pm$40.95$\pm$0.00 0.42$\pm$0.01 -1%$\pm$1-1%$\pm$10.81$\pm$0.00
+airfoil_self_noise 21%$\pm$21-1%$\pm$215%$\pm$150.97$\pm$0.02 0.36$\pm$0.06 -64%$\pm$32-73%$\pm$410.86$\pm$0.05
+QSAR_fish_toxicity 34%$\pm$8-5%$\pm$329%$\pm$50.87$\pm$0.02 0.33$\pm$0.01 3%$\pm$31%$\pm$40.55$\pm$0.02
 MultiETs
-space_ga 9%±4% -2%±1% 9%±4% 0.91±0.01 2.04±0.07 0.22±0.01 -16%±2% -18%±3% 0.54±0.02
-cpu_activity 27%±3% -8%±0% 27%±2% 0.90±0.01 0.47±0.01 $0$14%±2% -20%±2% 0.98±0.00
-naval_propulsion_plant 219%±36% -9%±0% 71%±3% 0.91±0.00 0.66±0.07 0.19±0.02 -320%±45% -409%±40% 0.96±0.01
-miami_housing 69%±4% -13%±0% 48%±1% 0.86±0.00 0.65±0.02 $0$10%±2% -20%±2% 0.90±0.00
-kin8nm 10%±2% -9%±0% 17%±2% 0.88±0.01 2.27±0.05 0.44±0.01 -35%±2% -38%±2% 0.48±0.01
-concrete_compressive_strength 9%±10% -12%±2% 19%±9% 0.81±0.02 1.18±0.11 0.24±0.02 -67%±7% -94%±10% 0.76±0.02
-cars 11%±7% 16%±2% -5%±7% 0.83±0.02 0.53±0.04 0.09±0.01 3%±4% 1%±2% 0.95±0.00
-energy_efficiency -4%±6% 0%±8% -5%±15% 0.65±0.04 0.08±0.01 0.02±0.00 3%±4% 0%±2% 1.00±0.00
-california_housing 31%±2% -7%±0% 29%±1% 0.90±0.00 1.70±0.03 0.41±0.01 -15%±1% -21%±1% 0.71±0.01
-airfoil_self_noise 18%±28% -15%±5% 25%±18% 0.83±0.05 1.22±0.31 0.26±0.07 -117%±50% -141%±57% 0.80±0.09
-QSAR_fish_toxicity 12%±2% -9%±1% 19%±2% 0.82±0.01 1.81±0.02 $0$6%±1% -11%±1% 0.53±0.01
+space_ga -6%$\pm$31%$\pm$1-7%$\pm$30.96$\pm$0.00 0.29$\pm$0.01 -15%$\pm$2-18%$\pm$30.54$\pm$0.02
+cpu_activity 11%$\pm$2-4%$\pm$013%$\pm$20.94$\pm$0.00 $0$14%$\pm$1-20%$\pm$20.98$\pm$0.00
+naval_propulsion_plant 137%$\pm$221%$\pm$057%$\pm$40.96$\pm$0.00 0.27$\pm$0.02 -316%$\pm$45-406%$\pm$430.96$\pm$0.01
+miami_housing -9%$\pm$1-1%$\pm$0-9%$\pm$10.95$\pm$0.00 $0$10%$\pm$2-20%$\pm$20.90$\pm$0.00
+kin8nm -10%$\pm$11%$\pm$1-12%$\pm$20.94$\pm$0.01 0.53$\pm$0.01 -34%$\pm$2-38%$\pm$20.48$\pm$0.01
+concrete_compressive_strength -8%$\pm$5-3%$\pm$2-6%$\pm$60.89$\pm$0.02 0.31$\pm$0.02 -67%$\pm$6-94%$\pm$80.76$\pm$0.01
+cars -25%$\pm$50%$\pm$3-34%$\pm$140.92$\pm$0.02 0.15$\pm$0.01 3%$\pm$41%$\pm$20.95$\pm$0.00
+energy_efficiency 10%$\pm$6-4%$\pm$212%$\pm$60.94$\pm$0.02 0.06$\pm$0.00 3%$\pm$21%$\pm$11.00$\pm$0.00
+california_housing 1%$\pm$2-2%$\pm$03%$\pm$20.95$\pm$0.00 0.55$\pm$0.01 -15%$\pm$1-21%$\pm$10.71$\pm$0.01
+airfoil_self_noise -16%$\pm$10-3%$\pm$2-17%$\pm$120.96$\pm$0.02 0.43$\pm$0.03 -118%$\pm$43-141%$\pm$490.80$\pm$0.08
+QSAR_fish_toxicity -6%$\pm$9-5%$\pm$2-2%$\pm$120.88$\pm$0.02 0.36$\pm$0.02 -6%$\pm$1-11%$\pm$10.53$\pm$0.01
 MultiMLPs
-space_ga 8%±1% 0%±3% 7%±3% 0.81±0.02 1.25±0.02 0.14±0.00 0%±1% -1%±1% 0.75±0.01
-cpu_activity 8%±3% -9%±4% 15%±6% 0.72±0.02 0.24±0.01 0.05±0.00 7%±2% 5%±2% 0.98±0.00
-naval_propulsion_plant 21%±20% 0%±3% 16%±15% 0.92±0.01 0.22±0.03 0.07±0.01 -45%±28% -36%±23% 1.00±0.00
-miami_housing 9%±4% 2%±3% 6%±6% 0.81±0.02 0.41±0.03 $0$3%±2% 4%±2% 0.91±0.00
-kin8nm 2%±8% 8%±6% -6%±12% 0.81±0.02 0.68±0.03 0.13±0.01 7%±3% 7%±2% 0.93±0.00
-concrete_compressive_strength 98%±15% -23%±8% 61%±5% 0.56±0.04 0.28±0.01 0.06±0.00 12%±6% 15%±4% 0.91±0.01
-cars 0%±15% 6%±13% -9%±33% 0.76±0.09 0.40±0.10 0.07±0.02 -3%±3% 0%±3% 0.95±0.00
-energy_efficiency 82%±20% 25%±26% 31%±12% 0.62±0.05 0.06±0.00 0.02±0.00 32%±4% 33%±3% 1.00±0.00
-california_housing -14%±3% 8%±1% -26%±4% 0.89±0.01 1.31±0.02 0.31±0.01 7%±0% 9%±1% 0.82±0.00
-airfoil_self_noise 74%±46% 0%±7% 39%±17% 0.75±0.02 0.35±0.01 0.08±0.00 18%±4% 16%±5% 0.97±0.00
-QSAR_fish_toxicity -6%±4% 11%±4% -19%±5% 0.76±0.04 1.49±0.06 0.23±0.01 4%±5% 7%±5% 0.55±0.04
+space_ga 6%$\pm$3-1%$\pm$16%$\pm$20.95$\pm$0.01 0.23$\pm$0.00 0%$\pm$1-1%$\pm$10.75$\pm$0.01
+cpu_activity -7%$\pm$60%$\pm$1-9%$\pm$70.95$\pm$0.00 0.10$\pm$0.00 7%$\pm$25%$\pm$20.98$\pm$0.00
+naval_propulsion_plant 4%$\pm$161%$\pm$00%$\pm$140.96$\pm$0.00 0.08$\pm$0.01 -45%$\pm$25-36%$\pm$201.00$\pm$0.00
+miami_housing -38%$\pm$30%$\pm$1-61%$\pm$90.95$\pm$0.00 0.15$\pm$0.01 -3%$\pm$24%$\pm$20.91$\pm$0.00
+kin8nm 8%$\pm$50%$\pm$18%$\pm$50.95$\pm$0.01 0.20$\pm$0.01 7%$\pm$27%$\pm$20.93$\pm$0.00
+concrete_compressive_strength 11%$\pm$6-2%$\pm$211%$\pm$60.94$\pm$0.02 0.29$\pm$0.03 12%$\pm$515%$\pm$40.91$\pm$0.01
+cars 3%$\pm$12-3%$\pm$34%$\pm$110.92$\pm$0.02 0.13$\pm$0.01 -3%$\pm$30%$\pm$30.95$\pm$0.00
+energy_efficiency 53%$\pm$210%$\pm$434%$\pm$100.96$\pm$0.02 0.05$\pm$0.00 32%$\pm$333%$\pm$31.00$\pm$0.00
+california_housing -12%$\pm$30%$\pm$1-14%$\pm$40.95$\pm$0.00 0.43$\pm$0.01 7%$\pm$09%$\pm$10.82$\pm$0.00
+airfoil_self_noise 68%$\pm$28-2%$\pm$140%$\pm$110.97$\pm$0.01 0.18$\pm$0.01 18%$\pm$416%$\pm$50.97$\pm$0.00
+QSAR_fish_toxicity 5%$\pm$101%$\pm$13%$\pm$100.89$\pm$0.03 0.35$\pm$0.04 4%$\pm$57%$\pm$40.55$\pm$0.04
 5.2 Missing Data
@@
 Note that relative metrics for SEMF are compared with the best result from any baseline imputer on that metric, regardless of how the imputer performed on the other metrics.
-Overall, the results are less consistent than with complete data.
-MultiXGBs maintained strong performance on some datasets, such as naval_propulsion_plant, but declined on others, like cpu_activity.
-MultiETs exhibited narrow predictions, improved absolute PICP in some cases, such as energy_efficiency, and improved relative point predictions.
-Conversely, MultiMLPs, while offering narrow prediction intervals, had higher PICP, indicating decreased predictive power and increased model uncertainty.
+Overall, the results are worse and less consistent than with complete data.
+MultiXGBs maintained good performance on some datasets, such as naval_propulsion_plant and energy_efficiency, but declined on others, like cpu_activity.
+MultiETs continued to exhibit mixed results, with only naval_propulsion_plant offering marginally better performance over the baseline.
+Similarly, MultiMLPs worked well on only one dataset, namely concrete_compressive_strength, while the other datasets had increased model uncertainty.
 Table 3: Test results for all models with missing data at 95% quantiles aggregated over five seeds.
-For each metric, the mean and standard deviation of the performance across the seeds are separated by ±.
+For each metric, the mean and standard deviation of the performance across the seeds are separated by $\pm$.
 Performance over the baseline is highlighted in bold.
@@
 Relative Absolute Relative Absolute
-Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP MPIW NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
+Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
 MultiXGBs
-space_ga -5%±6% 4%±2% -12%±7% 0.86±0.02 2.05±0.16 0.22±0.02 -7%±5% -9%±3% 0.45±0.07
-cpu_activity -5%±21% 12%±9% -37%±52% 0.89±0.07 0.76±0.24 0.14±0.05 -1%±24% -17%±14% 0.83±0.08
-naval_propulsion_plant 105%±33% -16%±7% 56%±6% 0.70±0.07 0.62±0.06 0.18±0.02 -14%±22% -29%±39% 0.77±0.06
-miami_housing 16%±14% -5%±6% 10%±18% 0.82±0.07 0.77±0.15 0.10±0.02 -14%±12% -18%±11% 0.72±0.11
-kin8nm -5%±2% 3%±2% -14%±5% 0.88±0.01 2.46±0.07 0.48±0.01 -3%±3% -4%±5% 0.40±0.02
-concrete_compressive_strength 2%±3% 8%±4% -10%±4% 0.77±0.04 1.47±0.03 0.30±0.01 -16%±14% -20%±16% 0.58±0.09
-cars -15%±15% 26%±10% -63%±29% 0.85±0.04 1.58±0.18 0.27±0.03 -30%±25% -35%±27% 0.66±0.17
-energy_efficiency 10%±29% 11%±6% -24%±48% 0.92±0.03 0.77±0.19 0.22±0.05 -186%±296% -153%±210% 0.91±0.09
-california_housing 26%±8% -2%±3% 21%±7% 0.88±0.03 1.67±0.12 0.40±0.03 -4%±4% -5%±5% 0.69±0.05
-airfoil_self_noise -2%±21% -3%±4% -4%±19% 0.73±0.05 1.59±0.33 0.34±0.07 -37%±34% -43%±40% 0.41±0.13
-QSAR_fish_toxicity -2%±3% 0%±5% -6%±8% 0.73±0.03 1.70±0.13 0.26±0.02 -4%±5% -2%±4% 0.36±0.05
+space_ga 0%$\pm$3-2%$\pm$12%$\pm$30.94$\pm$0.01 0.32$\pm$0.01 -7%$\pm$4-9%$\pm$30.45$\pm$0.06
+cpu_activity -14%$\pm$90%$\pm$2-20%$\pm$140.94$\pm$0.03 0.18$\pm$0.02 -1%$\pm$21-17%$\pm$130.83$\pm$0.07
+naval_propulsion_plant 18%$\pm$23-3%$\pm$212%$\pm$180.93$\pm$0.04 0.59$\pm$0.17 -14%$\pm$20-29%$\pm$350.77$\pm$0.05
+miami_housing -18%$\pm$13-1%$\pm$1-25%$\pm$220.94$\pm$0.01 0.20$\pm$0.02 -14%$\pm$11-18%$\pm$100.72$\pm$0.10
+kin8nm -9%$\pm$31%$\pm$1-13%$\pm$50.96$\pm$0.01 0.63$\pm$0.03 -3%$\pm$3-4%$\pm$40.40$\pm$0.02
+concrete_compressive_strength 1%$\pm$9-2%$\pm$21%$\pm$100.95$\pm$0.01 0.57$\pm$0.02 -16%$\pm$12-20%$\pm$140.58$\pm$0.08
+cars 9%$\pm$27-7%$\pm$46%$\pm$230.90$\pm$0.04 0.34$\pm$0.02 -30%$\pm$22-35%$\pm$240.66$\pm$0.15
+energy_efficiency 26%$\pm$24-4%$\pm$418%$\pm$180.95$\pm$0.04 0.27$\pm$0.05 -186%$\pm$265-153%$\pm$1880.91$\pm$0.08
+california_housing -6%$\pm$6-1%$\pm$1-8%$\pm$80.95$\pm$0.02 0.59$\pm$0.03 -4%$\pm$4-5%$\pm$50.69$\pm$0.05
+airfoil_self_noise -24%$\pm$6-1%$\pm$2-32%$\pm$120.95$\pm$0.03 0.76$\pm$0.07 -37%$\pm$31-43%$\pm$360.41$\pm$0.12
+QSAR_fish_toxicity 0%$\pm$8-4%$\pm$50%$\pm$110.91$\pm$0.05 0.50$\pm$0.11 -4%$\pm$4-2%$\pm$30.36$\pm$0.04
 MultiETs
-space_ga 4%±4% -5%±1% 7%±4% 0.88±0.02 2.23±0.10 0.24±0.01 -12%±4% -14%±5% 0.40±0.06
-cpu_activity -16%±12% -5%±2% -20%±21% 0.93±0.01 1.07±0.20 0.20±0.04 -30%±26% -54%±22% 0.79±0.11
-naval_propulsion_plant 58%±5% -17%±5% 46%±3% 0.82±0.05 1.38±0.10 0.41±0.03 -36%±23% -93%±53% 0.71±0.06
-miami_housing 20%±16% -10%±5% 21%±14% 0.87±0.05 1.15±0.26 0.14±0.03 -27%±12% -37%±12% 0.67±0.12
-kin8nm 5%±3% -8%±2% 12%±4% 0.87±0.02 2.55±0.10 0.50±0.02 -11%±3% -13%±3% 0.32±0.03
-concrete_compressive_strength 1%±6% -14%±4% 13%±2% 0.80±0.02 1.83±0.06 0.37±0.01 -35%±23% -50%±28% 0.49±0.07
-cars -23%±7% -8%±6% -35%±16% 0.83±0.04 1.41±0.15 0.24±0.02 -33%±20% -31%±20% 0.66±0.10
-energy_efficiency -15%±18% 0%±10% -36%±47% 0.85±0.06 0.51±0.05 0.15±0.01 -73%±82% -84%±72% 0.94±0.04
-california_housing 21%±7% -7%±2% 22%±5% 0.91±0.02 2.19±0.13 0.52±0.03 -14%±3% -17%±4% 0.61±0.06
-airfoil_self_noise -7%±20% -20%±2% 8%±18% 0.76±0.03 1.96±0.39 0.42±0.08 -65%±60% -93%±78% 0.33±0.12
-QSAR_fish_toxicity 10%±8% -9%±3% 16%±4% 0.82±0.04 2.12±0.20 0.33±0.03 -8%±9% -10%±10% 0.39±0.10
+space_ga -11%$\pm$20%$\pm$2-14%$\pm$50.95$\pm$0.01 0.34$\pm$0.02 -12%$\pm$4-13%$\pm$50.40$\pm$0.06
+cpu_activity -15%$\pm$8-3%$\pm$2-18%$\pm$130.94$\pm$0.02 0.20$\pm$0.02 -21%$\pm$18-44%$\pm$140.82$\pm$0.07
+naval_propulsion_plant 4%$\pm$16-4%$\pm$34%$\pm$180.95$\pm$0.03 0.73$\pm$0.13 -37%$\pm$21-95%$\pm$480.71$\pm$0.06
+miami_housing -33%$\pm$100%$\pm$3-55%$\pm$240.95$\pm$0.01 0.27$\pm$0.02 -25%$\pm$10-39%$\pm$140.68$\pm$0.12
+kin8nm -14%$\pm$11%$\pm$1-17%$\pm$20.95$\pm$0.01 0.66$\pm$0.02 -11%$\pm$2-13%$\pm$30.32$\pm$0.03
+concrete_compressive_strength -20%$\pm$3-2%$\pm$4-27%$\pm$90.95$\pm$0.02 0.61$\pm$0.02 -37%$\pm$18-51%$\pm$220.48$\pm$0.06
+cars -32%$\pm$12-4%$\pm$4-53%$\pm$320.93$\pm$0.02 0.38$\pm$0.05 -38%$\pm$15-32%$\pm$140.62$\pm$0.16
+energy_efficiency -30%$\pm$28-4%$\pm$2-82%$\pm$1200.96$\pm$0.03 0.32$\pm$0.22 -69%$\pm$60-75%$\pm$430.94$\pm$0.03
+california_housing -6%$\pm$4-2%$\pm$1-4%$\pm$40.96$\pm$0.01 0.70$\pm$0.03 -14%$\pm$3-17%$\pm$30.61$\pm$0.05
+airfoil_self_noise -31%$\pm$4-3%$\pm$1-43%$\pm$90.94$\pm$0.02 0.75$\pm$0.05 -62%$\pm$51-89%$\pm$660.35$\pm$0.10
+QSAR_fish_toxicity -7%$\pm$7-3%$\pm$3-7%$\pm$60.92$\pm$0.04 0.50$\pm$0.08 -9%$\pm$9-10%$\pm$90.39$\pm$0.09
 MultiMLPs
-space_ga -11%±6% -15%±2% 1%±7% 0.68±0.04 1.44±0.09 0.16±0.01 -40%±26% -34%±17% 0.21±0.25
-cpu_activity -50%±18% 22%±16% -193%±132% 0.86±0.11 0.97±0.42 0.18±0.08 -15%±21% -23%±19% 0.75±0.14
-naval_propulsion_plant -52%±21% -14%±27% -102%±62% 0.63±0.20 1.17±0.28 0.34±0.08 -108%±74% -183%±218% -0.03±0.53
-miami_housing -33%±17% 13%±11% -92%±78% 0.85±0.06 0.99±0.32 0.12±0.04 -25%±9% -28%±12% 0.67±0.11
-kin8nm -11%±6% -5%±2% -8%±6% 0.77±0.01 1.76±0.09 0.35±0.02 -16%±9% -15%±10% 0.37±0.04
-concrete_compressive_strength -15%±14% -13%±31% -10%±54% 0.60±0.25 1.11±0.59 0.22±0.12 -12%±6% -12%±7% 0.54±0.06
-cars -51%±5% 43%±8% -211%±38% 0.93±0.02 1.74±0.22 0.29±0.04 -28%±14% -26%±15% 0.66±0.11
-energy_efficiency -13%±20% 9%±17% -30%±28% 0.66±0.15 0.33±0.15 0.09±0.04 1%±10% -1%±18% 0.95±0.03
-california_housing -17%±7% 4%±4% -28%±16% 0.88±0.03 1.84±0.17 0.44±0.04 -5%±3% -3%±4% 0.61±0.06
-airfoil_self_noise 2%±18% -2%±6% -1%±23% 0.75±0.08 1.15±0.41 0.25±0.09 -15%±9% -9%±9% 0.55±0.20
-QSAR_fish_toxicity -17%±5% 6%±7% -30%±12% 0.76±0.04 1.82±0.18 0.28±0.03 -13%±7% -14%±9% 0.34±0.07
+space_ga -25%$\pm$11-3%$\pm$2-34%$\pm$250.94$\pm$0.01 0.38$\pm$0.06 -40%$\pm$23-34%$\pm$150.21$\pm$0.23
+cpu_activity -40%$\pm$90%$\pm$2-73%$\pm$240.95$\pm$0.01 0.25$\pm$0.02 -15%$\pm$19-23%$\pm$170.75$\pm$0.12
+naval_propulsion_plant -43%$\pm$11-1%$\pm$2-89%$\pm$440.95$\pm$0.02 1.33$\pm$0.19 -108%$\pm$66-183%$\pm$195-0.03$\pm$0.47
+miami_housing -36%$\pm$7-1%$\pm$1-57%$\pm$200.94$\pm$0.02 0.22$\pm$0.02 -25%$\pm$8-28%$\pm$110.67$\pm$0.10
+kin8nm -19%$\pm$30%$\pm$0-24%$\pm$50.96$\pm$0.01 0.70$\pm$0.02 -16%$\pm$8-15%$\pm$90.37$\pm$0.04
+concrete_compressive_strength 5%$\pm$10-3%$\pm$25%$\pm$100.94$\pm$0.02 0.54$\pm$0.04 -12%$\pm$5-12%$\pm$60.54$\pm$0.05
+cars -22%$\pm$17-1%$\pm$4-38%$\pm$300.95$\pm$0.03 0.35$\pm$0.06 -28%$\pm$12-26%$\pm$130.66$\pm$0.10
+energy_efficiency -7%$\pm$11-4%$\pm$3-11%$\pm$140.95$\pm$0.03 0.31$\pm$0.08 1%$\pm$9-1%$\pm$160.95$\pm$0.03
+california_housing -24%$\pm$50%$\pm$1-34%$\pm$100.96$\pm$0.01 0.68$\pm$0.02 -5%$\pm$2-3%$\pm$30.61$\pm$0.06
+airfoil_self_noise -8%$\pm$6-4%$\pm$3-7%$\pm$80.93$\pm$0.04 0.61$\pm$0.07 -15%$\pm$8-9%$\pm$80.55$\pm$0.18
+QSAR_fish_toxicity -11%$\pm$5-2%$\pm$3-14%$\pm$80.92$\pm$0.04 0.53$\pm$0.11 -13%$\pm$7-14%$\pm$80.34$\pm$0.06
 5.3 Discussion
-The results indicate that specific models, particularly MultiXGBs, perform strongly with complete data.
-However, their robustness diminishes with increasing amounts of missing data.
-Further, we observe that MultiETs can produce narrow prediction intervals at the expense of reduced coverage.
-In contrast, despite the experiment design prioritizing interval estimation over point accuracy, MultiMLPs also deliver good point predictions.
-Precise choices, such as processing the columns separately, using the same hyper-parameters for both the complete and incomplete data, and fixing the same number of latent nodes per input ($m_{k}$), might have further constrained the models’ ability to tailor their performance to varying dataset complexities.
-The stability of SEMF across different seeds, as indicated by the absolute PICP and NMPIW, contrasts with the significant variability observed in baseline models, suggesting that SEMF offers a more reliable performance framework.
-The method’s sampling operation, akin to cross-validation, helps combat overfitting, ensuring robust results despite variations in the data.
-Nonetheless, most models do not achieve the desired 95% prediction intervals.
-This shortfall could be addressed by integrating conformal prediction techniques with the intervals generated by SEMF, enhancing the reliability and coverage of prediction intervals across different datasets and missing data scenarios [34].
+Our results indicate that SEMF, when combined with XGBoost and MLPs, performs strongly on datasets with complete data.
+Both models produce better prediction intervals than traditional quantile regression methods in the complete case.
+MultiMLPs also deliver good point predictions for some datasets despite the experimental design prioritizing interval estimation over point accuracy.
+This could indicate either the effectiveness of the chosen $\sigma_{k}$, which (indirectly) led to narrower prediction intervals or underfitting of the MLPs.
+The truth may lie somewhere in between; SEMF’s sampling operation, akin to cross-validation, partially helps combat overfitting, ensuring robust results despite variations in the data.
+In our preliminary experiments, we observed that increasing the depth of the baseline MLPs did not help and eventually led to overfitting.
+ETs exhibit more mixed performance, highly dependent on the dataset’s characteristics.
+One possible explanation behind the larger variations in predictive power compared to other models may be ETs’ reliance on randomized splits, which introduces significant variability in the prediction process.
+Consequently, ETs may not benefit from the iterative sampling and refinement of predictions inherent in SEMF.
+The robustness of SEMF diminishes when applied to datasets with missing data.
+Despite not offering improvements over the baseline in the presence of missing data, we have presented these results to transparently illustrate our framework’s current capabilities of handling missing inputs.
+From a theoretical standpoint, we believe there is value in further investigating how a single loss function that leverages EM’s missing data capabilities can be effectively applied.
+Given that SEMF performs well with complete data, a practical alternative might involve using the best imputation method for the data at hand before applying SEMF, effectively treating the data as complete.
+We expect our ablation study to perform well, given the complete data results.
+An important observation is the stability of SEMF across different random seeds, as indicated by consistent PICP and NMPIW metrics (full results in Appendix E), contrasting the significant variability observed in the baseline models.
+This suggests that SEMF offers a more reliable performance framework.
+Additionally, it is worth noting that our experiments did not precisely tune for conformalized prediction intervals but used the non-conformalized (raw) quantiles from SEMF.
+The non-conformalized intervals offer less coverage but produce better CWR and NMPIW than the baseline (Appendix F).
+Certain experimental choices—such as processing the columns separately, using the same hyperparameters for both complete and incomplete data, and fixing the number of latent nodes per input ($m_{k}$)—may have further constrained the models’ ability to tailor their performance to the varying complexities of different datasets.
 6 Conclusion
@@
 Due to SEMF’s iterative simulation technique, training and inference can be done with complete and incomplete data.
-A comprehensive set of 330 experimental runs on 11 datasets with three different model types showed that SEMF is superior to quantile regression, particularly for models such as XGBoost, which intrinsically lacks latent representations.
-This research underscores the potential of SEMF in various application domains and opens new avenues for further exploration for supervised latent representation learning and uncertainty estimation.
+A comprehensive set of 330 experimental runs on 11 datasets with three different model types showed that SEMF, in the case of complete data, outperforms quantile regression, particularly on complete datasets and when using XGBoost, which intrinsically lacks latent representations.
+The results of the missing data are less positive and require further investigation.
+This research underscores SEMF’s potential in various application domains and opens new avenues for further exploration of supervised latent representation learning and uncertainty estimation.
 7 Limitations & Future Work
 The primary limitation of this study was its reliance on the normality assumption, which may not fully capture the potential of SEMF across diverse data distributions.
-Although in Appendix F we demonstrate that the framework can learn non-normal patterns, further investigation and exploration of SEMF under other distributions, such as uniform and log-normal, are needed.
+Although in Appendix G we demonstrate that the framework can learn non-normal patterns, further investigation and exploration of SEMF under other distributions, such as uniform, log-normal, and generalized extreme value distributions, are needed.
 The computational complexity of the approach presents another significant challenge, as the current implementation can be optimized for large-scale applications.
@@
 Evaluating CWR under various distributional assumptions would provide a more comprehensive assessment of its implications.
+Additionally, SEMF has only been evaluated on MCAR data and does not address missing at random cases (MAR), which require further investigation for real-world applicability.
 Finally, applying the same hyper-parameters across all datasets without specific tuning for incomplete data likely contributes to the observed decline in accuracy and robustness.
@@
 Further, the number of epochs in the case of MultiMLPs is set as 1000, except for energy_efficiency and QSAR_fish_toxicity, where this is changed to 5000.
-Any model-specific hyperparameter we did not specify in this paper remains at the implementation’s default value (e.g., the number of leaves in XGBoost [50]).
-Along with the supplementary code, we provide two additional CSV files: one for the optimal hyperparameters of SEMF models and the other for the results and hyperparameters of all 330 runs.
+Any model-specific hyperparameter we did not specify in this paper remains at the implementation’s default value (e.g., the number of leaves in XGBoost from [50]).
+Along with the supplementary code, we provide three additional CSV files: one for the results and hyperparameters of all 330 runs and the other two for the optimal hyperparameters of SEMF models, both raw (directly from SEMF) and conformalized.
 Additional conditions are applied only for experiments with missing data.
@@
 In instances where no configuration meets the initial improvement criteria for both metrics, we relax the requirement for positive $\Delta$PICP to accept values greater than -5% and subsequently -10%, allowing us to consider configurations where SEMF significantly improves CWR, even if the PICP improvement is less marked but remains within an acceptable range for drawing comparisons.
-Appendix E Full results
+Appendix E Full conformalized results
 Table 6: Test results for MultiXGBs with complete data at 95% quantiles for seeds 0, 10, 20, 30, 40, with rows ordered by seed (ascending). Performance over the baseline is highlighted in bold.
@@
 Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP MPIW NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
-space_ga -1% 6% -7% 0.88 1.85 0.20 -10% -13% 0.60
-space_ga -4% 5% -9% 0.89 1.85 0.20 -8% -10% 0.59
-space_ga 1% 9% -8% 0.90 1.76 0.19 -10% -10% 0.62
-space_ga -7% 5% -12% 0.90 1.92 0.21 -8% -10% 0.59
-space_ga 8% 3% 4% 0.87 1.61 0.17 -9% -8% 0.60
-cpu_activity 27% 7% 16% 0.88 0.37 0.07 21% 0% 0.98
-cpu_activity 33% 11% 17% 0.89 0.37 0.07 23% -2% 0.98
-cpu_activity 30% 10% 15% 0.88 0.36 0.07 20% -5% 0.98
-cpu_activity 16% 9% 7% 0.89 0.38 0.07 20% -3% 0.98
-cpu_activity 18% 9% 8% 0.90 0.38 0.07 22% 2% 0.98
-naval_propulsion_plant 176% 8% 61% 0.96 0.39 0.12 -20% -14% 0.99
-naval_propulsion_plant 151% 7% 58% 0.96 0.40 0.12 -9% -15% 0.99
-naval_propulsion_plant 136% 8% 54% 0.96 0.40 0.12 -15% -12% 0.99
-naval_propulsion_plant 156% 7% 58% 0.95 0.38 0.11 -12% -9% 0.99
-naval_propulsion_plant 161% 9% 58% 0.96 0.38 0.11 -15% -11% 0.99
-miami_housing 50% -5% 37% 0.82 0.46 0.06 3% -1% 0.91
-miami_housing 65% -7% 44% 0.83 0.45 0.06 4% 1% 0.91
-miami_housing 62% -10% 44% 0.81 0.44 0.06 7% 7% 0.90
-miami_housing 62% -7% 43% 0.82 0.44 0.05 9% 5% 0.91
-miami_housing 65% -9% 45% 0.81 0.44 0.05 -4% 0% 0.90
-kin8nm 12% 4% 8% 0.89 1.88 0.37 -17% -20% 0.64
-kin8nm 11% 5% 6% 0.89 1.91 0.37 -21% -22% 0.63
-kin8nm 11% 5% 6% 0.90 1.92 0.38 -19% -21% 0.64
-kin8nm 5% 7% -3% 0.91 2.06 0.40 -20% -23% 0.62
-kin8nm 8% 3% 5% 0.88 1.91 0.37 -21% -24% 0.62
-concrete_compressive_strength 7% 34% -26% 0.91 1.26 0.26 -34% -52% 0.86
-concrete_compressive_strength 13% 22% -7% 0.89 1.25 0.25 -29% -52% 0.84
-concrete_compressive_strength 13% 18% -5% 0.91 1.20 0.24 -21% -34% 0.86
-concrete_compressive_strength 20% 31% -9% 0.93 1.27 0.26 -16% -22% 0.86
-concrete_compressive_strength 14% 26% -10% 0.90 1.25 0.25 -28% -39% 0.83
-cars 35% 14% 16% 0.89 0.67 0.11 -5% -1% 0.95
-cars 16% 10% 5% 0.85 0.67 0.11 -4% -4% 0.95
-cars 13% 20% -6% 0.91 0.78 0.13 -2% 0% 0.95
-cars 12% 16% -4% 0.88 0.68 0.11 2% 5% 0.96
-cars 16% 18% -2% 0.91 0.71 0.12 -7% -4% 0.95
-energy_efficiency 255% 24% 64% 0.88 0.14 0.04 -50% -29% 1.00
-energy_efficiency 320% 19% 71% 0.91 0.14 0.04 -3% -13% 1.00
-energy_efficiency 257% 17% 67% 0.90 0.16 0.05 -21% -31% 1.00
-energy_efficiency 322% 16% 73% 0.85 0.12 0.04 17% 16% 1.00
-energy_efficiency 78% 34% 26% 0.89 0.22 0.06 -17% -22% 1.00
-california_housing 24% 1% 18% 0.88 1.19 0.28 -2% -3% 0.81
-california_housing 28% 1% 21% 0.88 1.16 0.28 0% -1% 0.82
-california_housing 33% 0% 25% 0.89 1.19 0.28 0% 0% 0.81
-california_housing 30% 1% 22% 0.87 1.16 0.28 -1% -1% 0.82
-california_housing 38% -2% 29% 0.87 1.15 0.27 -1% -2% 0.81
-airfoil_self_noise 68% 6% 37% 0.83 0.80 0.17 -41% -43% 0.89
-airfoil_self_noise 134% 5% 55% 0.81 0.59 0.13 -18% -18% 0.93
-airfoil_self_noise 13% -3% 15% 0.72 1.10 0.23 -109% -138% 0.78
-airfoil_self_noise 38% 20% 13% 0.91 1.12 0.24 -73% -76% 0.85
-airfoil_self_noise 51% -6% 38% 0.76 0.84 0.18 -82% -93% 0.86
-QSAR_fish_toxicity 11% 17% -5% 0.77 1.48 0.23 3% 3% 0.53
-QSAR_fish_toxicity 11% 9% 3% 0.75 1.42 0.22 8% 7% 0.58
-QSAR_fish_toxicity 4% 8% -5% 0.77 1.48 0.23 4% 3% 0.57
-QSAR_fish_toxicity 10% 9% 1% 0.77 1.45 0.23 -2% -1% 0.54
-QSAR_fish_toxicity 13% 15% -1% 0.74 1.48 0.23 0% -6% 0.52
+space_ga 2% -2% 4% 0.95 2.40 0.26 -10% -13% 0.60
+space_ga 10% -1% 10% 0.95 2.34 0.25 -10% -10% 0.62
+space_ga 8% -2% 9% 0.93 2.24 0.24 -9% -8% 0.60
+space_ga 6% -2% 7% 0.94 2.38 0.26 -8% -10% 0.59
+space_ga 7% -1% 7% 0.95 2.55 0.28 -8% -10% 0.59
+cpu_activity 7% 1% 5% 0.95 0.49 0.09 22% 2% 0.98
+cpu_activity 20% 1% 17% 0.93 0.45 0.09 20% -5% 0.98
+cpu_activity 20% 0% 17% 0.93 0.45 0.09 21% 0% 0.98
+cpu_activity 21% 1% 16% 0.94 0.45 0.09 23% -2% 0.98
+cpu_activity 10% 2% 7% 0.95 0.47 0.09 20% -3% 0.98
+naval_propulsion_plant 163% -1% 62% 0.95 0.36 0.11 -15% -12% 0.99
+naval_propulsion_plant 190% 0% 66% 0.95 0.37 0.11 -20% -14% 0.99
+naval_propulsion_plant 151% 0% 60% 0.96 0.40 0.12 -12% -9% 0.99
+naval_propulsion_plant 185% -1% 65% 0.94 0.35 0.10 -9% -15% 0.99
+naval_propulsion_plant 169% -1% 63% 0.96 0.37 0.11 -15% -11% 0.99
+miami_housing -2% 0% -2% 0.95 0.99 0.12 4% 1% 0.91
+miami_housing -10% 0% -10% 0.95 1.03 0.13 7% 7% 0.90
+miami_housing -6% 1% -8% 0.95 0.99 0.12 9% 5% 0.91
+miami_housing -7% 0% -8% 0.95 0.99 0.12 3% -1% 0.91
+miami_housing -10% 0% -12% 0.95 1.05 0.13 -4% 0% 0.90
+kin8nm 6% 0% 6% 0.94 2.27 0.45 -21% -24% 0.62
+kin8nm 4% 1% 3% 0.95 2.39 0.47 -19% -21% 0.64
+kin8nm 9% 1% 8% 0.94 2.20 0.43 -17% -20% 0.64
+kin8nm 9% 1% 7% 0.95 2.36 0.46 -21% -22% 0.63
+kin8nm 3% 1% 2% 0.94 2.33 0.46 -20% -23% 0.62
+concrete_compressive_strength 40% -3% 31% 0.93 1.42 0.29 -16% -22% 0.86
+concrete_compressive_strength 47% -6% 36% 0.91 1.48 0.30 -29% -52% 0.84
+concrete_compressive_strength 59% -2% 39% 0.96 1.60 0.32 -21% -34% 0.86
+concrete_compressive_strength 35% -3% 28% 0.95 1.48 0.30 -28% -39% 0.83
+concrete_compressive_strength 24% 0% 19% 0.97 1.59 0.32 -34% -52% 0.86
+cars 56% -4% 38% 0.91 0.70 0.12 -7% -4% 0.95
+cars 36% -5% 30% 0.90 0.86 0.14 -4% -4% 0.95
+cars 30% -3% 25% 0.91 0.76 0.13 2% 5% 0.96
+cars 65% -1% 40% 0.93 0.75 0.13 -5% -1% 0.95
+cars 16% -4% 17% 0.91 0.87 0.15 -2% 0% 0.95
+energy_efficiency 165% -5% 64% 0.92 0.20 0.06 -50% -29% 1.00
+energy_efficiency 288% 0% 74% 0.91 0.14 0.04 -3% -13% 1.00
+energy_efficiency 217% -1% 69% 0.96 0.18 0.05 -21% -31% 1.00
+energy_efficiency 253% -3% 73% 0.92 0.16 0.04 17% 16% 1.00
+energy_efficiency 185% -10% 68% 0.91 0.23 0.07 -17% -22% 1.00
+california_housing -5% 0% -6% 0.95 1.82 0.43 -2% -3% 0.81
+california_housing 4% 0% 4% 0.95 1.72 0.41 -1% -2% 0.81
+california_housing -4% 1% -5% 0.95 1.76 0.42 0% -1% 0.82
+california_housing 1% 0% 1% 0.95 1.74 0.41 0% 0% 0.81
+california_housing -3% 1% -4% 0.95 1.79 0.42 -1% -1% 0.82
+airfoil_self_noise 15% -4% 16% 0.95 1.73 0.37 -82% -93% 0.86
+airfoil_self_noise 45% 1% 30% 0.97 1.42 0.30 -18% -18% 0.93
+airfoil_self_noise 5% 1% 4% 0.98 1.86 0.40 -73% -76% 0.85
+airfoil_self_noise -6% 0% -7% 0.98 2.15 0.46 -109% -138% 0.78
+airfoil_self_noise 46% -2% 33% 0.95 1.37 0.29 -41% -43% 0.89
+QSAR_fish_toxicity 21% -2% 19% 0.88 2.12 0.33 3% 3% 0.53
+QSAR_fish_toxicity 39% -9% 35% 0.85 1.97 0.31 8% 7% 0.58
+QSAR_fish_toxicity 43% -3% 32% 0.89 2.21 0.34 4% 3% 0.57
+QSAR_fish_toxicity 38% -3% 30% 0.89 2.11 0.33 -2% -1% 0.54
+QSAR_fish_toxicity 29% -6% 28% 0.85 2.11 0.33 0% -6% 0.52
 Table 7: Test results for MultiETs with complete data at 95% quantiles for seeds 0, 10, 20, 30, 40, with rows ordered by seed (ascending). Performance over the baseline is highlighted in bold.
@@
 Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP MPIW NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
-space_ga 3% 0% 3% 0.92 2.15 0.23 -19% -24% 0.51
-space_ga 8% -1% 9% 0.91 2.04 0.22 -15% -18% 0.54
-space_ga 10% -1% 10% 0.91 2.04 0.22 -16% -17% 0.54
-space_ga 11% -2% 11% 0.91 2.01 0.22 -14% -17% 0.55
-space_ga 12% -3% 13% 0.90 1.96 0.21 -14% -15% 0.55
-cpu_activity 24% -8% 26% 0.91 0.48 0.09 -15% -22% 0.98
-cpu_activity 24% -8% 25% 0.91 0.49 0.09 -17% -22% 0.97
-cpu_activity 30% -9% 30% 0.90 0.45 0.09 -13% -18% 0.98
-cpu_activity 29% -8% 28% 0.90 0.47 0.09 -13% -18% 0.98
-cpu_activity 27% -8% 27% 0.91 0.48 0.09 -14% -20% 0.98
-naval_propulsion_plant 265% -9% 75% 0.91 0.58 0.17 -268% -371% 0.97
-naval_propulsion_plant 205% -9% 70% 0.90 0.69 0.20 -332% -419% 0.96
-naval_propulsion_plant 178% -9% 67% 0.91 0.74 0.22 -382% -464% 0.95
-naval_propulsion_plant 198% -9% 69% 0.91 0.69 0.20 -336% -422% 0.96
-naval_propulsion_plant 248% -10% 74% 0.90 0.58 0.17 -283% -368% 0.97
-miami_housing 70% -12% 48% 0.86 0.66 0.08 -12% -22% 0.90
-miami_housing 63% -12% 46% 0.86 0.68 0.08 -9% -21% 0.90
-miami_housing 68% -13% 48% 0.86 0.65 0.08 -9% -20% 0.90
-miami_housing 71% -13% 49% 0.86 0.63 0.08 -7% -18% 0.90
-miami_housing 72% -12% 49% 0.86 0.65 0.08 -12% -21% 0.89
-kin8nm 12% -9% 19% 0.88 2.20 0.43 -32% -36% 0.50
-kin8nm 9% -9% 17% 0.88 2.28 0.45 -36% -40% 0.46
-kin8nm 11% -10% 18% 0.88 2.24 0.44 -34% -37% 0.49
-kin8nm 9% -9% 16% 0.89 2.29 0.45 -34% -36% 0.49
-kin8nm 7% -9% 15% 0.88 2.33 0.46 -37% -40% 0.47
-concrete_compressive_strength 25% -15% 32% 0.79 1.02 0.21 -59% -85% 0.78
-concrete_compressive_strength 9% -12% 19% 0.81 1.19 0.24 -71% -96% 0.76
-concrete_compressive_strength 4% -11% 14% 0.83 1.22 0.25 -62% -92% 0.77
-concrete_compressive_strength -2% -9% 8% 0.83 1.31 0.27 -77% -110% 0.74
-concrete_compressive_strength 11% -12% 20% 0.81 1.14 0.23 -64% -88% 0.77
-cars 3% 15% -11% 0.83 0.54 0.09 3% 2% 0.95
-cars 16% 18% -2% 0.85 0.52 0.09 3% 0% 0.95
-cars 21% 13% 7% 0.80 0.47 0.08 4% -1% 0.95
-cars 6% 17% -10% 0.84 0.56 0.10 -2% 0% 0.94
-cars 10% 17% -7% 0.85 0.54 0.09 8% 5% 0.95
-energy_efficiency -3% -3% 0% 0.63 0.07 0.02 3% 0% 1.00
-energy_efficiency -3% 1% -5% 0.66 0.08 0.02 7% 3% 1.00
-energy_efficiency -14% 9% -29% 0.70 0.09 0.03 0% 0% 1.00
-energy_efficiency 2% -11% 14% 0.59 0.07 0.02 -3% -2% 1.00
-energy_efficiency -5% 1% -5% 0.65 0.08 0.02 6% 0% 1.00
-california_housing 33% -7% 30% 0.91 1.71 0.41 -14% -19% 0.72
-california_housing 28% -7% 27% 0.91 1.74 0.41 -16% -22% 0.70
-california_housing 34% -7% 31% 0.90 1.67 0.40 -14% -20% 0.72
-california_housing 29% -7% 28% 0.91 1.72 0.41 -16% -22% 0.71
-california_housing 31% -8% 30% 0.90 1.69 0.40 -15% -20% 0.71
-airfoil_self_noise 8% -21% 27% 0.77 1.20 0.26 -168% -189% 0.74
-airfoil_self_noise -16% -14% -3% 0.84 1.68 0.36 -170% -214% 0.67
-airfoil_self_noise 20% -8% 24% 0.89 1.22 0.26 -95% -105% 0.84
-airfoil_self_noise 62% -18% 49% 0.81 0.82 0.18 -57% -81% 0.89
-airfoil_self_noise 18% -13% 26% 0.86 1.19 0.25 -93% -118% 0.85
-QSAR_fish_toxicity 11% -10% 19% 0.81 1.82 0.28 -6% -13% 0.53
-QSAR_fish_toxicity 13% -10% 21% 0.82 1.79 0.28 -5% -11% 0.54
-QSAR_fish_toxicity 14% -10% 21% 0.82 1.82 0.28 -8% -12% 0.51
-QSAR_fish_toxicity 11% -8% 17% 0.82 1.79 0.28 -5% -11% 0.53
-QSAR_fish_toxicity 9% -8% 16% 0.84 1.84 0.29 -6% -10% 0.52
+space_ga -8% 2% -11% 0.95 2.59 0.28 -13% -15% 0.55
+space_ga -3% 0% -3% 0.95 2.77 0.30 -14% -17% 0.55
+space_ga -2% 1% -4% 0.96 2.61 0.28 -13% -13% 0.57
+space_ga -9% 1% -11% 0.96 2.81 0.30 -15% -18% 0.54
+space_ga -7% 0% -8% 0.96 2.79 0.30 -19% -24% 0.51
+cpu_activity 11% -4% 13% 0.95 0.57 0.11 -13% -20% 0.98
+cpu_activity 14% -5% 16% 0.94 0.54 0.10 -13% -18% 0.98
+cpu_activity 9% -4% 11% 0.95 0.58 0.11 -13% -18% 0.98
+cpu_activity 10% -4% 13% 0.94 0.57 0.11 -15% -22% 0.97
+cpu_activity 10% -4% 13% 0.94 0.57 0.11 -16% -22% 0.97
+naval_propulsion_plant 131% 1% 56% 0.96 0.96 0.28 -320% -418% 0.96
+naval_propulsion_plant 144% 1% 59% 0.96 0.88 0.26 -286% -369% 0.97
+naval_propulsion_plant 174% 0% 63% 0.96 0.81 0.24 -257% -348% 0.97
+naval_propulsion_plant 127% 0% 56% 0.96 0.94 0.28 -324% -426% 0.96
+naval_propulsion_plant 108% 1% 52% 0.96 1.03 0.30 -391% -469% 0.95
+miami_housing -7% 0% -7% 0.95 1.22 0.15 -11% -20% 0.90
+miami_housing -10% -1% -10% 0.95 1.24 0.15 -6% -17% 0.90
+miami_housing -8% -1% -8% 0.95 1.25 0.15 -13% -23% 0.89
+miami_housing -9% -1% -10% 0.95 1.22 0.15 -9% -19% 0.90
+miami_housing -9% 0% -9% 0.94 1.24 0.15 -10% -22% 0.89
+kin8nm -10% 1% -12% 0.95 2.75 0.54 -36% -40% 0.46
+kin8nm -8% 0% -9% 0.94 2.66 0.52 -33% -36% 0.49
+kin8nm -11% 3% -15% 0.95 2.76 0.54 -32% -35% 0.50
+kin8nm -9% 2% -12% 0.94 2.67 0.52 -36% -40% 0.48
+kin8nm -10% 1% -13% 0.95 2.73 0.53 -33% -36% 0.49
+concrete_compressive_strength -18% -3% -18% 0.89 1.67 0.34 -78% -110% 0.74
+concrete_compressive_strength -5% -3% -3% 0.89 1.51 0.30 -71% -97% 0.76
+concrete_compressive_strength -6% -6% 0% 0.87 1.42 0.29 -63% -87% 0.77
+concrete_compressive_strength -6% -3% -4% 0.90 1.47 0.30 -63% -91% 0.77
+concrete_compressive_strength -6% 1% -7% 0.92 1.59 0.32 -61% -88% 0.77
+cars -35% 5% -60% 0.95 1.01 0.17 6% 3% 0.95
+cars -25% 1% -34% 0.91 0.83 0.14 8% 4% 0.95
+cars -19% 0% -24% 0.92 0.80 0.14 4% -2% 0.95
+cars -24% -3% -28% 0.90 0.86 0.15 2% 0% 0.95
+cars -21% -3% -22% 0.91 0.91 0.15 -3% 1% 0.94
+energy_efficiency 16% -8% 21% 0.90 0.18 0.05 0% 0% 1.00
+energy_efficiency 6% -2% 6% 0.97 0.21 0.06 4% 1% 1.00
+energy_efficiency 8% -4% 11% 0.95 0.21 0.06 5% 2% 1.00
+energy_efficiency 3% -3% 5% 0.96 0.21 0.06 1% 0% 1.00
+energy_efficiency 17% -4% 18% 0.95 0.19 0.06 5% 0% 1.00
+california_housing -2% -2% 1% 0.95 2.38 0.57 -16% -22% 0.70
+california_housing 3% -2% 5% 0.95 2.31 0.55 -13% -19% 0.72
+california_housing 2% -2% 4% 0.95 2.29 0.55 -15% -20% 0.71
+california_housing 0% -3% 3% 0.95 2.32 0.55 -16% -22% 0.71
+california_housing 1% -3% 4% 0.95 2.32 0.55 -14% -20% 0.72
+airfoil_self_noise -25% -5% -26% 0.94 2.17 0.46 -166% -184% 0.74
+airfoil_self_noise 0% -1% 1% 0.98 1.78 0.38 -59% -80% 0.89
+airfoil_self_noise -15% -1% -16% 0.98 1.98 0.42 -97% -106% 0.84
+airfoil_self_noise -28% -5% -32% 0.93 2.20 0.47 -169% -213% 0.67
+airfoil_self_noise -14% -3% -13% 0.96 1.93 0.41 -98% -125% 0.84
+QSAR_fish_toxicity -16% -2% -17% 0.88 2.45 0.38 -6% -10% 0.53
+QSAR_fish_toxicity -10% -2% -9% 0.89 2.43 0.38 -7% -12% 0.52
+QSAR_fish_toxicity 3% -6% 9% 0.88 2.38 0.37 -8% -12% 0.50
+QSAR_fish_toxicity -12% -4% -9% 0.88 2.41 0.38 -5% -10% 0.55
+QSAR_fish_toxicity 8% -9% 16% 0.84 2.03 0.32 -6% -13% 0.53
 Table 8: Test results for MultiMLPs with complete data at 95% quantiles for seeds 0, 10, 20, 30, 40, with rows ordered by seed (ascending). Performance over the baseline is highlighted in bold.
@@
 Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP MPIW NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
-space_ga 7% -4% 10% 0.79 1.22 0.13 -1% -1% 0.75
-space_ga 8% 2% 5% 0.82 1.25 0.14 0% 0% 0.76
-space_ga 7% 4% 3% 0.83 1.27 0.14 -3% -3% 0.74
-space_ga 8% -1% 8% 0.79 1.24 0.14 0% -1% 0.74
-space_ga 9% -2% 10% 0.81 1.27 0.14 1% 0% 0.76
-cpu_activity 8% -11% 16% 0.72 0.25 0.05 6% 5% 0.98
-cpu_activity 14% -15% 25% 0.68 0.23 0.04 8% 5% 0.98
-cpu_activity 8% -8% 15% 0.73 0.24 0.04 8% 7% 0.98
-cpu_activity 5% -3% 8% 0.74 0.25 0.05 6% 5% 0.98
-cpu_activity 6% -8% 12% 0.70 0.24 0.04 4% 2% 0.98
-naval_propulsion_plant 50% 0% 34% 0.92 0.19 0.06 -44% -48% 1.00
-naval_propulsion_plant 20% 0% 17% 0.93 0.20 0.06 -35% -13% 1.00
-naval_propulsion_plant 3% 1% 3% 0.92 0.26 0.08 -51% -45% 0.99
-naval_propulsion_plant 3% 3% 0% 0.94 0.23 0.07 -86% -63% 1.00
-naval_propulsion_plant 29% -6% 27% 0.91 0.23 0.07 -9% -11% 1.00
-miami_housing 13% -2% 13% 0.78 0.37 0.05 -2% 7% 0.91
-miami_housing 10% 3% 6% 0.81 0.40 0.05 -6% 1% 0.91
-miami_housing 14% 1% 10% 0.82 0.42 0.05 -2% 3% 0.92
-miami_housing 8% 1% 7% 0.82 0.43 0.05 -1% 4% 0.91
-miami_housing 3% 6% -4% 0.83 0.44 0.06 -4% 6% 0.91
-kin8nm -5% 9% -15% 0.81 0.69 0.13 6% 5% 0.93
-kin8nm 15% 1% 12% 0.78 0.62 0.12 7% 7% 0.93
-kin8nm 2% 4% -2% 0.82 0.67 0.13 10% 10% 0.94
-kin8nm 2% 14% -12% 0.83 0.69 0.14 10% 6% 0.94
-kin8nm -3% 12% -15% 0.83 0.71 0.14 4% 5% 0.93
-concrete_compressive_strength 89% -14% 54% 0.62 0.30 0.06 4% 10% 0.90
-concrete_compressive_strength 82% -21% 57% 0.55 0.29 0.06 9% 15% 0.91
-concrete_compressive_strength 120% -17% 62% 0.60 0.28 0.06 18% 19% 0.92
-concrete_compressive_strength 100% -33% 66% 0.52 0.27 0.06 9% 12% 0.91
-concrete_compressive_strength 101% -30% 65% 0.52 0.28 0.06 18% 20% 0.91
-cars 10% 3% 7% 0.77 0.38 0.06 -4% 5% 0.95
-cars -23% 28% -65% 0.88 0.54 0.09 1% 1% 0.96
-cars 16% -4% 17% 0.65 0.29 0.05 -6% -2% 0.95
-cars 2% -6% 8% 0.68 0.34 0.06 -2% -1% 0.95
-cars -5% 8% -13% 0.81 0.45 0.08 -5% -2% 0.95
-energy_efficiency 93% 59% 17% 0.70 0.07 0.02 37% 38% 1.00
-energy_efficiency 83% 27% 29% 0.61 0.06 0.02 28% 32% 1.00
-energy_efficiency 109% 28% 39% 0.64 0.06 0.02 35% 32% 1.00
-energy_efficiency 65% 26% 24% 0.58 0.07 0.02 31% 31% 1.00
-energy_efficiency 60% -15% 47% 0.58 0.06 0.02 31% 32% 1.00
-california_housing -13% 9% -25% 0.90 1.33 0.32 7% 8% 0.81
-california_housing -15% 6% -24% 0.89 1.27 0.30 7% 9% 0.82
-california_housing -11% 9% -23% 0.90 1.29 0.31 8% 11% 0.82
-california_housing -14% 8% -26% 0.89 1.31 0.31 7% 8% 0.81
-california_housing -18% 9% -33% 0.90 1.32 0.32 7% 9% 0.82
-airfoil_self_noise 140% -10% 62% 0.74 0.34 0.07 18% 18% 0.97
-airfoil_self_noise 42% 9% 23% 0.73 0.34 0.07 11% 9% 0.97
-airfoil_self_noise 43% -1% 31% 0.77 0.36 0.08 22% 22% 0.97
-airfoil_self_noise 39% 3% 26% 0.73 0.35 0.07 17% 12% 0.97
-airfoil_self_noise 105% -2% 52% 0.77 0.38 0.08 21% 19% 0.97
-QSAR_fish_toxicity -6% 8% -14% 0.72 1.40 0.22 -4% 6% 0.52
-QSAR_fish_toxicity 0% 17% -17% 0.82 1.50 0.23 6% 11% 0.54
-QSAR_fish_toxicity -5% 8% -13% 0.77 1.57 0.24 0% -1% 0.57
-QSAR_fish_toxicity -8% 13% -23% 0.78 1.51 0.23 10% 9% 0.61
-QSAR_fish_toxicity -12% 10% -25% 0.73 1.49 0.23 5% 11% 0.51
+space_ga 5% -2% 7% 0.95 2.11 0.23 0% -1% 0.74
+space_ga 3% -1% 3% 0.96 2.13 0.23 -1% -1% 0.75
+space_ga 6% -1% 7% 0.94 2.04 0.22 -3% -3% 0.74
+space_ga 5% -2% 6% 0.95 2.15 0.23 0% 0% 0.76
+space_ga 10% 0% 9% 0.95 2.06 0.22 1% 0% 0.76
+cpu_activity -18% 1% -22% 0.96 0.58 0.11 4% 2% 0.98
+cpu_activity -3% 0% -2% 0.95 0.51 0.10 8% 7% 0.98
+cpu_activity -9% 0% -10% 0.95 0.53 0.10 6% 5% 0.98
+cpu_activity -6% 0% -7% 0.95 0.54 0.10 6% 5% 0.98
+cpu_activity -2% 0% -2% 0.95 0.53 0.10 8% 5% 0.98
+naval_propulsion_plant 32% 1% 24% 0.95 0.23 0.07 -44% -48% 1.00
+naval_propulsion_plant 0% 1% -2% 0.96 0.31 0.09 -9% -11% 1.00
+naval_propulsion_plant 7% 0% 5% 0.95 0.24 0.07 -35% -13% 1.00
+naval_propulsion_plant -8% 1% -11% 0.96 0.28 0.08 -86% -63% 1.00
+naval_propulsion_plant -13% 2% -16% 0.96 0.34 0.10 -51% -45% 0.99
+miami_housing -38% 0% -60% 0.95 1.18 0.15 -6% 1% 0.91
+miami_housing -37% 0% -59% 0.95 1.16 0.14 -2% 7% 0.91
+miami_housing -43% 0% -74% 0.95 1.25 0.15 -4% 6% 0.91
+miami_housing -38% 1% -62% 0.95 1.15 0.14 -2% 3% 0.92
+miami_housing -33% -1% -47% 0.95 1.10 0.14 -1% 4% 0.91
+kin8nm 10% 1% 9% 0.95 1.03 0.20 4% 5% 0.93
+kin8nm 2% 1% 1% 0.96 1.07 0.21 10% 10% 0.94
+kin8nm 4% 1% 3% 0.95 1.06 0.21 10% 6% 0.94
+kin8nm 10% -3% 12% 0.93 0.99 0.19 6% 5% 0.93
+kin8nm 15% -1% 13% 0.94 1.03 0.20 7% 7% 0.93
+concrete_compressive_strength 7% -3% 9% 0.91 1.24 0.25 18% 19% 0.92
+concrete_compressive_strength 18% 1% 14% 0.97 1.39 0.28 9% 12% 0.91
+concrete_compressive_strength 2% 1% 1% 0.97 1.63 0.33 9% 15% 0.91
+concrete_compressive_strength 7% -5% 11% 0.94 1.62 0.33 4% 10% 0.90
+concrete_compressive_strength 19% -3% 18% 0.93 1.26 0.26 18% 20% 0.91
+cars 5% -9% 13% 0.88 0.66 0.11 -5% -2% 0.95
+cars 0% -3% 2% 0.93 0.80 0.14 -6% -2% 0.95
+cars -9% 0% -10% 0.93 0.80 0.14 1% 1% 0.96
+cars -8% -3% -4% 0.93 0.83 0.14 -2% -1% 0.95
+cars 25% -1% 21% 0.94 0.79 0.13 -4% 5% 0.95
+energy_efficiency 76% -3% 46% 0.94 0.16 0.04 31% 31% 1.00
+energy_efficiency 46% 7% 25% 0.99 0.19 0.06 35% 32% 1.00
+energy_efficiency 34% -1% 27% 0.95 0.17 0.05 28% 32% 1.00
+energy_efficiency 29% -3% 25% 0.95 0.16 0.05 31% 32% 1.00
+energy_efficiency 79% -3% 45% 0.96 0.17 0.05 37% 38% 1.00
+california_housing -13% 1% -16% 0.95 1.88 0.45 8% 11% 0.82
+california_housing -8% 0% -9% 0.95 1.78 0.42 7% 8% 0.81
+california_housing -16% 1% -20% 0.95 1.83 0.44 7% 9% 0.82
+california_housing -$-1$2% 0.94 1.82 0.43 7% 8% 0.81
+california_housing -12% 0% -13% 0.94 1.76 0.42 7% 9% 0.82
+airfoil_self_noise 98% -2% 51% 0.97 0.82 0.18 21% 19% 0.97
+airfoil_self_noise 33% 0% 24% 0.98 0.90 0.19 17% 12% 0.97
+airfoil_self_noise 36% -3% 28% 0.95 0.79 0.17 11% 9% 0.97
+airfoil_self_noise 83% -4% 47% 0.95 0.76 0.16 22% 22% 0.97
+airfoil_self_noise 93% -1% 49% 0.98 0.88 0.19 18% 18% 0.97
+QSAR_fish_toxicity -11% 2% -16% 0.91 2.53 0.39 5% 11% 0.51
+QSAR_fish_toxicity 3% 2% 1% 0.91 2.38 0.37 0% -1% 0.57
+QSAR_fish_toxicity 7% -1% 7% 0.91 2.31 0.36 6% 11% 0.54
+QSAR_fish_toxicity 18% 2% 14% 0.90 2.10 0.33 -4% 6% 0.52
+QSAR_fish_toxicity 9% -1% 9% 0.83 1.84 0.29 10% 9% 0.61
 Table 9: Test results for MultiXGBs with 50% missing data at 95% quantiles for seeds 0, 10, 20, 30, 40, with rows ordered by seed (ascending). Performance over the baseline is highlighted in bold.
@@
 Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP MPIW NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
-space_ga -12% 5% -20% 0.87 2.25 0.24 -11% -13% 0.35
-space_ga -9% 6% -18% 0.87 2.15 0.23 -12% -11% 0.44
-space_ga 3% 6% -5% 0.85 1.90 0.21 -1% -6% 0.46
-space_ga -8% 4% -14% 0.89 2.06 0.22 -5% -11% 0.54
-space_ga 0% 1% -4% 0.82 1.88 0.20 -6% -6% 0.45
-cpu_activity 2% 2% -1% 0.79 0.58 0.11 -30% -30% 0.73
-cpu_activity 9% 7% -12% 0.87 0.61 0.12 -7% -23% 0.94
-cpu_activity -41% 27% -127% 0.96 1.17 0.22 -1% -26% 0.81
-cpu_activity -8% 14% -34% 0.94 0.78 0.15 36% 5% 0.88
-cpu_activity 12% 8% -12% 0.91 0.66 0.12 -2% -11% 0.81
-naval_propulsion_plant 82% -9% 47% 0.76 0.64 0.19 -30% -42% 0.83
-naval_propulsion_plant 121% -13% 59% 0.76 0.65 0.19 5% -6% 0.84
-naval_propulsion_plant 127% -13% 58% 0.73 0.66 0.20 6% 6% 0.76
-naval_propulsion_plant 134% -21% 63% 0.64 0.52 0.15 -7% -11% 0.70
-naval_propulsion_plant 58% -25% 53% 0.62 0.64 0.19 -45% -90% 0.73
-miami_housing 16% -6% 13% 0.84 0.86 0.11 -4% -4% 0.72
-miami_housing -9% 4% -22% 0.90 0.94 0.12 -14% -20% 0.79
-miami_housing 25% -11% 17% 0.72 0.60 0.07 -35% -35% 0.54
-miami_housing 23% -4% 21% 0.85 0.83 0.10 -7% -12% 0.80
-miami_housing 25% -9% 21% 0.78 0.63 0.08 -10% -19% 0.76
-kin8nm -3% 2% -12% 0.88 2.41 0.47 0% 1% 0.40
-kin8nm -6% 3% -14% 0.89 2.47 0.48 -6% -8% 0.39
-kin8nm -4% 2% -10% 0.88 2.48 0.48 -5% -6% 0.37
-kin8nm -9% 7% -23% 0.90 2.57 0.50 -5% -7% 0.44
-kin8nm -3% 3% -10% 0.87 2.38 0.47 1% 2% 0.41
-concrete_compressive_strength -2% 0% -6% 0.71 1.51 0.30 -19% -22% 0.48
-concrete_compressive_strength 2% 11% -15% 0.81 1.44 0.29 -4% -3% 0.69
-concrete_compressive_strength 4% 11% -12% 0.76 1.44 0.29 -15% -24% 0.50
-concrete_compressive_strength 1% 8% -9% 0.77 1.47 0.30 -4% -9% 0.64
-concrete_compressive_strength 7% 8% -7% 0.80 1.49 0.30 -37% -44% 0.57
-cars -3% 24% -34% 0.85 1.34 0.23 -9% -16% 0.75
-cars -35% 9% -98% 0.82 1.78 0.30 -68% -80% 0.55
-cars -6% 29% -52% 0.88 1.54 0.26 -8% -14% 0.76
-cars -29% 32% -90% 0.81 1.73 0.29 -28% -39% 0.40
-cars -5% 34% -40% 0.92 1.49 0.25 -37% -25% 0.81
-energy_efficiency 20% 15% -13% 0.93 0.71 0.21 -37% -39% 0.94
-energy_efficiency 35% 3% 23% 0.88 0.58 0.17 -15% -24% 0.97
-energy_efficiency 37% 7% 19% 0.92 0.63 0.18 -16% -44% 0.96
-energy_efficiency -21% 18% -81% 0.95 1.01 0.29 -155% -137% 0.93
-energy_efficiency -22% 14% -66% 0.91 0.92 0.27 -705% -520% 0.76
-california_housing 30% -8% 28% 0.82 1.52 0.36 -4% -7% 0.61
-california_housing 26% -2% 21% 0.89 1.61 0.39 0% 1% 0.73
-california_housing 34% -1% 26% 0.90 1.66 0.40 -11% -13% 0.73
-california_housing 13% 0% 11% 0.90 1.85 0.44 -3% -3% 0.65
-california_housing 27% 0% 21% 0.91 1.70 0.41 -1% -3% 0.73
-airfoil_self_noise -8% 4% -14% 0.77 1.70 0.36 -14% -8% 0.52
-airfoil_self_noise -18% -2% -20% 0.80 1.77 0.38 -94% -110% 0.50
-airfoil_self_noise -12% -7% -11% 0.70 1.86 0.40 -43% -48% 0.19
-airfoil_self_noise 34% -5% 29% 0.69 1.02 0.22 -19% -29% 0.38
-airfoil_self_noise -6% -6% -5% 0.71 1.58 0.34 -14% -22% 0.43
-QSAR_fish_toxicity -3% -4% -1% 0.68 1.56 0.24 -4% -5% 0.38
-QSAR_fish_toxicity -1% -4% 2% 0.75 1.83 0.28 -9% -3% 0.35
-QSAR_fish_toxicity 3% 9% -8% 0.75 1.66 0.26 4% 4% 0.43
-QSAR_fish_toxicity -6% 1% -19% 0.72 1.61 0.25 -7% -3% 0.35
-QSAR_fish_toxicity -1% -3% -2% 0.74 1.84 0.29 -5% -3% 0.30
+space_ga 4% -4% 7% 0.93 2.82 0.31 -11% -13% 0.35
+space_ga 2% -1% 3% 0.94 2.97 0.32 -6% -6% 0.45
+space_ga -4% -3% -2% 0.94 3.06 0.33 -12% -11% 0.44
+space_ga -1% -1% -1% 0.95 2.97 0.32 -1% -6% 0.46
+space_ga 2% 0% 2% 0.95 2.77 0.30 -5% -11% 0.54
+cpu_activity -9% 1% -11% 0.95 0.89 0.17 -7% -23% 0.94
+cpu_activity -26% 1% -44% 0.93 1.02 0.19 -1% -26% 0.81
+cpu_activity -15% -2% -14% 0.89 0.80 0.15 -30% -30% 0.73
+cpu_activity -1% -1% -3% 0.95 0.86 0.16 36% 5% 0.88
+cpu_activity -19% 2% -26% 0.98 1.12 0.21 -2% -11% 0.81
+naval_propulsion_plant -8% 0% -11% 0.98 2.93 0.86 -45% -90% 0.73
+naval_propulsion_plant 31% -2% 24% 0.93 1.50 0.44 5% -6% 0.84
+naval_propulsion_plant 18% -2% 17% 0.96 2.06 0.61 -30% -42% 0.83
+naval_propulsion_plant 53% -5% 37% 0.87 1.27 0.37 6% 6% 0.76
+naval_propulsion_plant -5% -3% -6% 0.93 2.22 0.65 -7% -11% 0.70
+miami_housing -10% -2% -11% 0.94 1.45 0.18 -7% -12% 0.80
+miami_housing -38% 0% -63% 0.93 1.86 0.23 -35% -35% 0.54
+miami_housing -26% -1% -33% 0.95 1.74 0.22 -10% -19% 0.76
+miami_housing 1% -3% 2% 0.95 1.35 0.17 -14% -20% 0.79
+miami_housing -17% 1% -22% 0.94 1.64 0.20 -4% -4% 0.72
+kin8nm -6% 0% -8% 0.97 3.36 0.66 0% 1% 0.40
+kin8nm -15% 1% -22% 0.97 3.46 0.68 -6% -8% 0.39
+kin8nm -8% 3% -12% 0.95 3.00 0.59 1% 2% 0.41
+kin8nm -9% 2% -14% 0.96 3.11 0.61 -5% -7% 0.44
+kin8nm -8% 1% -9% 0.95 3.23 0.63 -5% -6% 0.37
+concrete_compressive_strength -8% -3% -8% 0.95 2.81 0.57 -19% -22% 0.48
+concrete_compressive_strength 7% -5% 11% 0.93 2.79 0.56 -37% -44% 0.57
+concrete_compressive_strength 14% -3% 13% 0.97 2.69 0.55 -4% -9% 0.64
+concrete_compressive_strength 4% -1% 3% 0.95 2.83 0.57 -15% -24% 0.50
+concrete_compressive_strength -11% 0% -14% 0.97 3.01 0.61 -4% -3% 0.69
+cars 31% -3% 24% 0.93 1.91 0.32 -9% -16% 0.75
+cars 51% -7% 38% 0.93 1.84 0.31 -37% -25% 0.81
+cars -16% -11% -17% 0.89 2.17 0.37 -68% -80% 0.55
+cars -13% -3% -21% 0.94 2.18 0.37 -8% -14% 0.76
+cars -5% -12% 5% 0.83 2.00 0.34 -28% -39% 0.40
+energy_efficiency -11% 0% -13% 1.00 1.15 0.33 -15% -24% 0.97
+energy_efficiency 20% -10% 25% 0.89 0.84 0.24 -705% -520% 0.76
+energy_efficiency 42% -1% 23% 0.95 0.78 0.23 -37% -39% 0.94
+energy_efficiency 61% -4% 40% 0.94 0.70 0.20 -16% -44% 0.96
+energy_efficiency 20% -3% 16% 0.97 1.12 0.33 -155% -137% 0.93
+california_housing -14% 0% -17% 0.96 2.67 0.64 -3% -3% 0.65
+california_housing 0% -2% 0% 0.96 2.41 0.57 -11% -13% 0.73
+california_housing -7% -1% -8% 0.96 2.48 0.59 0% 1% 0.73
+california_housing 2% -3% 3% 0.92 2.28 0.54 -4% -7% 0.61
+california_housing -12% -1% -16% 0.97 2.63 0.63 -1% -3% 0.73
+airfoil_self_noise -30% -2% -43% 0.97 3.95 0.84 -94% -110% 0.50
+airfoil_self_noise -18% -1% -22% 0.91 2.98 0.64 -19% -29% 0.38
+airfoil_self_noise -23% -4% -25% 0.91 3.49 0.75 -43% -48% 0.19
+airfoil_self_noise -31% 1% -49% 0.97 3.61 0.77 -14% -8% 0.52
+airfoil_self_noise -17% -1% -21% 0.97 3.81 0.81 -14% -22% 0.43
+QSAR_fish_toxicity -8% 3% -14% 0.98 4.22 0.66 4% 4% 0.43
+QSAR_fish_toxicity -6% 1% -7% 0.97 3.93 0.61 -9% -3% 0.35
+QSAR_fish_toxicity 8% -5% 10% 0.88 2.59 0.40 -4% -5% 0.38
+QSAR_fish_toxicity -5% -5% -5% 0.88 2.83 0.44 -7% -3% 0.35
+QSAR_fish_toxicity 12% -11% 15% 0.85 2.58 0.40 -5% -3% 0.30
 Table 10: Test results for MultiETs with 50% missing data at 95% quantiles for seeds 0, 10, 20, 30, 40, with rows ordered by seed (ascending). Performance over the baseline is highlighted in bold.
@@
 Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP MPIW NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
-space_ga 7% -6% 12% 0.86 2.17 0.23 -17% -18% 0.30
-space_ga 3% -5% 8% 0.87 2.14 0.23 -14% -15% 0.41
-space_ga 2% -5% 5% 0.88 2.23 0.24 -9% -10% 0.41
-space_ga 7% -5% 9% 0.86 2.21 0.24 -7% -7% 0.40
-space_ga -1% -3% 0% 0.91 2.39 0.26 -14% -18% 0.47
-cpu_activity -8% -5% -8% 0.93 0.90 0.17 -36% -46% 0.93
-cpu_activity -8% -6% -4% 0.92 0.89 0.17 2% -23% 0.87
-cpu_activity -30% -3% -43% 0.94 1.33 0.25 -36% -80% 0.73
-cpu_activity -27% -3% -41% 0.94 1.21 0.23 -13% -49% 0.74
-cpu_activity -6% -6% -3% 0.91 1.00 0.19 -65% -70% 0.67
-naval_propulsion_plant 55% -22% 48% 0.77 1.29 0.38 -18% -50% 0.64
-naval_propulsion_plant 55% -10% 42% 0.89 1.52 0.45 -66% -158% 0.79
-naval_propulsion_plant 54% -15% 44% 0.84 1.44 0.42 -29% -73% 0.74
-naval_propulsion_plant 67% -16% 50% 0.83 1.29 0.38 -56% -142% 0.74
-naval_propulsion_plant 57% -20% 48% 0.78 1.38 0.41 -14% -42% 0.66
-miami_housing 20% -6% 22% 0.92 1.33 0.17 -13% -26% 0.78
-miami_housing 32% -16% 31% 0.79 0.87 0.11 -41% -50% 0.48
-miami_housing -4% -5% -1% 0.92 1.39 0.17 -38% -51% 0.67
-miami_housing 37% -14% 36% 0.84 0.88 0.11 -21% -34% 0.79
-miami_housing 16% -7% 16% 0.89 1.30 0.16 -20% -26% 0.65
-kin8nm 5% -9% 13% 0.86 2.48 0.49 -10% -12% 0.32
-kin8nm 4% -8% 9% 0.87 2.58 0.51 -10% -12% 0.33
-kin8nm 5% -8% 12% 0.87 2.62 0.51 -13% -15% 0.30
-kin8nm 8% -11% 17% 0.84 2.41 0.47 -8% -8% 0.30
-kin8nm 1% -5% 6% 0.90 2.66 0.52 -14% -18% 0.37
-concrete_compressive_strength 4% -11% 15% 0.79 1.88 0.38 -24% -36% 0.46
-concrete_compressive_strength -3% -17% 12% 0.78 1.74 0.35 -26% -36% 0.42
-concrete_compressive_strength 4% -12% 13% 0.81 1.83 0.37 -30% -47% 0.56
-concrete_compressive_strength 7% -11% 15% 0.83 1.81 0.37 -19% -31% 0.58
-concrete_compressive_strength -7% -18% 9% 0.80 1.87 0.38 -76% -99% 0.45
-cars -18% 1% -39% 0.89 1.29 0.22 -16% -15% 0.81
-cars -18% -11% -11% 0.84 1.39 0.23 -23% -30% 0.72
-cars -33% -6% -54% 0.84 1.63 0.28 -59% -59% 0.60
-cars -28% -12% -41% 0.80 1.27 0.21 -18% -10% 0.59
-cars -19% -14% -29% 0.79 1.48 0.25 -49% -42% 0.59
-energy_efficiency -11% 13% -35% 0.91 0.50 0.14 -7% -13% 0.98
-energy_efficiency 0% -7% 0% 0.83 0.49 0.14 -8% -20% 0.95
-energy_efficiency -9% -14% 0% 0.76 0.47 0.14 -112% -137% 0.93
-energy_efficiency -11% 4% -30% 0.87 0.59 0.17 -38% -76% 0.96
-energy_efficiency -46% 2% -114% 0.89 0.52 0.15 -198% -176% 0.86
-california_housing 30% -7% 28% 0.92 2.09 0.50 -11% -13% 0.64
-california_housing 28% -9% 28% 0.87 2.02 0.48 -14% -18% 0.50
-california_housing 13% -6% 17% 0.92 2.24 0.53 -19% -23% 0.64
-california_housing 19% -5% 20% 0.93 2.23 0.53 -12% -16% 0.64
-california_housing 16% -6% 18% 0.93 2.35 0.56 -11% -16% 0.60
-airfoil_self_noise -7% -18% 10% 0.77 2.02 0.43 -27% -51% 0.36
-airfoil_self_noise -26% -18% -11% 0.80 2.21 0.47 -171% -229% 0.30
-airfoil_self_noise 25% -20% 36% 0.74 1.28 0.27 -28% -38% 0.41
-airfoil_self_noise -22% -24% -2% 0.73 2.26 0.48 -54% -86% 0.13
-airfoil_self_noise -7% -18% 10% 0.79 2.02 0.43 -44% -58% 0.43
-QSAR_fish_toxicity 11% -8% 15% 0.85 2.37 0.37 -11% -10% 0.40
-QSAR_fish_toxicity 13% -8% 18% 0.82 2.02 0.31 -1% -5% 0.47
-QSAR_fish_toxicity 19% -7% 22% 0.84 1.90 0.29 -3% -3% 0.48
-QSAR_fish_toxicity -3% -15% 12% 0.76 2.02 0.32 -24% -27% 0.25
-QSAR_fish_toxicity 8% -9% 14% 0.83 2.29 0.36 -3% -5% 0.36
+space_ga -7% -2% -6% 0.94 3.08 0.33 -18% -19% 0.29
+space_ga -12% 3% -18% 0.97 3.41 0.37 -7% -7% 0.40
+space_ga -13% 0% -16% 0.95 3.05 0.33 -13% -11% 0.42
+space_ga -13% 0% -18% 0.96 3.21 0.35 -8% -8% 0.41
+space_ga -10% 0% -11% 0.96 2.91 0.32 -15% -18% 0.46
+cpu_activity -18% -3% -23% 0.95 1.03 0.19 -40% -56% 0.92
+cpu_activity -12% -6% -11% 0.91 1.04 0.20 -15% -44% 0.81
+cpu_activity -5% -5% -2% 0.92 1.00 0.19 -43% -59% 0.75
+cpu_activity -13% -2% -14% 0.96 0.97 0.18 2% -19% 0.87
+cpu_activity -30% -1% -42% 0.96 1.31 0.25 -10% -43% 0.76
+naval_propulsion_plant -19% 0% -23% 0.98 3.17 0.94 -63% -158% 0.71
+naval_propulsion_plant 11% -5% 13% 0.94 2.24 0.66 -31% -78% 0.73
+naval_propulsion_plant 14% -2% 14% 0.97 2.26 0.67 -61% -147% 0.80
+naval_propulsion_plant 24% -9% 25% 0.89 1.99 0.59 -13% -42% 0.66
+naval_propulsion_plant -10% -2% -11% 0.95 2.75 0.81 -19% -52% 0.64
+miami_housing -15% -2% -17% 0.94 1.86 0.23 -11% -25% 0.79
+miami_housing -44% 5% -87% 0.96 2.39 0.30 -43% -64% 0.47
+miami_housing -41% 0% -69% 0.96 2.30 0.29 -27% -42% 0.72
+miami_housing -35% -2% -52% 0.96 2.18 0.27 -21% -33% 0.79
+miami_housing -31% 1% -48% 0.95 2.24 0.28 -22% -30% 0.64
+kin8nm -15% -1% -19% 0.96 3.55 0.69 -8% -8% 0.29
+kin8nm -14% 2% -19% 0.96 3.43 0.67 -10% -14% 0.32
+kin8nm -11% 0% -13% 0.94 3.20 0.63 -9% -12% 0.32
+kin8nm -14% 0% -18% 0.95 3.30 0.65 -14% -18% 0.37
+kin8nm -14% 2% -19% 0.96 3.42 0.67 -13% -16% 0.30
+concrete_compressive_strength -20% 4% -35% 0.96 3.01 0.61 -30% -43% 0.41
+concrete_compressive_strength -16% -8% -10% 0.92 3.08 0.62 -72% -94% 0.47
+concrete_compressive_strength -18% -3% -24% 0.96 2.99 0.60 -21% -33% 0.56
+concrete_compressive_strength -25% -4% -34% 0.94 2.86 0.58 -26% -36% 0.41
+concrete_compressive_strength -22% 2% -31% 0.96 3.05 0.62 -35% -50% 0.53
+cars -24% -6% -27% 0.92 1.99 0.34 -22% -23% 0.73
+cars -49% 4% -103% 0.97 2.70 0.46 -20% -17% 0.80
+cars -14% -5% -12% 0.93 2.02 0.34 -59% -56% 0.60
+cars -36% -6% -54% 0.93 2.44 0.41 -37% -37% 0.66
+cars -37% -6% -68% 0.89 2.12 0.36 -50% -25% 0.34
+energy_efficiency 2% -3% 5% 0.97 0.62 0.18 -18% -32% 0.98
+energy_efficiency -42% -6% -62% 0.94 0.85 0.25 -159% -104% 0.90
+energy_efficiency -4% -6% 2% 0.93 0.69 0.20 -8% -20% 0.95
+energy_efficiency -29% -4% -41% 0.94 0.79 0.23 -39% -82% 0.95
+energy_efficiency -75% 0% -316% 1.00 2.61 0.76 -122% -136% 0.92
+california_housing -2% -2% 0% 0.96 2.88 0.69 -11% -16% 0.60
+california_housing -9% -2% -8% 0.96 2.90 0.69 -19% -23% 0.64
+california_housing -5% -2% -4% 0.97 3.02 0.72 -11% -13% 0.64
+california_housing -1% -3% 1% 0.93 2.77 0.66 -15% -18% 0.50
+california_housing -10% -1% -10% 0.97 3.07 0.73 -13% -17% 0.64
+airfoil_self_noise -33% -3% -46% 0.97 3.62 0.77 -163% -218% 0.34
+airfoil_self_noise -36% -2% -56% 0.93 3.42 0.73 -27% -36% 0.42
+airfoil_self_noise -31% -5% -40% 0.92 3.42 0.73 -50% -83% 0.17
+airfoil_self_noise -32% -2% -45% 0.97 3.91 0.83 -45% -57% 0.43
+airfoil_self_noise -22% -2% -30% 0.94 3.29 0.70 -27% -51% 0.36
+QSAR_fish_toxicity 2% 1% 0% 0.98 4.12 0.64 4% 2% 0.51
+QSAR_fish_toxicity -10% -4% -7% 0.93 3.43 0.53 -12% -12% 0.40
+QSAR_fish_toxicity -7% 1% -11% 0.93 2.87 0.45 -6% -8% 0.45
+QSAR_fish_toxicity -17% -7% -16% 0.86 2.99 0.47 -23% -26% 0.25
+QSAR_fish_toxicity 0% -3% -1% 0.90 2.63 0.41 -4% -6% 0.34
 Table 11: Test results for MultiMLPs with 50% missing data at 95% quantiles for seeds 0, 10, 20, 30, 40, with rows ordered by seed (ascending). Performance over the baseline is highlighted in bold.
@@
 Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP MPIW NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
-space_ga -12% -16% 4% 0.68 1.50 0.16 -41% -40% 0.15
-space_ga -21% -17% -11% 0.64 1.41 0.15 -85% -61% -0.21
-space_ga -5% -14% 6% 0.65 1.30 0.14 -21% -20% 0.37
-space_ga -9% -13% 4% 0.70 1.50 0.16 -28% -25% 0.35
-space_ga -9% -13% 1% 0.73 1.49 0.16 -25% -26% 0.38
-cpu_activity -38% 14% -89% 0.87 0.75 0.14 14% 1% 0.92
-cpu_activity -60% 32% -240% 0.94 1.08 0.20 -20% -17% 0.78
-cpu_activity -67% 36% -355% 0.95 1.50 0.28 -30% -34% 0.72
-cpu_activity -24% -2% -28% 0.68 0.39 0.07 -3% -16% 0.78
-cpu_activity -63% 28% -251% 0.85 1.13 0.21 -37% -49% 0.54
-naval_propulsion_plant -41% 0% -70% 0.63 0.85 0.25 -62% -68% -0.21
-naval_propulsion_plant -53% -13% -87% 0.82 1.60 0.47 -141% -163% 0.50
-naval_propulsion_plant -48% -3% -85% 0.72 1.23 0.36 -99% -94% 0.16
-naval_propulsion_plant -88% -61% -210% 0.29 1.10 0.33 -215% -562% -0.85
-naval_propulsion_plant -33% 4% -55% 0.70 1.04 0.31 -24% -27% 0.27
-miami_housing -48% 22% -148% 0.93 1.33 0.17 -21% -20% 0.74
-miami_housing -54% 27% -202% 0.89 1.33 0.17 -41% -48% 0.53
-miami_housing -19% 11% -40% 0.84 0.75 0.09 -18% -19% 0.75
-miami_housing -17% 6% -32% 0.82 0.69 0.09 -23% -22% 0.77
-miami_housing -28% 1% -39% 0.76 0.82 0.10 -24% -32% 0.57
-kin8nm -7% -5% -3% 0.75 1.67 0.33 -12% -10% 0.39
-kin8nm -11% -6% -8% 0.77 1.80 0.35 -20% -22% 0.36
-kin8nm -19% -4% -19% 0.77 1.88 0.37 -20% -19% 0.31
-kin8nm -4% -2% -5% 0.77 1.68 0.33 -3% 0% 0.43
-kin8nm -15% -9% -8% 0.78 1.79 0.35 -26% -24% 0.36
-concrete_compressive_strength -15% 5% -26% 0.74 1.53 0.31 -9% -6% 0.51
-concrete_compressive_strength -5% -38% 34% 0.40 0.57 0.12 -10% -10% 0.46
-concrete_compressive_strength -32% 21% -80% 0.84 1.69 0.34 -5% -9% 0.60
-concrete_compressive_strength 2% -53% 54% 0.27 0.39 0.08 -20% -23% 0.52
-concrete_compressive_strength -25% 0% -33% 0.76 1.37 0.28 -15% -13% 0.59
-cars -50% 38% -176% 0.93 1.62 0.27 -36% -38% 0.75
-cars -44% 49% -169% 0.93 1.55 0.26 -3% -5% 0.75
-cars -53% 32% -232% 0.94 2.04 0.34 -34% -33% 0.70
-cars -49% 44% -223% 0.89 1.59 0.27 -32% -15% 0.48
-cars -57% 51% -257% 0.93 1.90 0.32 -32% -37% 0.63
-energy_efficiency -35% 4% -68% 0.81 0.43 0.13 -8% 1% 0.97
-energy_efficiency -12% 25% -47% 0.60 0.27 0.08 8% 2% 0.92
-energy_efficiency -19% -14% -5% 0.57 0.31 0.09 -1% -32% 0.93
-energy_efficiency -18% 7% -29% 0.83 0.50 0.14 -8% 6% 0.96
-energy_efficiency 20% 26% -3% 0.51 0.12 0.03 14% 18% 0.98
-california_housing -16% 1% -20% 0.88 1.76 0.42 -7% -3% 0.63
-california_housing -14% 4% -25% 0.83 1.66 0.40 -4% -3% 0.54
-california_housing -9% 0% -10% 0.89 1.79 0.43 -1% 0% 0.68
-california_housing -20% 5% -30% 0.90 1.91 0.45 -2% -1% 0.65
-california_housing -27% 10% -52% 0.89 2.09 0.50 -8% -9% 0.55
-airfoil_self_noise 3% -1% 0% 0.76 1.22 0.26 -3% 1% 0.54
-airfoil_self_noise 27% -7% 26% 0.77 0.84 0.18 -20% -11% 0.84
-airfoil_self_noise 8% -7% 14% 0.62 0.63 0.14 -22% -13% 0.31
-airfoil_self_noise -9% 1% -12% 0.76 1.44 0.31 -23% -20% 0.44
-airfoil_self_noise -20% 6% -33% 0.83 1.62 0.35 -9% -1% 0.62
-QSAR_fish_toxicity -17% -1% -22% 0.75 1.89 0.29 -21% -28% 0.28
-QSAR_fish_toxicity -24% 9% -45% 0.79 1.97 0.31 -8% -9% 0.44
-QSAR_fish_toxicity -19% 0% -23% 0.70 1.56 0.24 -6% -8% 0.39
-QSAR_fish_toxicity -17% 16% -41% 0.76 1.72 0.27 -21% -17% 0.28
-QSAR_fish_toxicity -9% 8% -18% 0.81 1.98 0.31 -10% -9% 0.32
-Appendix F Example of learning with non-normality
+space_ga -25% -3% -32% 0.95 3.45 0.37 -41% -40% 0.15
+space_ga -22% -3% -27% 0.96 3.58 0.39 -28% -25% 0.35
+space_ga -45% -2% -82% 0.95 4.52 0.49 -85% -61% -0.21
+space_ga -13% -2% -13% 0.95 3.12 0.34 -21% -20% 0.37
+space_ga -20% -6% -18% 0.92 3.07 0.33 -25% -26% 0.38
+cpu_activity -22% -2% -32% 0.95 1.14 0.21 14% 1% 0.92
+cpu_activity -44% 2% -80% 0.95 1.40 0.27 -30% -34% 0.72
+cpu_activity -49% 1% -100% 0.94 1.44 0.27 -37% -49% 0.54
+cpu_activity -45% 2% -89% 0.96 1.23 0.23 -20% -17% 0.78
+cpu_activity -38% -1% -61% 0.96 1.29 0.24 -3% -16% 0.78
+naval_propulsion_plant -36% -4% -52% 0.95 5.40 1.59 -215% -562% -0.85
+naval_propulsion_plant -44% -2% -78% 0.95 4.14 1.22 -99% -94% 0.16
+naval_propulsion_plant -64% -1% -175% 0.98 4.66 1.37 -141% -163% 0.50
+naval_propulsion_plant -35% 2% -67% 0.92 3.53 1.04 -24% -27% 0.27
+naval_propulsion_plant -38% -2% -70% 0.94 4.75 1.40 -62% -68% -0.21
+miami_housing -32% -1% -46% 0.94 1.48 0.18 -21% -20% 0.74
+miami_housing -46% 1% -88% 0.96 1.98 0.25 -41% -48% 0.53
+miami_housing -25% -2% -31% 0.95 1.69 0.21 -18% -19% 0.75
+miami_housing -34% -2% -49% 0.96 1.82 0.23 -23% -22% 0.77
+miami_housing -42% -2% -72% 0.91 1.91 0.24 -24% -32% 0.57
+kin8nm -14% 1% -17% 0.98 3.68 0.72 -3% 0% 0.43
+kin8nm -18% 0% -22% 0.97 3.67 0.72 -20% -22% 0.36
+kin8nm -17% 1% -22% 0.96 3.56 0.70 -12% -10% 0.39
+kin8nm -24% 0% -32% 0.96 3.46 0.68 -26% -24% 0.36
+kin8nm -20% 0% -25% 0.95 3.51 0.69 -20% -19% 0.31
+concrete_compressive_strength -10% 2% -13% 0.95 2.80 0.57 -9% -6% 0.51
+concrete_compressive_strength 9% -3% 8% 0.95 2.68 0.54 -15% -13% 0.59
+concrete_compressive_strength -1% -4% 3% 0.95 2.86 0.58 -20% -23% 0.52
+concrete_compressive_strength 16% -5% 16% 0.91 2.49 0.50 -10% -10% 0.46
+concrete_compressive_strength 13% -3% 11% 0.96 2.40 0.48 -5% -9% 0.60
+cars -23% -3% -33% 0.95 1.93 0.33 -3% -5% 0.75
+cars -28% -2% -43% 0.97 1.94 0.33 -36% -38% 0.75
+cars -27% -3% -39% 0.94 2.06 0.35 -34% -33% 0.70
+cars -41% 6% -85% 0.97 2.70 0.46 -32% -37% 0.63
+cars 8% -4% 10% 0.90 1.68 0.28 -32% -15% 0.48
+energy_efficiency -12% 0% -14% 0.99 1.44 0.42 -8% 1% 0.97
+energy_efficiency -4% -2% -2% 0.97 0.69 0.20 14% 18% 0.98
+energy_efficiency -24% -3% -37% 0.94 1.23 0.36 8% 2% 0.92
+energy_efficiency 10% -3% -1% 0.96 1.07 0.31 -8% 6% 0.96
+energy_efficiency -7% -10% 0% 0.89 0.84 0.24 -1% -32% 0.93
+california_housing -28% 1% -42% 0.95 3.00 0.71 -8% -9% 0.55
+california_housing -20% -2% -25% 0.97 2.85 0.68 -1% 0% 0.68
+california_housing -30% 0% -42% 0.96 2.90 0.69 -7% -3% 0.63
+california_housing -16% -1% -19% 0.93 2.70 0.64 -4% -3% 0.54
+california_housing -28% 0% -40% 0.96 2.83 0.67 -2% -1% 0.65
+airfoil_self_noise -14% 0% -16% 1.00 3.46 0.74 -20% -11% 0.84
+airfoil_self_noise 2% -6% 7% 0.89 2.60 0.56 -22% -13% 0.31
+airfoil_self_noise -12% -8% -9% 0.88 2.56 0.55 -23% -20% 0.44
+airfoil_self_noise -13% -5% -12% 0.93 2.94 0.63 -9% -1% 0.62
+airfoil_self_noise -5% 1% -8% 0.94 2.77 0.59 -3% 1% 0.54
+QSAR_fish_toxicity -16% 1% -20% 0.99 4.78 0.74 -8% -9% 0.44
+QSAR_fish_toxicity -7% -5% -7% 0.92 3.31 0.51 -21% -28% 0.28
+QSAR_fish_toxicity -5% -2% -8% 0.95 3.17 0.49 -6% -8% 0.39
+QSAR_fish_toxicity -8% -5% -8% 0.86 2.74 0.43 -21% -17% 0.28
+QSAR_fish_toxicity -17% 2% -28% 0.90 3.03 0.47 -10% -9% 0.32
+Appendix F Aggregated results for non-conformalized predictions
+Table 12: Test results for all models with complete-raw data at 95% quantiles aggregated over five seeds.
+For each metric, the mean and standard deviation of the performance across the seeds are separated by $\pm$.
+Performance over the baseline is highlighted in bold.
+Interval Predictions Point Predictions
+Relative Absolute Relative Absolute
+Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
+MultiXGBs
+space_ga -1%$\pm$56%$\pm$2-7%$\pm$60.89$\pm$0.01 0.20$\pm$0.01 -9%$\pm$1-10%$\pm$20.60$\pm$0.01
+cpu_activity 25%$\pm$69%$\pm$112%$\pm$40.89$\pm$0.01 0.07$\pm$0.00 21%$\pm$1-1%$\pm$20.98$\pm$0.00
+naval_propulsion_plant 156%$\pm$138%$\pm$158%$\pm$20.96$\pm$0.00 $0$14%$\pm$4-12%$\pm$20.99$\pm$0.00
+miami_housing 61%$\pm$5-7%$\pm$243%$\pm$30.82$\pm$0.01 0.06$\pm$0.00 4%$\pm$43%$\pm$30.91$\pm$0.01
+kin8nm 10%$\pm$35%$\pm$24%$\pm$40.90$\pm$0.01 0.38$\pm$0.01 -20%$\pm$1-22%$\pm$10.63$\pm$0.01
+concrete_compressive_strength 13%$\pm$426%$\pm$6-11%$\pm$70.91$\pm$0.01 $0$26%$\pm$6-40%$\pm$110.85$\pm$0.01
+cars 18%$\pm$815%$\pm$42%$\pm$80.89$\pm$0.02 0.12$\pm$0.01 -3%$\pm$3-1%$\pm$30.95$\pm$0.00
+energy_efficiency 247%$\pm$8922%$\pm$660%$\pm$170.88$\pm$0.02 0.05$\pm$0.01 -15%$\pm$22-16%$\pm$171.00$\pm$0.00
+california_housing 31%$\pm$50%$\pm$123%$\pm$40.88$\pm$0.01 $0$1%$\pm$1-1%$\pm$10.81$\pm$0.00
+airfoil_self_noise 61%$\pm$414%$\pm$931%$\pm$160.81$\pm$0.07 0.19$\pm$0.04 -64%$\pm$32-73%$\pm$410.86$\pm$0.05
+QSAR_fish_toxicity 10%$\pm$311%$\pm$3-1%$\pm$30.76$\pm$0.01 0.23$\pm$0.00 3%$\pm$31%$\pm$40.55$\pm$0.02
+MultiETs
+space_ga 9%$\pm$3-2%$\pm$19%$\pm$40.91$\pm$0.01 0.22$\pm$0.01 -16%$\pm$2-18%$\pm$30.54$\pm$0.02
+cpu_activity 27%$\pm$3-8%$\pm$027%$\pm$20.90$\pm$0.00 $0$14%$\pm$1-20%$\pm$20.98$\pm$0.00
+naval_propulsion_plant 219%$\pm$32-9%$\pm$071%$\pm$30.91$\pm$0.00 0.19$\pm$0.02 -320%$\pm$41-409%$\pm$360.96$\pm$0.01
+miami_housing 69%$\pm$3-13%$\pm$048%$\pm$10.86$\pm$0.00 $0$10%$\pm$2-20%$\pm$10.90$\pm$0.00
+kin8nm 10%$\pm$2-9%$\pm$017%$\pm$20.88$\pm$0.01 0.44$\pm$0.01 -35%$\pm$2-38%$\pm$20.48$\pm$0.01
+concrete_compressive_strength 9%$\pm$9-12%$\pm$219%$\pm$80.81$\pm$0.01 0.24$\pm$0.02 -67%$\pm$7-94%$\pm$90.76$\pm$0.01
+cars 11%$\pm$616%$\pm$2-5%$\pm$70.83$\pm$0.02 0.09$\pm$0.01 3%$\pm$31%$\pm$20.95$\pm$0.00
+energy_efficiency -4%$\pm$50%$\pm$7-5%$\pm$140.65$\pm$0.03 0.02$\pm$0.00 3%$\pm$30%$\pm$11.00$\pm$0.00
+california_housing 31%$\pm$2-7%$\pm$029%$\pm$10.90$\pm$0.00 0.41$\pm$0.01 -15%$\pm$1-21%$\pm$10.71$\pm$0.01
+airfoil_self_noise 18%$\pm$25-15%$\pm$425%$\pm$170.83$\pm$0.04 0.26$\pm$0.06 -117%$\pm$45-141%$\pm$510.80$\pm$0.08
+QSAR_fish_toxicity 12%$\pm$2-9%$\pm$119%$\pm$20.82$\pm$0.01 $0$6%$\pm$1-11%$\pm$10.53$\pm$0.01
+MultiMLPs
+space_ga 8%$\pm$10%$\pm$37%$\pm$30.81$\pm$0.01 0.14$\pm$0.00 0%$\pm$1-1%$\pm$10.75$\pm$0.01
+cpu_activity 8%$\pm$3-9%$\pm$415%$\pm$60.72$\pm$0.02 0.05$\pm$0.00 7%$\pm$25%$\pm$20.98$\pm$0.00
+naval_propulsion_plant 21%$\pm$180%$\pm$316%$\pm$130.92$\pm$0.01 0.07$\pm$0.01 -45%$\pm$25-36%$\pm$201.00$\pm$0.00
+miami_housing 9%$\pm$42%$\pm$36%$\pm$60.81$\pm$0.02 $0$3%$\pm$24%$\pm$20.91$\pm$0.00
+kin8nm 2%$\pm$78%$\pm$5-6%$\pm$100.81$\pm$0.02 0.13$\pm$0.01 7%$\pm$27%$\pm$20.93$\pm$0.00
+concrete_compressive_strength 98%$\pm$13-23%$\pm$761%$\pm$50.56$\pm$0.04 0.06$\pm$0.00 12%$\pm$515%$\pm$40.91$\pm$0.01
+cars 0%$\pm$146%$\pm$12-9%$\pm$300.76$\pm$0.08 0.07$\pm$0.01 -3%$\pm$30%$\pm$30.95$\pm$0.00
+energy_efficiency 82%$\pm$1825%$\pm$2331%$\pm$110.62$\pm$0.04 0.02$\pm$0.00 32%$\pm$333%$\pm$31.00$\pm$0.00
+california_housing -14%$\pm$28%$\pm$1-26%$\pm$30.89$\pm$0.01 0.31$\pm$0.01 7%$\pm$09%$\pm$10.82$\pm$0.00
+airfoil_self_noise 74%$\pm$410%$\pm$639%$\pm$150.75$\pm$0.02 0.08$\pm$0.00 18%$\pm$416%$\pm$50.97$\pm$0.00
+QSAR_fish_toxicity -6%$\pm$411%$\pm$3-19%$\pm$50.76$\pm$0.03 0.23$\pm$0.01 4%$\pm$57%$\pm$40.55$\pm$0.04
+Table 13: Test results for all models with missing-raw data at 95% quantiles aggregated over five seeds.
+For each metric, the mean and standard deviation of the performance across the seeds are separated by $\pm$.
+Performance over the baseline is highlighted in bold.
+Interval Predictions Point Predictions
+Relative Absolute Relative Absolute
+Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
+MultiXGBs
+space_ga -5%$\pm$64%$\pm$2-12%$\pm$70.86$\pm$0.02 0.22$\pm$0.01 -7%$\pm$4-9%$\pm$30.45$\pm$0.06
+cpu_activity -5%$\pm$1912%$\pm$8-37%$\pm$460.89$\pm$0.06 0.14$\pm$0.04 -1%$\pm$21-17%$\pm$130.83$\pm$0.07
+naval_propulsion_plant 105%$\pm$29-16%$\pm$656%$\pm$60.70$\pm$0.06 0.18$\pm$0.02 -14%$\pm$20-29%$\pm$350.77$\pm$0.05
+miami_housing 16%$\pm$13-5%$\pm$510%$\pm$160.82$\pm$0.06 0.10$\pm$0.02 -14%$\pm$11-18%$\pm$100.72$\pm$0.10
+kin8nm -5%$\pm$23%$\pm$2-14%$\pm$50.88$\pm$0.01 0.48$\pm$0.01 -3%$\pm$3-4%$\pm$40.40$\pm$0.02
+concrete_compressive_strength 2%$\pm$38%$\pm$4-10%$\pm$30.77$\pm$0.03 0.30$\pm$0.01 -16%$\pm$12-20%$\pm$140.58$\pm$0.08
+cars -15%$\pm$1426%$\pm$9-63%$\pm$260.85$\pm$0.04 0.27$\pm$0.03 -30%$\pm$22-35%$\pm$240.66$\pm$0.15
+energy_efficiency 10%$\pm$2611%$\pm$6-24%$\pm$430.92$\pm$0.02 0.22$\pm$0.05 -186%$\pm$265-153%$\pm$1880.91$\pm$0.08
+california_housing 26%$\pm$7-2%$\pm$321%$\pm$60.88$\pm$0.03 0.40$\pm$0.03 -4%$\pm$4-5%$\pm$50.69$\pm$0.05
+airfoil_self_noise -2%$\pm$18-3%$\pm$4-4%$\pm$170.73$\pm$0.04 0.34$\pm$0.06 -37%$\pm$31-43%$\pm$360.41$\pm$0.12
+QSAR_fish_toxicity -2%$\pm$30%$\pm$5-6%$\pm$70.73$\pm$0.03 0.26$\pm$0.02 -4%$\pm$4-2%$\pm$30.36$\pm$0.04
+MultiETs
+space_ga 4%$\pm$3-5%$\pm$17%$\pm$40.88$\pm$0.02 0.24$\pm$0.01 -12%$\pm$4-14%$\pm$40.40$\pm$0.05
+cpu_activity -16%$\pm$10-5%$\pm$1-20%$\pm$180.93$\pm$0.01 0.20$\pm$0.03 -30%$\pm$23-54%$\pm$200.79$\pm$0.10
+naval_propulsion_plant 58%$\pm$5-17%$\pm$446%$\pm$30.82$\pm$0.05 0.41$\pm$0.03 -36%$\pm$21-93%$\pm$480.71$\pm$0.05
+miami_housing 20%$\pm$14-10%$\pm$421%$\pm$130.87$\pm$0.05 0.14$\pm$0.03 -27%$\pm$11-37%$\pm$110.67$\pm$0.11
+kin8nm 5%$\pm$2-8%$\pm$212%$\pm$40.87$\pm$0.02 0.50$\pm$0.02 -11%$\pm$2-13%$\pm$30.32$\pm$0.03
+concrete_compressive_strength 1%$\pm$5-14%$\pm$313%$\pm$20.80$\pm$0.02 0.37$\pm$0.01 -35%$\pm$21-50%$\pm$250.49$\pm$0.06
+cars -23%$\pm$6-8%$\pm$5-35%$\pm$140.83$\pm$0.04 0.24$\pm$0.02 -33%$\pm$18-31%$\pm$180.66$\pm$0.09
+energy_efficiency -15%$\pm$160%$\pm$9-36%$\pm$420.85$\pm$0.05 0.15$\pm$0.01 -73%$\pm$74-84%$\pm$640.94$\pm$0.04
+california_housing 21%$\pm$7-7%$\pm$222%$\pm$50.91$\pm$0.02 0.52$\pm$0.03 -14%$\pm$3-17%$\pm$30.61$\pm$0.05
+airfoil_self_noise -7%$\pm$18-20%$\pm$28%$\pm$160.76$\pm$0.03 0.42$\pm$0.08 -65%$\pm$54-93%$\pm$700.33$\pm$0.11
+QSAR_fish_toxicity 10%$\pm$7-9%$\pm$316%$\pm$30.82$\pm$0.03 0.33$\pm$0.03 -8%$\pm$8-10%$\pm$90.39$\pm$0.09
+MultiMLPs
+space_ga -11%$\pm$5-15%$\pm$21%$\pm$60.68$\pm$0.03 0.16$\pm$0.01 -40%$\pm$23-34%$\pm$150.21$\pm$0.23
+cpu_activity -50%$\pm$1622%$\pm$14-193%$\pm$1180.86$\pm$0.10 0.18$\pm$0.07 -15%$\pm$19-23%$\pm$170.75$\pm$0.12
+naval_propulsion_plant -52%$\pm$19-14%$\pm$24-102%$\pm$560.63$\pm$0.18 0.34$\pm$0.07 -108%$\pm$66-183%$\pm$195-0.03$\pm$0.47
+miami_housing -33%$\pm$1513%$\pm$10-92%$\pm$700.85$\pm$0.06 0.12$\pm$0.04 -25%$\pm$8-28%$\pm$110.67$\pm$0.10
+kin8nm -11%$\pm$5-5%$\pm$2-8%$\pm$60.77$\pm$0.01 0.35$\pm$0.02 -16%$\pm$8-15%$\pm$90.37$\pm$0.04
+concrete_compressive_strength -15%$\pm$13-13%$\pm$28-10%$\pm$480.60$\pm$0.22 0.22$\pm$0.11 -12%$\pm$5-12%$\pm$60.54$\pm$0.05
+cars -51%$\pm$443%$\pm$7-211%$\pm$340.93$\pm$0.02 0.29$\pm$0.03 -28%$\pm$12-26%$\pm$130.66$\pm$0.10
+energy_efficiency -13%$\pm$189%$\pm$15-30%$\pm$250.66$\pm$0.13 0.09$\pm$0.04 1%$\pm$9-1%$\pm$160.95$\pm$0.03
+california_housing -17%$\pm$64%$\pm$3-28%$\pm$140.88$\pm$0.03 0.44$\pm$0.04 -5%$\pm$2-3%$\pm$30.61$\pm$0.06
+airfoil_self_noise 2%$\pm$16-2%$\pm$5-1%$\pm$210.75$\pm$0.07 0.25$\pm$0.08 -15%$\pm$8-9%$\pm$80.55$\pm$0.18
+QSAR_fish_toxicity -17%$\pm$56%$\pm$6-30%$\pm$110.76$\pm$0.04 0.28$\pm$0.03 -13%$\pm$7-14%$\pm$80.34$\pm$0.06
+Appendix G Example of learning with non-normality
 To illustrate how SEMF adapts to non-normal outcomes, we provide an example from the naval_propulsion_plant dataset [57].

### v4

@@
 Abstract
-This work introduces the Supervised Expectation-Maximization Framework (SEMF), a versatile and model-agnostic approach for generating prediction intervals in datasets with complete or missing data.
+This work introduces the Supervised Expectation-Maximization Framework (SEMF), a versatile and model-agnostic approach for generating prediction intervals with any ML model.
 SEMF extends the Expectation-Maximization algorithm, traditionally used in unsupervised learning, to a supervised context, leveraging latent variable modeling for uncertainty estimation.
-Extensive empirical evaluations across 11 tabular datasets show that SEMF often achieves narrower normalized prediction intervals and higher coverage rates than traditional quantile regression methods [^1] Code: https://github.com/Unco3892/semf-paper.
-Furthermore, SEMF can be integrated with machine learning models like gradient-boosted trees and neural networks, highlighting its practical applicability.
-The results indicate that SEMF enhances uncertainty quantification, particularly in scenarios with complete data.
+Through extensive empirical evaluation of diverse simulated distributions and 11 real-world tabular datasets, SEMF consistently produces narrower prediction intervals while maintaining the desired coverage probability, outperforming traditional quantile regression methods.
+Furthermore, without using the quantile (pinball) loss, SEMF allows point predictors, including gradient-boosted trees and neural networks, to be calibrated with conformal quantile regression.
+The results indicate that SEMF enhances uncertainty quantification under diverse data distributions and is particularly effective for models that otherwise struggle with inherent uncertainty representation.
+[^1] Code: https://github.com/Unco3892/semf-paper
 1 Introduction
@@
 In this paper, we introduce the Supervised Expectation-Maximization Framework (SEMF) based on the Expectation-Maximization (EM) algorithm [5].
-Traditionally recognized as a clustering technique, EM is used for supervised learning in SEMF, allowing for both point estimates and prediction intervals using any ML model (model-agnostic).
-SEMF generates representations for latent or missing modalities, which can be relevant for incomplete data and holds potential for multi-modal data applications, though multi-modal settings are left for future exploration.
+Traditionally recognized as a clustering technique, EM is used for supervised learning in SEMF, allowing for both point estimates and prediction intervals using any ML model (model-agnostic), with a particular focus on uncertainty quantification.
 This paper details the methodology behind the framework and proposes a training algorithm based on Monte Carlo (MC) sampling, also used in variational inference for Variational Auto-Encoders (VAEs) [6, 7].
-SEMF differs from prominent supervised EM approaches such as [8], which focus on point prediction using Gaussian Mixture Models (GMMs).
-Additionally, our method operates in a frequentist paradigm, directly maximizing the likelihood function through iterative EM steps without integrating over posterior distributions.
+The EM algorithm has already been used for supervised learning tasks using specific models [8, 9, 10], minimal supervision [11], and using labeled and unlabeled data in semi-supervised settings for both single and multiple modalities [12, 13].
+Our method, SEMF, differs from prominent supervised EM approaches such as ghahramani1993em [8], which focus on point prediction using Gaussian Mixture Models (GMMs).
+Furthermore, SEMF operates in a frequentist paradigm, directly maximizing the likelihood function through iterative EM steps without integrating over posterior distributions.
 Although SEMF can be extended to a Bayesian framework as its likelihood component, this extension lies beyond the scope of this paper.
-The remainder of this paper is organized as follows: Section 2 details the theory and the methodology of SEMF.
-Section 3 reviews related works in latent representation learning, uncertainty estimation, and handling of missing data.
-Section 4 describes the experimental setup, including datasets and evaluation metrics.
+Finally, SEMF generates representations for latent modalities through specialized models, holding potential for multi-modal data applications, though multi-modal settings are left for future exploration.
+The remainder of this paper is organized as follows: Section 2 details the theory and methodology of SEMF.
+Section 3 reviews related work in latent representation learning and uncertainty estimation.
+Section 4 describes the experimental setup, including synthetic and benchmark datasets, as well as evaluation metrics.
 Section 5 discusses the results, demonstrating the efficacy of SEMF.
@@
 This section presents the founding principles of SEMF from its parameters, training, and inference procedure with, at its core, the EM algorithm.
-This algorithm, first introduced by [5], is an unsupervised method for handling latent variables and incomplete data.
 Invented to maximize the model likelihood, it builds a sequence of parameters that guarantee an increase in the log-likelihood [14] by iterating between the Expectation (E) and the Maximization (M) steps.
@@
 where $p^{\prime}$ stands for the current estimates, $\log p(x,z)$ is the log-likelihood of the complete observation $(x,z)$, and $z$ is a latent variable.
-The M-step maximizes this $Q$-function: $p^{\prime}\leftarrow\arg\max_{p}Q(p|p^{\prime})$ The sequence is repeated until convergence.
+The M-step updates the current estimates with the arguments maximizing the $Q$-function: $p^{\prime}\leftarrow\arg\max_{p}Q(p|p^{\prime})$ The sequence is repeated until convergence.
 2.1 Problem Scenario
 Let $x=(x_{1},x_{2},\ldots,x_{K})$ denote $K$ inputs and the output be $y$.
-For simplicity, we limit $y$ to be numerical, although it could be categorical without loss of generality.
+For simplicity, we limit $y$ to be continuous numerical, although this assumption could be relaxed to discrete or categorical without loss of generality.
 Component $x_{k}$ is a source: a modality, a single or group of variables, or an unstructured input such as an image or text.
 For clarity, we limit to $K=2$, where $x_{1}$ and $x_{2}$ are single variables.
-We assume that only $x_{1}$ may contain missing values, either at random or partially at random.
 […]
+The contribution to the log-likelihood of a complete observation $(y,z,x)$ is $\log p(y,z|x)=\log p(y|z)+\log p(z|x)$. In the E-step, we compute
 […]
-where $p^{\prime}$ is our current estimate.
 Eq.
@@
 […]
-If $x_{1}$ is missing, a similar development leads to
-[…]
-where $x_{1,r}$ and $z_{r}$ are respectively sampled from $p^{\prime}(x_{1}|x_{2})$ and $p^{\prime}(z|x_{1,r},x_{2})$.
 2.2 Objective Function
-Adapting Eq. 3 and Eq. 4 for the observed data $\{(y_{i},x_{i})\}_{i=1}^{N}$, the overall loss function, $\mathcal{L}$, is
-$\mathcal{L}(\phi,\theta,\xi)=-\sum_{i=1}^{N}\sum_{r=1}^{R}\left\{\log p_{\phi}(z_{i,r}|x_{i,r})+\log p_{\theta}(y_{i}|z_{i,r})+\mathds{1}_{\{i\in I_{m}\}}\log p_{\xi}(x_{1,i,r}|x_{2,i})\right\}w_{i,r},$ (5)
-where $I_{m}$ is the set of those $i$’s such that $x_{1,i}$ is missing.
-The models of $p(y|z)$, $p(z|x)$, and $p(x_{1}|x_{2})$ inherit parameters $\theta$, $\phi$, and $\xi$, respectively.
-Also, $x_{1,i,r}$ is sampled from $p_{\xi^{\prime}}(x_{1}|x_{2,i})$, if $x_{1,i}$ is missing, and $z_{1,i,r}$ and $z_{2,i,r}$ are sampled from $p_{\phi^{\prime}}(z_{1}|x_{1,i,r})$ and $p_{\phi^{\prime}}(z_{2}|x_{2,i})$.
-Furthermore, for compactness of notation, $x_{i,r}$ is $(x_{1,i},x_{2,i})$ if $x_{1,i}$ is observed, and $(x_{1,i,r},x_{2,i})$ if $x_{1,i,r}$ is missing.
-Finally, the weights are
-$w_{i,r}=\frac{p_{\theta^{\prime}}(y_{i}|z_{i,r})}{\sum_{t=1}^{R}p_{\theta^{\prime}}(y_{i}|z_{i,t})}.$ (6)
+Adapting Eq. 3 for the observed data $\{(y_{i},x_{i})\}_{i=1}^{N}$, the overall loss function, $\mathcal{L}$, is
+$\mathcal{L}(\phi,\theta)=-\sum_{i=1}^{N}\sum_{r=1}^{R}\left\{\log p_{\phi}(z_{i,r}|x_{i})+\log p_{\theta}(y_{i}|z_{i,r})\right\}w_{i,r},$ (4)
+where the models of $p(y|z)$ and $p(z|x)$ inherit parameters $\theta$ and $\phi$, respectively. The weights are
+$w_{i,r}=\frac{p_{\theta^{\prime}}(y_{i}|z_{i,r})}{\sum_{t=1}^{R}p_{\theta^{\prime}}(y_{i}|z_{i,t})}$ (5)
+where $z_{i},r\stackrel{{\scriptstyle ind.}}{{\sim}}p_{\phi^{\prime}}(z|x_{i})$, $r=1,\ldots,R$.
 Eq.
-5 shows that ${\cal L}$ is a sum of losses associated with the encoder model, $p_{\phi}$, for each source, the decoder model, $p_{\theta}$, from the latent variables to the output, and, if applicable, the model handling missing data, $p_{\xi}$.
-At each M-step, ${\cal L}$ is minimized with respect to $\theta$, $\phi$, and $\xi$.
-Then, $\theta^{\prime}$, $\phi^{\prime}$, and $\xi^{\prime}$ are updated, as well as the weights and the sampling.
-Then the process is iterated until convergence.
-2.2.1 Example: $\mathcal{L}(\phi,\theta,\xi)$ under normality
-Similar to [7], we develop further ${\cal L}$ under normality assumptions for the encoder, $p_{\phi}$, and the decoder, $p_{\theta}$.
-These simple cases are illustrative, though any other distributions could be adopted, including non-continuous or non-numerical outcomes.
+4 shows that ${\cal L}$ is a sum of losses associated with the encoder model, $p_{\phi}$, for each source, and the decoder model, $p_{\theta}$, from the latent variables to the output.
+$p_{\theta}$ and $p_{\phi}$ are referred to as encoder and decoder in the spirit of auto-encoder models.
+At each M-step, ${\cal L}$ is minimized with respect to $\theta$ and $\phi$.
+Then, $\theta^{\prime}$ and $\phi^{\prime}$ are updated, as well as the weights and the sampled $z$.
+Then, the process is iterated until convergence.
+2.2.1 Example: $\mathcal{L}(\phi,\theta)$ under normality
+For illustration purposes, we develop below the case where $p_{\phi}$ and $p_{\theta}$ are normal distributions, similar to kingma2014vae [7].
+As a reminder, any other distributions could be adopted, including non-continuous or non-numerical outcomes.
 Encoder $p_{\phi}(z|x)$.
@@
 where $J_{m_{k}}$ is the $m_{k}\times m_{k}$ identity matrix. In particular,
-$\log p_{\phi}(z_{k}|x_{k})$ $=$ $-\frac{m_{k}}{2}\log 2\pi-\frac{1}{2}\log\sigma_{k}^{2}-\frac{1}{2\sigma_{k}^{2}}\sum_{j=1}^{m_{k}}\{z_{k,j}-g_{\phi_{k},j}(x_{k})\}^{2},\quad k=1,2.$ (8)
+$\log p_{\phi}(z_{k}|x_{k})$ $=$ $-\frac{m_{k}}{2}\log 2\pi-\frac{m_{k}}{2}\log\sigma_{k}^{2}-\frac{1}{2\sigma_{k}^{2}}\sum_{j=1}^{m_{k}}\{z_{k,j}-g_{\phi_{k},j}(x_{k})\}^{2},\quad k=1,2.$ (7)
 The mean $g_{\phi_{k}}(x_{k})$ can be any model, such as a neural network, with output of length $m_{k}$, $k=1,2$.
@@
 It controls the amount of noise introduced in the latent dimension and is pivotal in determining the prediction interval width for $p(y|z)$.
-In this paper, $\sigma_{k}$ is fixed for simplicity.
+The code implementation allows for all three methods.
+This paper only presents the results for a fixed $\sigma_{k}$, as well as training a separate set of $g_{\phi_{k}}(x_{k})$ models to estimate varying $\sigma_{k}$s.
 Decoder $p_{\theta}(y|z)$.
@@
 Again, the mean $f_{\theta}(z)$ can be any model, such as a neural network.
-Model for missing data $p_{\xi}(x_{1}|x_{2})$.
-We use an empirical model for $X_{1}$ given $X_{2}=x_{2}$ where $p_{\xi}(x_{1}|x_{2})$ put masses only on those non-missing $x_{1}$’s in the training set.
-Let $I_{nm}$ be the set of indices such that $x_{1,j}$, $j\in I_{nm}$, are all the non-missing $x_{1}$ in the training set.
-Additionally, for a given $j\in I_{nm}$, $x_{1}[j]$ is the observed $x_{1}$ corresponding to $j$.
-For a given $x_{2}$, $p_{\xi}(x_{2})$ is a vector of length $|I_{nm}|$, the cardinality of $I_{nm}$, with components
-$p_{\xi}(j|x_{2})=\frac{\exp\{h_{\xi,j}(x_{2})\}}{\sum_{t\in I_{nm}}\exp\{h_{\xi,t}(x_{2})\}},\quad j\in I_{nm},$ (11)
-where $h_{\xi}$ is a vector of length $|I_{nm}|$, typically a neural network with input $x_{2}$ and output on $I_{nm}$. Now, the probability of $X_{1}=x_{1}$ given $X_{2}=x_{2}$ is
-$p_{\xi}(x_{1}|x_{2})=\sum_{j\in I_{nm}}p_{\xi}(j|x_{2})\cdot\mathds{1}\{x_{1}=x_{1}^{(nm)}[j]\}.$ (12)
 Summary.
@@
 $(\sigma^{*})^{2}=\frac{1}{N}\sum_{i,r}w_{i,r}\{y_{i}-f_{\theta^{*}}(z_{i,r})\}^{2},$ (12)
-$\xi^{*}=\arg\max_{\xi}\sum_{t\in D_{x}}w_{t}\log p_{\xi}(j_{t}|x_{2,t}).$ (16)
-When $x_{1,i}$ is missing, the sampling is enriched by simulated $x_{1,i,r}$.
-Eq.
-16 selects $j^{*}$ from $I_{nm}$ based on $p_{\xi^{\prime}}(j|x_{2})$ according to Eq.
-12 and Eq.
-11.
-The $D_{x}$ above is a subset of the data for $\xi$, learned on the missing data part.
-We also note that learning the parameters above is parallelizable.
 2.3 Training
@@
 The process iterates for each batch until the maximum number of steps is reached or an early stopping criterion is satisfied.
-The full details are given in Algorithm 1 (Appendix A).
+The full details are given in Algorithm 1.
 The framework requires tuning hyper-parameters such as the number of MC samples $R$, the number of latent nodes $m_{k}$, and the standard deviation $\sigma_{k}$ of $Z_{k}$.
@@
 Moreover, due to the generative nature of SEMF, the variation resulting from the initial random seed is measured in Subsection 4.1.
-Additionally.
-the model-specific hyper-parameters ($p_{\phi}$, $p_{\theta}$ and $p_{\xi}$) are also discussed in the same Subsection.
+Additionally, the model-specific hyper-parameters ($p_{\phi}$ and $p_{\theta}$) are also discussed in the same subsection.
+Algorithm 1 SEMF Training: two input sources
+1:$y,x_{1},x_{2},R$
+$\theta,\phi_{1},\phi_{2}$
+3:Initialize $\theta,\phi_{1},\phi_{2}$
+4:Initialize $D_{y},D_{z_{1}},D_{z_{2}},D_{x},D_{i}$ to $\emptyset$
+5:Split $I=\{1,\ldots,N\}$ into $L$ batches $\{b_{1},\ldots,b_{L}\}$
+6:for $\ell=1$, …, $L$ do
+7:   for all $i$ in $b_{\ell}$ do
+8:      for $r=1$, …, $R$ do
+9:         Simulate $z_{1,i,r}\sim p_{\phi_{1}}(\cdot|x_{1,i})$
+10:         Simulate $z_{2,i,r}\sim p_{\phi_{2}}(\cdot|x_{2,i})$
+11:         Set $z_{i,r}=[z_{1,i,r},z_{2,i,r}]$
+12:      end for
+13:      for $r=1$, …, $R$ do
+14:         Compute
+$w_{i,r}=\frac{p_{\theta}(y_{i}|z_{i,r})}{\sum_{r=1}^{R}p_{\theta}(y_{i}|z_{i,r})}$
+15:         Update $D_{y}\leftarrow D_{y}\cup[y_{i}|z_{i,r}|w_{i,r}]$
+16:         Update $D_{z_{2}}\leftarrow D_{z_{2}}\cup[z_{2,i,r}|x_{2,i}|w_{i,r}]$
+17:         Update $D_{z_{1}}\leftarrow D_{z_{1}}\cup[z_{1,i,r}|x_{1,i}|w_{i,r}]$
+18:      end for
+19:      Update $\theta\leftarrow Q_{y}(\theta,D_{y})$
+20:      Update $\phi_{1}\leftarrow Q_{1}(\phi_{1},D_{z_{1}})$
+21:      Update $\phi_{2}\leftarrow Q_{2}(\phi_{2},D_{z_{2}})$
+22:   end for
+23:end for
+24:Check convergence; Go to step 4 if not
 2.4 Inference
 The encoder-decoder structure of SEMF entails the simulations of $z_{r}$ during inference, as depicted in Figure 1.
-In theory, any inference can be performed for $\hat{y}$, for instance the mean value $\hat{y}=\frac{1}{R}\sum_{r=1}^{R}f_{\theta}(z_{r})$, where $z_{r}\sim p_{\phi}(z|x)$ (see Algorithm 2 in Appendix A).
-For prediction intervals, a double simulation scheme is used,
+In theory, any inference on $y$ given $x$ can be performed for $\hat{y}$, for instance the mean value $\hat{y}=\frac{1}{R}\sum_{r=1}^{R}f_{\theta}(z_{r})$, where $z_{r}\sim p_{\phi}(z|x)$ (see Algorithm 2 for the simulation scheme).
+For prediction intervals, a second simulation step is used,
 $z_{r}\sim p(z|x),\quad\hat{y}_{r,s}\sim p_{\theta}(y|z_{r}),\quad r,s=1,\ldots,R.$ (13)
-Prediction interval at a given level of certainty $\alpha$ follows as
+Prediction interval on $y$ given $x$ at a given level $\alpha$ follows as
 $PI=\text{quantile}\left(\{\hat{y}_{r,s}\};\frac{\alpha}{2},1-\frac{\alpha}{2}\right).$ (14)
-Remark.
-We denote the $R$ for inference as $R_{\mathrm{infer}}$.
-Figure 1: Inference procedure with the SEMF’s learnable parameters $\phi_{k},\theta$ and $\xi$.
-Here, we illustrate the number of inputs $k$ as $k=1,2$, assuming that $x_{1}$ may contain missing values
+In order to calibrate these intervals, we can use conformal prediction [15], which we review in Subsection 3.2 and subsequently explain in Subsection 4.1.
+Figure 1: Inference procedure with the SEMF’s learnable parameters $\phi_{k}$ and $\theta$. Here, we illustrate the number of input sources $k$ as $k=1,2$
+Algorithm 2 SEMF Simulation
+1:$\theta^{*},\phi_{1}^{*},\phi_{2}^{*},x_{1},x_{2},R$
+2:$z_{i,r}$
+3:for $r=1,\ldots,R$ do
+4:   Simulate $z_{1,i,r}\sim p_{\phi_{1}^{*}}(\cdot|x_{1,i})$
+5:   Simulate $z_{2,i,r}\sim p_{\phi_{2}^{*}}(\cdot|x_{2,i})$
+6:   Set $z_{i,r}=[z_{1,i,r},z_{2,i,r}]$
+7:end for
 3 Related Work
 3.1 Latent Representation Learning
-Latent representation learning involves modeling hidden variables from observed data for various ML tasks, most notably Auto-Encoders (AEs) and VAEs.
-AEs are neural networks that reconstruct inputs by learning an intermediate latent representation.
-The encoder $g_{\phi}(\cdot)$ in an AE embeds the input $x$ into a latent variable $z$, which then passes through the decoder $f_{\theta}(\cdot)$, reconstructing $x$ as $\hat{x}=f_{\theta}(z)$.
-The training process optimizes the parameters $\phi$ and $\theta$ by minimizing the reconstruction loss.
-Unlike AEs, which focus on reconstruction, VAEs use variational methods to fit distributions of latent variables and the output [7].
-Given a sample $x$ with a latent variable $z$, VAEs model the marginal likelihood $p_{\theta}(x)$.
+Latent-representation learning typically relies on the encoder-decoder architecture introduced in Section 2.
+Classic examples are Auto-Encoders (AEs) and VAEs: the encoder $g_{\phi}$ maps a sample input $x$ to a latent variable $z$, and the decoder $f_{\theta}$ reconstructs the input $\hat{x}=f_{\theta}(z)$.
+Training jointly optimizes the parameter vectors $\phi$ and $\theta$ by minimizing the reconstruction loss.
+Unlike AEs, which focus solely on reconstruction, VAEs introduce a variational objective that aims to model the data distribution through the marginal likelihood $p_{\theta}(x)$ while fitting approximate posterior distributions over the latent variables [7, 16].
 However, directly maximizing this likelihood is difficult [16].
@@
 3.2 Prediction Intervals
-Crucial for estimating uncertainty, prediction intervals in regression are often derived using methods such as Bayesian approaches [22, 23, 24], ensemble techniques [25, 26, 27], or quantile regression [28, 29].
-Additionally, these methods can be complemented with conformal prediction, a framework for calibrating any point predictor to produce prediction intervals [30, 31], making it highly relevant for enhancing reliability in applications requiring rigorous uncertainty quantification.
-A key component of quantile regression is the pinball loss function, which effectively balances the residuals to capture the desired quantiles even for non-parametric models [32], making it ideal for asymmetric distributions where tail behavior is of critical importance [29].
-This loss function is pivotal not only for single model scenarios but also enhances ensemble methods by refining their quantile estimate [33].
-Conformal prediction further extends the applicability of these intervals by providing a layer of calibration that adjusts intervals obtained from any predictive model, ensuring they cover the true value with a pre-specified probability [34].
-3.3 Missing Data
-Managing missing values is a pivotal aspect when dealing with real-world data.
-Naive methods such as discarding instances or mean/median-imputation may be infeasible or carry the risk of changing the data distribution [35, 36].
-The chosen technique for handling missing data should adhere to the dataset’s characteristics and mechanisms behind the missing data [37].
-More advanced approaches, like the Iterative Imputer from scikit-learn [38], an implementation of Multiple Imputation by Chained Equations (MICE) [39], expand on the simple imputations by iteratively modeling each feature with missing values as a function of other features [40, 41].
-Due to its complexity, MICE is best compared with alternatives such as K-means clustering and artificial neural networks (ANNs).
-K-means assigns missing values based on cluster centroids [42] as opposed to ANNs, which learn complex, non-linear relationships between variables [43].
-ANNs effectively predict or reconstruct missing values and thus are particularly useful in datasets where relationships between variables are intricate and not easily captured by straightforward imputation methods.
-Accordingly, ANNs have demonstrated superior performance over MICE and GMM (with EM for missing values) in scenarios where a large proportion of the data is missing [44].
+Crucial for estimating uncertainty, prediction intervals provide a range for regression outcomes.
+Common approaches include Bayesian methods [23, 24, 25], ensemble techniques [26, 27, 28], and quantile regression [29, 30].
+Quantile regression specifically utilizes the pinball loss function to target desired quantiles, proving effective even for non-parametric models [31] and asymmetric distributions [30].
+This loss is valuable both for individual models and for refining quantile estimates when used with ensembles [32]—a loss that can also be learned jointly as demonstrated by Simultaneous Quantile Regression (SQR) [33].
+SQR represents a powerful neural network approach that jointly estimates multiple quantiles while maintaining their proper ordering, making it particularly effective for constructing prediction intervals and a strong baseline for uncertainty quantification tasks.
+Complementary to these approaches, conformal prediction (CP) offers a general framework to construct and calibrate prediction intervals for any model [34].
+Assuming data exchangeability, CP adjusts intervals to ensure they meet a pre-specified coverage probability, thereby enhancing reliability in applications demanding rigorous uncertainty quantification.
+Recent advances in CP include its integration with ensemble models for time series [35, 36] and Bayesian approaches such as Conformal Bayesian model averaging (CBMA) aggregates conformity scores from several Bayesian models to obtain optimal prediction sets with guarantees [37].
+Furthermore, several CP frameworks incorporate data-dependent score functions have recently been proposed to improve the efficiency and validity of intervals derived via quantile estimators [38], boosting methods (XGBoost) to refine conformity scores [39] and locally adaptive techniques to enhance conditional validity [40].
 4 Experimental Setup
-4.1 Datasets
-We systematically curate a subset of datasets from the OpenML-CTR23 [45] benchmark suite to evaluate and carry out our experiments.
-Initially comprising 35 datasets, we apply an exclusion criteria to refine this collection to 11 datasets.
-The details and overview are in Appendix B.
-We remove duplicated rows from all the datasets and carry out the standardization (scaling) of all predictors, including the outcome, which we transformed to have zero means and unit variances.
-The features of these datasets are then treated as separate inputs to SEMF.
-In the second stage of our experiments, we artificially introduce 50% missing values in our datasets for any predictor except for the first feature of a randomly chosen row, which emulates missing completely at random (MCAR) data.
-In all our datasets, 70% of the data is used to train all models, 15% as a hold-out validation set to monitor SEMF’s performance, and 15% to evaluate the models.
-To combat overfitting, baseline models that benefit from early stopping are allocated another 15% from the training data.
-Lastly, it is essential to note that all data in SEMF are processed batch-wise, without employing mini-batch training, to ensure consistency and stability in the training process.
 4.1 Models
-Our baseline consists of both point and quantile regression eXtreme Gradient Boosting (XGBoost) [46], Extremely Randomized Trees (ET) [47], and neural networks [48], all summarized and depicted in Table 1.
+Our baseline consists of quantile regression models based on eXtreme Gradient Boosting (XGBoost) [41], Extremely Randomized Trees (ET) [42], and neural networks using SQR [33], all summarized and depicted in Table 1.
 To ensure consistency in our experimental setup, we align the families and hyper-parameters of $p_{\phi}$ and $p_{\theta}$ with our baseline models.
@@
 When establishing prediction intervals, we conformalize our prediction intervals according to [43] at an uncertainty tolerance of 5% for both the baseline and SEMF (Eq.
-18).
-The missing data simulator, $p_{\xi}$, is constructed using a shallow neural network, which employs the $\mathrm{SELU}$ activation function [49].
-It is experimented with two distinct node counts: 50 and 100.
-To constrain the breadth of our parameter exploration, the simulator for the missing model adopts the optimal set of hyper-parameters identified from analyses involving complete datasets.
-We target (larger) $\sigma_{k}$ values that introduce more noise and produce better intervals than point predictions.
-The optimal models are then trained and tested with five different seeds, and the results are averaged.
+14).
+On the notable hyper-parameters, we target larger $\sigma_{k}$ values that introduce more noise and produce better intervals than point predictions.
+For several datasets, we set $\sigma_{k}$ to train_residual_models, which trains separate models to predict the scale of the latent variables individually and is more suitable to heteroscedastic noise patterns.
+This approach often yields substantially improved interval predictions by capturing local uncertainty structures, though at the cost of longer computation times due to the additional model training.
+In our experience, using train_residual_models for $\sigma_{k}$ is recommended as a starting point for new applications, with fixed values being a computationally efficient alternative once the behavior of the dataset is better understood.
+The optimal models are then trained and tested with five different seeds, and the results are averaged, and the variability of the results is presented.
 The point prediction, $\hat{y}$, uses the mean inferred values.
-We then study the performance of SEMF against mean and median imputation techniques, five nearest neighbors, and MICE from [38].
-The imputers are used within the point and interval baseline models explained in the following subsection to form the missing value baseline.
-Appendix C contains more details on the hyper-parameters for each SEMF model and dataset.
-Table 1: SEMF models, baselines, and hyper-parameters.
-SEMF Point Prediction Baseline Interval Prediction Baseline
-MultiXGBs XGBoost [50]
+Appendix A contains more details on the hyper-parameters for each SEMF model and dataset introduced below.
+SEMF Base Model Interval Prediction Baseline
+MultiXGBs XGBoost
 Trees: 100, Maximum depth: 6, Early stopping steps: 10
@@
 Same as point prediction baseline, XGBoost
-MultiETs Extremely Randomized Trees [38]
+MultiETs Extremely Randomized Trees
 Trees: 100, Maximum depth: 10
-Quantile Extremely Randomized Trees [51]
+Quantile Extremely Randomized Trees
 Same as point prediction baseline, Extremely Randomized Trees
@@
 Hidden layers: 2, Nodes per layer: 100, Activation functions: ReLU, Epochs: 1000 or 5000, Learning rate: 0.001, Batch training, Early stopping steps: 100
-Simultaneous Quantile Regression [48]
+Simultaneous Quantile Regression
 Same as point prediction baseline, Deep Neural Network
+Table 1: SEMF models, baselines, and hyper-parameters.
+4.2 Datasets
+4.2.1 Simulations
+In our experiments, we generate synthetic datasets with $10^{3}$ observations and $k=2$ predictors from a standard normal distribution according to
+$y=f(x)+\epsilon,$ (15)
+where $f(x)$ in a simple setup is defined as
+$f(x)=\sum_{i=1}^{k}\cos(x_{i})$ (16)
+to isolate noise effects, and in a more complex setup, we instead generate
+$f(x)=\sum_{i=1}^{k}\Bigl{(}x_{i}^{2}+0.5\,\sin(3\,x_{i})\Bigr{)}$ (17)
+to introduce nonlinearity and heteroscedasticity.
+$\epsilon$ is drawn from one of four distributions: $\mathcal{N}(0,0.5)$, $\mathcal{U}(-0.5,0.5)$, a centered log-normal, or $\text{Gumbel}(0,0.5)$.
+Figure 2 illustrates an example of Eq.
+17 with $k=1$ (for simplicity) and 500 observations, where blue dots indicate predictions with 95% prediction intervals and the red curve denotes $f(x)$.
+For our simulation experiments, the underlying data is generated from a fixed seed with variability only in the model seeds (Subsection 4.1).
+We evaluate all three model variants (MultiXGBs, MultiETs, and MultiMLPs) against standard quantile regression baselines.
+For MultiXGBs and MultiETs, we use $R=10$ with 10 nodes per latent dimension, while MultiMLPs use $R=5$ with 20 nodes per latent dimension.
+All models employ early stopping with appropriate patience settings to prevent overfitting on the synthetic data.
+The data scaling and splits are the same as the benchmark data below.
+Figure 2: Prediction intervals under different noise distributions for a one-dimensional $x$ generated from 500 observations.
+Each panel shows predictions (blue dots) with 95% prediction intervals (shaded regions) for a model trained on data with (a) Normal, (b) Uniform, (c) Log-normal, and (d) Gumbel noise.
+The red curve denotes $f(x)$, the underlying (deterministic) function.
+4.2.2 Benchmark data
+We systematically curate a subset of datasets from the OpenML-CTR23 [44] benchmark suite to evaluate and carry out our experiments.
+Initially comprising 35 datasets, we apply an exclusion criteria to refine this collection to 11 datasets.
+The details and overview are in Appendix B.
+We remove duplicated rows from all the datasets and carry out the scaling of all predictors, including the outcome.
+The features of these datasets are then treated as separate inputs to SEMF.
+In all our datasets, 70% of the data is used to train all models, 15% as a hold-out validation set to monitor SEMF’s performance, and 15% to evaluate the models.
+To combat overfitting, baseline models that benefit from early stopping are allocated another 15% from the training data.
+Lastly, it is essential to note that all data in SEMF are processed batch-wise, without employing mini-batch training, to ensure consistency and stability in the training process.
 4.3 Metrics
-The evaluation of point predictions employs Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and R-squared ($\text{R}^{2}$).
-For prediction intervals, the chosen metrics, following [52] and [53], are the Prediction Interval Coverage Probability (PICP), and Normalized Mean Prediction Interval Width (NMPIW) as described in Subsection D.1.
-This paper also evaluates SEMF on our new metric, termed Coverage-Width Ratio (CWR),
+The evaluation of prediction intervals employs Prediction Interval Coverage Probability (PICP), Normalized Mean Prediction Interval Width (NMPIW), Continuous Ranked Probability Score (CRPS) and the quantile (pinball) loss.
+Detailed definitions and closed-form expressions for these metrics (with the CRPS computed under a uniformity assumption on the predictive distribution) are provided in Appendix C.
+In addition, we introduce the Coverage-Width Ratio (CWR)
 $\text{CWR}=\frac{\text{PICP}}{\text{NMPIW}},$ (18)
-which evaluates the coverage probability ratio to the prediction interval’s width.
-CWR provides a refined understanding of the balance between an interval’s accuracy (coverage) and precision (width).
-Though a larger value of this metric is better in higher confidence levels, the marginal increase in NMPIW is likely higher than that of PICP, resulting in decreasing CWR.
-In our case, measuring the performance of SEMF over the baseline models is far more critical than reviewing absolute metrics in isolation.
-For any metric above, except for ($\text{R}^{2}$), this is computed as
+which quantifies the trade-off between reliability and precision.
+In our case, measuring the performance of SEMF over the baseline models is far more critical than reviewing absolute metrics in isolation. For any metric above, this is computed as
 $\text{Metric}_{\Delta}(\%)=\left(\frac{\text{Metric}_{\text{SEMF}}-\text{Metric}_{\text{Baseline}}}{\text{Metric}_{\text{Baseline}}}\right)\times 100,$ (19)
-on which we base our decisions for selecting the best hyper-parameters as explained in Subsection D.2.
-5 Results
-We trained and tested 330 models corresponding to the three model types—MultiXGBs, MultiETs, and MultiMLPs—across 11 datasets with both complete and 50% missing data, using five seeds for each combination.
-Table 2 and Table 3 present the mean and standard deviation for our metrics aggregated over the five seeds.
+on which we base our decisions for selecting the best hyper-parameters as explained in Subsection C.2.
+5 Experiments
+5.1 Simulated Data
+Table 2 and Table 3 summarize the performance of our SEMF variants on synthetic datasets generated using two predictors.
+The PICP shows that on average, all models across all experiments were able to achieve at least the 95% desired coverage probability.
+$\Delta$CWR $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball $\Delta$PICP PICP
+MultiXGBs
+normal 18% [10:35] 16% [10:28] 6% [3:10] 11% [5:17] -1% [-3:0] 0.95$\pm$0.01
+uniform 25% [-2:58] 19% [-5:39] 10% [-1:23] 18% [0:32] -1% [-5:2] 0.96$\pm$0.03
+lognormal 28% [8:59] 21% [8:39] 13% [8:18] 7% [2:14] -1% [-3:0] 0.96$\pm$0.01
+gumbel 10% [-5:20] 8% [-9:18] 8% [1:14] 8% [0:16] 0% [-3:4] 0.96$\pm$0.01
+MultiETs
+normal 15% [1:32] 14% [0:27] 5% [0:8] 8% [0:16] -2% [-4:1] 0.95$\pm$0.01
+uniform 16% [3:35] 13% [-1:28] 6% [3:10] 13% [8:21] 0% [-3:4] 0.97$\pm$0.01
+lognormal 5% [-14:18] 3% [-18:17] 9% [5:13] -3% [-10:2] 0% [-2:2] 0.97$\pm$0.01
+gumbel 3% [-4:14] 3% [-7:14] 8% [6:11] 1% [-2:5] 0% [-2:2] 0.96$\pm$0.01
+MultiMLPs
+normal 3% [-2:9] 3% [-2:8] 1% [-2:4] 2% [-3:5] 0% [-1:0] 0.96$\pm$0.01
+uniform 4% [-2:13] 4% [-2:10] 2% [-1:8] 7% [-5:23] 0% [-4:3] 0.97$\pm$0.02
+lognormal 7% [-4:18] 6% [-5:16] 3% [-2:7] 4% [0:9] 0% [-1:1] 0.97$\pm$0.01
+gumbel 5% [-3:13] 5% [-5:14] 0% [-2:3] 0% [-4:4] -1% [-2:2] 0.96$\pm$0.02
+Table 2: Test results (95% prediction intervals) for 1000 observations generated using cosine $f(x)$ with 2 predictors, and an additive noise ($\epsilon$) belonging to one of the four distributions.
+Relative metrics are shown over 5 seeds as ‘mean [min:max]’, and absolute metrics as ‘mean $\pm$ std’.
+Average performance over the baseline is highlighted in bold.
+$\Delta$CWR $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball $\Delta$PICP PICP
+MultiXGBs
+normal 39% [28:65] 28% [22:41] 29% [26:34] 23% [16:35] 0% [-3:0] 0.95$\pm$0.01
+uniform 65% [54:80] 39% [34:44] 41% [39:47] 25% [16:32] 0% [-1:1] 0.96$\pm$0.02
+lognormal 0% [-13:7] -2% [-16:6] 20% [14:27] 4% [-4:16] 2% [0:5] 0.96$\pm$0.02
+gumbel 23% [0:49] 17% [0:35] 20% [9:30] 16% [3:21] 0% [-3:4] 0.96$\pm$0.01
+MultiETs
+normal 23% [7:35] 19% [7:26] 14% [0:21] 12% [-4:23] -1% [-2:1] 0.95$\pm$0.02
+uniform 40% [25:71] 28% [20:42] 24% [16:33] 12% [-4:21] 0% [-3:2] 0.97$\pm$0.02
+lognormal 8% [-9:33] 5% [-13:27] 11% [0:16] 6% [-6:12] 1% [-2:3] 0.97$\pm$0.01
+gumbel 29% [-3:68] 19% [-4:42] 18% [5:30] 15% [0:28] 0% [-2:1] 0.97$\pm$0.01
+MultiMLPs
+normal 5% [-14:24] 1% [-20:21] 7% [2:13] 13% [6:20] 2% [-2:4] 0.95$\pm$0.01
+uniform 37% [15:66] 25% [13:41] 19% [13:27] 29% [9:40] 1% [-2:6] 0.97$\pm$0.02
+lognormal 1% [-12:10] 1% [-15:11] 4% [2:8] -1% [-10:4] 0% [-2:1] 0.96$\pm$0.01
+gumbel 15% [-6:38] 11% [-8:30] 10% [-3:23] 11% [-5:23] 0% [-3:2] 0.96$\pm$0.00
+Table 3: Test results (95% prediction intervals) for 1000 observations generated using quadratic $f(x)$ with 2 predictors, and an additive noise ($\epsilon$) belonging to one of the four distributions.
+Relative metrics are shown over 5 seeds as ‘mean [min:max]’, and absolute metrics as ‘mean $\pm$ std’.
+Average performance over the baseline is highlighted in bold.
+For the cosine experiments (Table 2), all three methods (MultiXGBs, MultiETs, and MultiMLPs) exhibit positive relative improvements over the baseline in terms of the CWR, NMPIW, CRPS, and pinball loss.
+For instance, under normal noise, the MultiXGBs variant improves the $\Delta$CWR by an average of 18% (ranging between 10% and 35%), while still preserving a PICP of approximately 0.95.
+Similar positive trends are seen across the uniform and lognormal noise conditions.
+Although under Gumbel noise the improvements are less pronounced, SEMF still delivers performance that is at least on par with the baseline.
+For the quadratic experiments (Table 3), which represents a more complex heteroscedastic case, the improvements are even more substantial.
+Under normal noise, MultiXGBs achieve an average improvement of 39% in $\Delta$CWR, and, when the noise is uniform, gains can be as high as 65%.
+There is some variability—for example, under lognormal noise, the improvement in $\Delta$CWR for MultiXGBs is close to zero—but overall, the results indicate that SEMF produces tighter and more reliable prediction intervals compared to the baseline.
+Similar to the cosine case, the Gumbel noise setting shows the smallest relative gains; nevertheless, the performance under Gumbel noise remains competitive with the baseline.
+5.2 Benchmark Data
+We trained and tested 165 models corresponding to the three model types—MultiXGBs, MultiETs, and MultiMLPs—across 11 datasets, using five seeds for each combination.
+Table 4 presents the means and standard deviations for our metrics aggregated over the five seeds.
 Appendix E includes the results from each individual run.
-For comparison, we also present the non-conformalized prediction intervals in Appendix F, though we solely discuss the results for conformalized intervals on both complete and incomplete data.
-5.1 Complete Data
-The results of models with complete data are in Table 2.
-Overall, MultiXGBs and MultiMLPs performed generally well in producing intervals compared to their baselines, as shown by positive $\Delta$CWR and $\Delta$NMPIW while achieving similar $\Delta$PICP to the baselines.
-Notably, all models attained suitable intervals on naval_propulsion_plant and energy_efficiency.
-Interestingly, MultiMLPs also attained good relative performance improvements for point prediction, which can indicate that the chosen $\sigma_{k}$ was too low for generating performant prediction intervals.
-MultiETs attained mixed results, with significant improvements in $\Delta$CWR for other datasets such as naval_propulsion_plant, but its overall performance remains less conclusive.
-Table 2: Test results for all models with complete data at 95% quantiles aggregated over five seeds.
-For each metric, the mean and standard deviation of the performance across the seeds are separated by $\pm$.
+Relative Metrics Absolute Metrics
+Dataset $\Delta$CWR $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball $\Delta$PICP CWR NMPIW PICP
+MultiXGBs
+space_ga 11% [4:21] 9% [2:19] 11% [10:12] 9% [6:13] 0% [-2:2] 4.51$\pm$1.64 0.24$\pm$0.08 0.95$\pm$0.01
+cpu_activity 5% [-3:11] 5% [-4:11] 16% [10:26] 15% [0:36] 0% [-1:1] 9.77$\pm$0.55 0.10$\pm$0.01 0.95$\pm$0.01
+naval_propulsion_plant 163% [114:200] 61% [53:68] 72% [66:74] 45% [31:50] 0% [-3:2] 8.27$\pm$0.68 0.12$\pm$0.01 0.95$\pm$0.01
+miami_housing 5% [-4:12] 5% [-4:10] 12% [8:17] -12% [-21:-5] 0% [-1:0] 8.75$\pm$0.36 0.11$\pm$0.00 0.95$\pm$0.00
+kin8nm 18% [16:19] 16% [14:17] 23% [20:25] 8% [3:11] -1% [-1:-1] 2.31$\pm$0.04 0.41$\pm$0.01 0.94$\pm$0.00
+concrete_compressive_strength 40% [24:73] 29% [20:45] 32% [27:35] 12% [4:25] -3% [-5:-1] 3.22$\pm$0.21 0.29$\pm$0.02 0.94$\pm$0.02
+cars 32% [10:67] 24% [10:42] 27% [20:42] 17% [6:32] -1% [-3:0] 5.33$\pm$0.55 0.18$\pm$0.02 0.95$\pm$0.01
+energy_efficiency 128% [72:230] 53% [39:68] 67% [59:81] 54% [43:73] 3% [-2:8] 13.95$\pm$2.46 0.07$\pm$0.01 0.96$\pm$0.00
+california_housing 1% [-1:4] 1% [-1:4] 18% [15:23] -8% [-10:-3] 0% [-1:0] 2.24$\pm$0.08 0.42$\pm$0.02 0.95$\pm$0.01
+airfoil_self_noise -8% [-32:6] -12% [-47:6] 2% [-36:20] -17% [-43:4] 0% [-1:0] 2.29$\pm$0.38 0.44$\pm$0.10 0.97$\pm$0.01
+QSAR_fish_toxicity 19% [2:53] 14% [1:35] 13% [8:24] 15% [8:30] 0% [-1:2] 1.71$\pm$0.10 0.57$\pm$0.04 0.98$\pm$0.01
+MultiETs
+space_ga 5% [-2:15] 6% [-2:16] 8% [6:11] -6% [-9:-3] -1% [-3:0] 4.38$\pm$1.82 0.26$\pm$0.09 0.95$\pm$0.02
+cpu_activity 2% [-1:5] 6% [2:10] 3% [-2:7] -18% [-22:-11] -4% [-5:-3] 8.25$\pm$0.23 0.11$\pm$0.00 0.95$\pm$0.01
+naval_propulsion_plant 144% [118:160] 60% [56:63] 68% [64:72] 44% [40:52] -2% [-6:0] 3.65$\pm$0.25 0.26$\pm$0.02 0.95$\pm$0.01
+miami_housing -9% [-15:-5] -10% [-19:-4] -3% [-7:0] -73% [-81:-67] 0% [0:1] 6.44$\pm$0.27 0.15$\pm$0.01 0.95$\pm$0.00
+kin8nm 7% [5:10] 7% [4:8] 14% [12:17] -1% [-3:0] 0% [0:1] 2.13$\pm$0.05 0.45$\pm$0.01 0.95$\pm$0.01
+concrete_compressive_strength 6% [-14:26] 4% [-20:23] 9% [4:12] -3% [-12:3] 0% [-3:3] 3.05$\pm$0.25 0.31$\pm$0.03 0.95$\pm$0.01
+cars 15% [-25:54] 7% [-40:36] 6% [-8:20] 8% [-5:28] -1% [-3:4] 4.89$\pm$0.81 0.20$\pm$0.04 0.95$\pm$0.02
+energy_efficiency 15% [1:26] 14% [1:23] 3% [-15:12] -11% [-80:25] -2% [-3:0] 15.93$\pm$2.57 0.06$\pm$0.01 0.95$\pm$0.03
+california_housing 1% [-11:9] 2% [-11:11] 21% [16:25] -22% [-26:-17] -2% [-3:-1] 1.70$\pm$0.09 0.56$\pm$0.04 0.95$\pm$0.01
+airfoil_self_noise -13% [-39:22] -20% [-66:18] -17% [-52:19] -26% [-61:18] -1% [-2:1] 2.25$\pm$0.46 0.45$\pm$0.10 0.97$\pm$0.01
+QSAR_fish_toxicity -3% [-20:17] -6% [-26:18] -3% [-16:6] -8% [-19:2] -1% [-4:0] 1.71$\pm$0.22 0.58$\pm$0.09 0.97$\pm$0.01
+MultiMLPs
+space_ga 5% [-6:25] 6% [-4:21] 2% [-9:13] 0% [-13:8] -2% [-3:0] 5.10$\pm$1.78 0.21$\pm$0.07 0.94$\pm$0.02
+cpu_activity -14% [-16:-9] -16% [-20:-11] -9% [-14:-1] -3% [-11:17] 0% [-1:1] 8.57$\pm$0.14 0.11$\pm$0.00 0.95$\pm$0.01
+naval_propulsion_plant 23% [-2:84] 15% [-1:46] 17% [-3:46] -7% [-25:33] 0% [-1:1] 12.78$\pm$1.86 0.08$\pm$0.01 0.95$\pm$0.01
+miami_housing -21% [-35:-13] -28% [-56:-15] -13% [-23:-9] -26% [-38:-11] 0% [-1:0] 9.06$\pm$1.06 0.11$\pm$0.02 0.95$\pm$0.01
+kin8nm 9% [2:18] 8% [2:16] 8% [5:14] 8% [0:15] 0% [-2:1] 4.74$\pm$0.17 0.20$\pm$0.01 0.94$\pm$0.01
+concrete_compressive_strength 31% [-5:73] 21% [-2:43] 16% [-2:35] 7% [-19:31] -2% [-4:2] 3.43$\pm$0.17 0.28$\pm$0.02 0.95$\pm$0.02
+cars -12% [-44:21] -21% [-89:22] -12% [-23:5] -19% [-29:-13] 0% [-5:7] 5.40$\pm$1.25 0.19$\pm$0.04 0.96$\pm$0.03
+energy_efficiency 87% [46:149] 45% [30:63] 42% [33:56] 23% [0:50] -1% [-6:5] 21.91$\pm$3.25 0.04$\pm$0.01 0.94$\pm$0.02
+california_housing -11% [-14:-7] -12% [-16:-6] -2% [-5:0] -9% [-13:-3] 0% [-1:0] 2.13$\pm$0.08 0.45$\pm$0.02 0.95$\pm$0.01
+airfoil_self_noise 42% [16:63] 28% [14:39] 30% [17:43] 23% [5:32] -1% [-2:1] 4.92$\pm$0.31 0.20$\pm$0.01 0.97$\pm$0.01
+QSAR_fish_toxicity 8% [-15:36] 6% [-19:28] 1% [-9:13] 3% [-11:15] -1% [-3:1] 1.71$\pm$0.22 0.58$\pm$0.08 0.97$\pm$0.01
+Table 4: Test results for all models with at 95% quantiles aggregated over five seeds.
+For relative metrics, values are shown as ‘mean[min:max]’, and absolute metrics as ‘mean $\pm$ std’.
 Performance over the baseline is highlighted in bold.
-Interval Predictions Point Predictions
-Relative Absolute Relative Absolute
-Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
-MultiXGBs
-space_ga 6%$\pm$3-1%$\pm$07%$\pm$20.94$\pm$0.01 0.26$\pm$0.01 -9%$\pm$1-10%$\pm$20.60$\pm$0.01
-cpu_activity 16%$\pm$61%$\pm$112%$\pm$50.94$\pm$0.01 0.09$\pm$0.00 21%$\pm$1-1%$\pm$20.98$\pm$0.00
-naval_propulsion_plant 172%$\pm$140%$\pm$163%$\pm$20.95$\pm$0.01 $0$14%$\pm$4-12%$\pm$20.99$\pm$0.00
-miami_housing -7%$\pm$30%$\pm$0-8%$\pm$30.95$\pm$0.00 0.13$\pm$0.00 4%$\pm$43%$\pm$30.91$\pm$0.01
-kin8nm 6%$\pm$21%$\pm$05%$\pm$20.94$\pm$0.01 0.45$\pm$0.01 -20%$\pm$1-22%$\pm$10.63$\pm$0.01
-concrete_compressive_strength 41%$\pm$12-3%$\pm$231%$\pm$70.94$\pm$0.02 0.31$\pm$0.01 -26%$\pm$6-40%$\pm$110.85$\pm$0.01
-cars 40%$\pm$18-3%$\pm$130%$\pm$80.91$\pm$0.01 0.13$\pm$0.01 -3%$\pm$3-1%$\pm$30.95$\pm$0.00
-energy_efficiency 222%$\pm$45-4%$\pm$370%$\pm$30.92$\pm$0.02 0.05$\pm$0.01 -15%$\pm$22-16%$\pm$171.00$\pm$0.00
-california_housing -1%$\pm$30%$\pm$0-2%$\pm$40.95$\pm$0.00 0.42$\pm$0.01 -1%$\pm$1-1%$\pm$10.81$\pm$0.00
-airfoil_self_noise 21%$\pm$21-1%$\pm$215%$\pm$150.97$\pm$0.02 0.36$\pm$0.06 -64%$\pm$32-73%$\pm$410.86$\pm$0.05
-QSAR_fish_toxicity 34%$\pm$8-5%$\pm$329%$\pm$50.87$\pm$0.02 0.33$\pm$0.01 3%$\pm$31%$\pm$40.55$\pm$0.02
-MultiETs
-space_ga -6%$\pm$31%$\pm$1-7%$\pm$30.96$\pm$0.00 0.29$\pm$0.01 -15%$\pm$2-18%$\pm$30.54$\pm$0.02
-cpu_activity 11%$\pm$2-4%$\pm$013%$\pm$20.94$\pm$0.00 $0$14%$\pm$1-20%$\pm$20.98$\pm$0.00
-naval_propulsion_plant 137%$\pm$221%$\pm$057%$\pm$40.96$\pm$0.00 0.27$\pm$0.02 -316%$\pm$45-406%$\pm$430.96$\pm$0.01
-miami_housing -9%$\pm$1-1%$\pm$0-9%$\pm$10.95$\pm$0.00 $0$10%$\pm$2-20%$\pm$20.90$\pm$0.00
-kin8nm -10%$\pm$11%$\pm$1-12%$\pm$20.94$\pm$0.01 0.53$\pm$0.01 -34%$\pm$2-38%$\pm$20.48$\pm$0.01
-concrete_compressive_strength -8%$\pm$5-3%$\pm$2-6%$\pm$60.89$\pm$0.02 0.31$\pm$0.02 -67%$\pm$6-94%$\pm$80.76$\pm$0.01
-cars -25%$\pm$50%$\pm$3-34%$\pm$140.92$\pm$0.02 0.15$\pm$0.01 3%$\pm$41%$\pm$20.95$\pm$0.00
-energy_efficiency 10%$\pm$6-4%$\pm$212%$\pm$60.94$\pm$0.02 0.06$\pm$0.00 3%$\pm$21%$\pm$11.00$\pm$0.00
-california_housing 1%$\pm$2-2%$\pm$03%$\pm$20.95$\pm$0.00 0.55$\pm$0.01 -15%$\pm$1-21%$\pm$10.71$\pm$0.01
-airfoil_self_noise -16%$\pm$10-3%$\pm$2-17%$\pm$120.96$\pm$0.02 0.43$\pm$0.03 -118%$\pm$43-141%$\pm$490.80$\pm$0.08
-QSAR_fish_toxicity -6%$\pm$9-5%$\pm$2-2%$\pm$120.88$\pm$0.02 0.36$\pm$0.02 -6%$\pm$1-11%$\pm$10.53$\pm$0.01
-MultiMLPs
-space_ga 6%$\pm$3-1%$\pm$16%$\pm$20.95$\pm$0.01 0.23$\pm$0.00 0%$\pm$1-1%$\pm$10.75$\pm$0.01
-cpu_activity -7%$\pm$60%$\pm$1-9%$\pm$70.95$\pm$0.00 0.10$\pm$0.00 7%$\pm$25%$\pm$20.98$\pm$0.00
-naval_propulsion_plant 4%$\pm$161%$\pm$00%$\pm$140.96$\pm$0.00 0.08$\pm$0.01 -45%$\pm$25-36%$\pm$201.00$\pm$0.00
-miami_housing -38%$\pm$30%$\pm$1-61%$\pm$90.95$\pm$0.00 0.15$\pm$0.01 -3%$\pm$24%$\pm$20.91$\pm$0.00
-kin8nm 8%$\pm$50%$\pm$18%$\pm$50.95$\pm$0.01 0.20$\pm$0.01 7%$\pm$27%$\pm$20.93$\pm$0.00
-concrete_compressive_strength 11%$\pm$6-2%$\pm$211%$\pm$60.94$\pm$0.02 0.29$\pm$0.03 12%$\pm$515%$\pm$40.91$\pm$0.01
-cars 3%$\pm$12-3%$\pm$34%$\pm$110.92$\pm$0.02 0.13$\pm$0.01 -3%$\pm$30%$\pm$30.95$\pm$0.00
-energy_efficiency 53%$\pm$210%$\pm$434%$\pm$100.96$\pm$0.02 0.05$\pm$0.00 32%$\pm$333%$\pm$31.00$\pm$0.00
-california_housing -12%$\pm$30%$\pm$1-14%$\pm$40.95$\pm$0.00 0.43$\pm$0.01 7%$\pm$09%$\pm$10.82$\pm$0.00
-airfoil_self_noise 68%$\pm$28-2%$\pm$140%$\pm$110.97$\pm$0.01 0.18$\pm$0.01 18%$\pm$416%$\pm$50.97$\pm$0.00
-QSAR_fish_toxicity 5%$\pm$101%$\pm$13%$\pm$100.89$\pm$0.03 0.35$\pm$0.04 4%$\pm$57%$\pm$40.55$\pm$0.04
-5.2 Missing Data
-The results of models with 50% missing data are in Table 3.
-Note that relative metrics for SEMF are compared with the best result from any baseline imputer on that metric, regardless of how the imputer performed on the other metrics.
-Overall, the results are worse and less consistent than with complete data.
-MultiXGBs maintained good performance on some datasets, such as naval_propulsion_plant and energy_efficiency, but declined on others, like cpu_activity.
-MultiETs continued to exhibit mixed results, with only naval_propulsion_plant offering marginally better performance over the baseline.
-Similarly, MultiMLPs worked well on only one dataset, namely concrete_compressive_strength, while the other datasets had increased model uncertainty.
-Table 3: Test results for all models with missing data at 95% quantiles aggregated over five seeds.
-For each metric, the mean and standard deviation of the performance across the seeds are separated by $\pm$.
-Performance over the baseline is highlighted in bold.
-Interval Predictions Point Predictions
-Relative Absolute Relative Absolute
-Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
-MultiXGBs
-space_ga 0%$\pm$3-2%$\pm$12%$\pm$30.94$\pm$0.01 0.32$\pm$0.01 -7%$\pm$4-9%$\pm$30.45$\pm$0.06
-cpu_activity -14%$\pm$90%$\pm$2-20%$\pm$140.94$\pm$0.03 0.18$\pm$0.02 -1%$\pm$21-17%$\pm$130.83$\pm$0.07
-naval_propulsion_plant 18%$\pm$23-3%$\pm$212%$\pm$180.93$\pm$0.04 0.59$\pm$0.17 -14%$\pm$20-29%$\pm$350.77$\pm$0.05
-miami_housing -18%$\pm$13-1%$\pm$1-25%$\pm$220.94$\pm$0.01 0.20$\pm$0.02 -14%$\pm$11-18%$\pm$100.72$\pm$0.10
-kin8nm -9%$\pm$31%$\pm$1-13%$\pm$50.96$\pm$0.01 0.63$\pm$0.03 -3%$\pm$3-4%$\pm$40.40$\pm$0.02
-concrete_compressive_strength 1%$\pm$9-2%$\pm$21%$\pm$100.95$\pm$0.01 0.57$\pm$0.02 -16%$\pm$12-20%$\pm$140.58$\pm$0.08
-cars 9%$\pm$27-7%$\pm$46%$\pm$230.90$\pm$0.04 0.34$\pm$0.02 -30%$\pm$22-35%$\pm$240.66$\pm$0.15
-energy_efficiency 26%$\pm$24-4%$\pm$418%$\pm$180.95$\pm$0.04 0.27$\pm$0.05 -186%$\pm$265-153%$\pm$1880.91$\pm$0.08
-california_housing -6%$\pm$6-1%$\pm$1-8%$\pm$80.95$\pm$0.02 0.59$\pm$0.03 -4%$\pm$4-5%$\pm$50.69$\pm$0.05
-airfoil_self_noise -24%$\pm$6-1%$\pm$2-32%$\pm$120.95$\pm$0.03 0.76$\pm$0.07 -37%$\pm$31-43%$\pm$360.41$\pm$0.12
-QSAR_fish_toxicity 0%$\pm$8-4%$\pm$50%$\pm$110.91$\pm$0.05 0.50$\pm$0.11 -4%$\pm$4-2%$\pm$30.36$\pm$0.04
-MultiETs
-space_ga -11%$\pm$20%$\pm$2-14%$\pm$50.95$\pm$0.01 0.34$\pm$0.02 -12%$\pm$4-13%$\pm$50.40$\pm$0.06
-cpu_activity -15%$\pm$8-3%$\pm$2-18%$\pm$130.94$\pm$0.02 0.20$\pm$0.02 -21%$\pm$18-44%$\pm$140.82$\pm$0.07
-naval_propulsion_plant 4%$\pm$16-4%$\pm$34%$\pm$180.95$\pm$0.03 0.73$\pm$0.13 -37%$\pm$21-95%$\pm$480.71$\pm$0.06
-miami_housing -33%$\pm$100%$\pm$3-55%$\pm$240.95$\pm$0.01 0.27$\pm$0.02 -25%$\pm$10-39%$\pm$140.68$\pm$0.12
-kin8nm -14%$\pm$11%$\pm$1-17%$\pm$20.95$\pm$0.01 0.66$\pm$0.02 -11%$\pm$2-13%$\pm$30.32$\pm$0.03
-concrete_compressive_strength -20%$\pm$3-2%$\pm$4-27%$\pm$90.95$\pm$0.02 0.61$\pm$0.02 -37%$\pm$18-51%$\pm$220.48$\pm$0.06
-cars -32%$\pm$12-4%$\pm$4-53%$\pm$320.93$\pm$0.02 0.38$\pm$0.05 -38%$\pm$15-32%$\pm$140.62$\pm$0.16
-energy_efficiency -30%$\pm$28-4%$\pm$2-82%$\pm$1200.96$\pm$0.03 0.32$\pm$0.22 -69%$\pm$60-75%$\pm$430.94$\pm$0.03
-california_housing -6%$\pm$4-2%$\pm$1-4%$\pm$40.96$\pm$0.01 0.70$\pm$0.03 -14%$\pm$3-17%$\pm$30.61$\pm$0.05
-airfoil_self_noise -31%$\pm$4-3%$\pm$1-43%$\pm$90.94$\pm$0.02 0.75$\pm$0.05 -62%$\pm$51-89%$\pm$660.35$\pm$0.10
-QSAR_fish_toxicity -7%$\pm$7-3%$\pm$3-7%$\pm$60.92$\pm$0.04 0.50$\pm$0.08 -9%$\pm$9-10%$\pm$90.39$\pm$0.09
-MultiMLPs
-space_ga -25%$\pm$11-3%$\pm$2-34%$\pm$250.94$\pm$0.01 0.38$\pm$0.06 -40%$\pm$23-34%$\pm$150.21$\pm$0.23
-cpu_activity -40%$\pm$90%$\pm$2-73%$\pm$240.95$\pm$0.01 0.25$\pm$0.02 -15%$\pm$19-23%$\pm$170.75$\pm$0.12
-naval_propulsion_plant -43%$\pm$11-1%$\pm$2-89%$\pm$440.95$\pm$0.02 1.33$\pm$0.19 -108%$\pm$66-183%$\pm$195-0.03$\pm$0.47
-miami_housing -36%$\pm$7-1%$\pm$1-57%$\pm$200.94$\pm$0.02 0.22$\pm$0.02 -25%$\pm$8-28%$\pm$110.67$\pm$0.10
-kin8nm -19%$\pm$30%$\pm$0-24%$\pm$50.96$\pm$0.01 0.70$\pm$0.02 -16%$\pm$8-15%$\pm$90.37$\pm$0.04
-concrete_compressive_strength 5%$\pm$10-3%$\pm$25%$\pm$100.94$\pm$0.02 0.54$\pm$0.04 -12%$\pm$5-12%$\pm$60.54$\pm$0.05
-cars -22%$\pm$17-1%$\pm$4-38%$\pm$300.95$\pm$0.03 0.35$\pm$0.06 -28%$\pm$12-26%$\pm$130.66$\pm$0.10
-energy_efficiency -7%$\pm$11-4%$\pm$3-11%$\pm$140.95$\pm$0.03 0.31$\pm$0.08 1%$\pm$9-1%$\pm$160.95$\pm$0.03
-california_housing -24%$\pm$50%$\pm$1-34%$\pm$100.96$\pm$0.01 0.68$\pm$0.02 -5%$\pm$2-3%$\pm$30.61$\pm$0.06
-airfoil_self_noise -8%$\pm$6-4%$\pm$3-7%$\pm$80.93$\pm$0.04 0.61$\pm$0.07 -15%$\pm$8-9%$\pm$80.55$\pm$0.18
-QSAR_fish_toxicity -11%$\pm$5-2%$\pm$3-14%$\pm$80.92$\pm$0.04 0.53$\pm$0.11 -13%$\pm$7-14%$\pm$80.34$\pm$0.06
+The results on OpenML-CTR23 datasets reveal distinct performance patterns across our model variants.
+MultiXGBs shows improvement over the baseline in most datasets, with notable results on naval_propulsion_plant ($\Delta$CWR: 163%, $\Delta$NMPIW: 61%) and energy_efficiency ($\Delta$CWR: 128%, $\Delta$NMPIW: 53%).
+For concrete_compressive_strength and cars, MultiXGBs yields moderate improvements in interval quality ($\Delta$CWR: 40% and 32%, respectively).
+On airfoil_self_noise, however, MultiXGBs performs slightly below the baseline.
+While maintaining PICP values comparable to the baseline (within ±3%), MultiXGBs generally produces narrower prediction intervals, resulting in higher CWR values.
+MultiETs shows more variable performance.
+It achieves relative improvements on naval_propulsion_plant ($\Delta$CWR: 144%, $\Delta$NMPIW: 60%) and modest gains on energy_efficiency and cars, but underperforms on datasets like miami_housing, airfoil_self_noise, as well as QSAR_fish_toxicity.
+Interestingly, MultiETs often shows positive $\Delta$CRPS values even when the $\Delta$Pinball is negative, suggesting a trade-off between interval sharpness and quantile calibration.
+MultiMLPs exhibits mixed results across datasets.
+It performs well on energy_efficiency ($\Delta$CWR: 87%, $\Delta$NMPIW: 45%) and airfoil_self_noise ($\Delta$CWR: 42%, $\Delta$NMPIW: 28%), but shows lower performance on cpu_activity, miami_housing, and california_housing.
+The absolute CWR for MultiMLPs on energy_efficiency (21.91) indicates efficient intervals that maintain the target coverage with minimal width.
+Across all models, PICP values remain within the target range of approximately 0.95±0.03, indicating that SEMF maintains proper coverage while often reducing interval widths.
+These results suggest that SEMF’s benefits are most pronounced when applied to XGBoost, followed by the randomized trees and neural networks, the order depending on the metric.
 5.3 Discussion
-Our results indicate that SEMF, when combined with XGBoost and MLPs, performs strongly on datasets with complete data.
-Both models produce better prediction intervals than traditional quantile regression methods in the complete case.
-MultiMLPs also deliver good point predictions for some datasets despite the experimental design prioritizing interval estimation over point accuracy.
-This could indicate either the effectiveness of the chosen $\sigma_{k}$, which (indirectly) led to narrower prediction intervals or underfitting of the MLPs.
-The truth may lie somewhere in between; SEMF’s sampling operation, akin to cross-validation, partially helps combat overfitting, ensuring robust results despite variations in the data.
-In our preliminary experiments, we observed that increasing the depth of the baseline MLPs did not help and eventually led to overfitting.
-ETs exhibit more mixed performance, highly dependent on the dataset’s characteristics.
-One possible explanation behind the larger variations in predictive power compared to other models may be ETs’ reliance on randomized splits, which introduces significant variability in the prediction process.
-Consequently, ETs may not benefit from the iterative sampling and refinement of predictions inherent in SEMF.
-The robustness of SEMF diminishes when applied to datasets with missing data.
-Despite not offering improvements over the baseline in the presence of missing data, we have presented these results to transparently illustrate our framework’s current capabilities of handling missing inputs.
-From a theoretical standpoint, we believe there is value in further investigating how a single loss function that leverages EM’s missing data capabilities can be effectively applied.
-Given that SEMF performs well with complete data, a practical alternative might involve using the best imputation method for the data at hand before applying SEMF, effectively treating the data as complete.
-We expect our ablation study to perform well, given the complete data results.
-An important observation is the stability of SEMF across different random seeds, as indicated by consistent PICP and NMPIW metrics (full results in Appendix E), contrasting the significant variability observed in the baseline models.
-This suggests that SEMF offers a more reliable performance framework.
-Additionally, it is worth noting that our experiments did not precisely tune for conformalized prediction intervals but used the non-conformalized (raw) quantiles from SEMF.
-The non-conformalized intervals offer less coverage but produce better CWR and NMPIW than the baseline (Appendix F).
-Certain experimental choices—such as processing the columns separately, using the same hyperparameters for both complete and incomplete data, and fixing the number of latent nodes per input ($m_{k}$)—may have further constrained the models’ ability to tailor their performance to the varying complexities of different datasets.
+Our results indicate that overall, SEMF enhances uncertainty quantification in regression tasks, with varying effectiveness depending on the base model type and dataset characteristics.
+The framework often produces narrower prediction intervals while maintaining coverage targets, addressing the trade-off between interval width and reliability.
+The mechanism behind SEMF’s effectiveness appears to be model-dependent.
+For tree-based methods like XGBoost, which have limited capacity to model uncertainty directly, SEMF’s latent representation learning introduces a form of distributional modeling that enables more efficient uncertainty quantification.
+The iterative sampling and weighting procedure in the EM algorithm allows XGBoost to better capture the uncertainty structure in the data.
+This explains why MultiXGBs show the most consistent improvements across datasets.
+Neural networks, which inherently have greater representational capacity, benefit differently from SEMF.
+The framework’s structured approach to uncertainty quantification through latent variables complements the neural network’s flexibility, particularly in datasets with complex feature interactions like energy_efficiency.
+However, this advantage is not universal across all datasets, suggesting that the interaction between SEMF’s latent structure and the neural network’s own representational capabilities may sometimes create redundancies.
+Even in datasets where SEMF shows more modest improvements (such as space_ga and kin8nm), it still performs at least on par with SQR, our strong neural network baseline for simultaneous quantile estimation.
+This performance floor across diverse datasets highlights SEMF’s robustness in different modeling scenarios.
+The more variability in the performance of MultiETs compared to XGBoost can be attributed to two factors.
+First, the randomized splitting mechanism in ETs introduces inherent variability that may interfere with SEMF’s iterative refinement process.
+Second, ETs sometimes had less stable estimates for the weights (Eq.
+5), affecting the quality of the latent representations learned during training.
+This hypothesis is supported by the higher standard deviations observed in MultiETs’ performance metrics compared to the other models.
+Dataset characteristics also influence SEMF’s effectiveness.
+The framework shows particular strength on datasets like naval_propulsion_plant and energy_efficiency, which feature complex interactions among input variables that benefit from latent representation learning.
+Particularly for naval_propulsion_plant, the overall $y$ output is uniformly distributed, which the trees baselines fail to capture, but SEMF is more robust to this.
+Conversely, on simpler datasets or those with highly linear relationships, the additional complexity introduced by SEMF may not provide significant advantages over traditional quantile regression approaches.
+The relationship between CRPS and pinball Loss also provides insight into SEMF’s behavior.
+For several dataset-model combinations (e.g., MultiXGBs on miami_housing and california_housing), we observe positive $\Delta$CRPS alongside negative $\Delta$Pinball values.
+This divergence stems from fundamental differences in what these metrics evaluate.
+CRPS assesses the overall quality of the prediction interval, favoring narrow intervals that are well-centered around the true value.
+In contrast, pinball Loss measures the calibration of individual quantile predictions, penalizing miscalibrations at the tails more heavily.
+SEMF typically produces sharper intervals that may sacrifice some calibration at the exact quantile levels, resulting in better CRPS but potentially worse pinball loss compared to baseline models with wider but more conservatively calibrated intervals.
+The stability of SEMF across different random seeds, evidenced by the relatively small standard deviations in metrics like PICP and NMPIW, contrasts with the higher variability in baseline models.
+This suggests that SEMF’s sampling-based approach leads to more consistent uncertainty estimates, an important consideration for applications where predictive stability is required.
+The tight clustering of PICP values around the target level (0.95) across diverse datasets demonstrates SEMF’s ability to maintain calibration while improving efficiency.
+Further, our experimental design focused on the conformalized performance of SEMF, with CP applied consistently across all models.
+The performance improvements observed with this standardized approach suggest that SEMF could benefit further from specialized calibration techniques designed to leverage its sampling-based uncertainty estimation.
+Future work might explore adaptive conformalization methods that account for the specific characteristics of SEMF’s predictive distributions.
 6 Conclusion
-This paper introduces the Supervised Expectation-Maximization Framework (SEMF), a novel model-agnostic approach for generating prediction intervals in datasets with missing values.
+This paper introduces the Supervised Expectation-Maximization Framework (SEMF), a novel model-agnostic approach for generating prediction intervals in datasets with any machine learning model.
 SEMF draws from the EM algorithm for supervised learning to devise latent representations that produce better prediction intervals than quantile regression.
-Due to SEMF’s iterative simulation technique, training and inference can be done with complete and incomplete data.
-A comprehensive set of 330 experimental runs on 11 datasets with three different model types showed that SEMF, in the case of complete data, outperforms quantile regression, particularly on complete datasets and when using XGBoost, which intrinsically lacks latent representations.
-The results of the missing data are less positive and require further investigation.
-This research underscores SEMF’s potential in various application domains and opens new avenues for further exploration of supervised latent representation learning and uncertainty estimation.
+Our comprehensive evaluation demonstrates SEMF’s effectiveness in two complementary settings.
+First, controlled simulations across different noise distributions (normal, uniform, log-normal, and Gumbel) and data generating functions (cosine and quadratic with periodic components) show SEMF’s robustness to varying data characteristics, with particularly strong improvements for heteroscedastic data patterns.
+Second, a set of 165 experimental runs on 11 real-world benchmark datasets with three different model types confirmed that SEMF outperforms quantile regression in practical applications, particularly when using XGBoost, which intrinsically lacks latent representations.
+The framework’s ability to maintain consistent performance across different noise distributions while producing narrower intervals underscores its potential in various application domains and opens new avenues for further exploration of supervised latent representation learning and uncertainty estimation.
 7 Limitations & Future Work
 The primary limitation of this study was its reliance on the normality assumption, which may not fully capture the potential of SEMF across diverse data distributions.
-Although in Appendix G we demonstrate that the framework can learn non-normal patterns, further investigation and exploration of SEMF under other distributions, such as uniform, log-normal, and generalized extreme value distributions, are needed.
+Furthermore, in our simulations, we show that our framework can learn non-normal patterns; however, further investigation examination of how SEMF under broader distributional parameters can be interesting.
 The computational complexity of the approach presents another significant challenge, as the current implementation can be optimized for large-scale applications.
-Additionally, while the CWR metric is useful, it implicitly assumes that a 1% drop in PICP equates to a 1% reduction in NMPIW, thus assuming a uniform distribution.
+Additionally, while the CWR metric is valid, it implicitly assumes that a 1% drop in PICP equates to a 1% reduction in NMPIW, thus assuming a uniform distribution.
 Evaluating CWR under various distributional assumptions would provide a more comprehensive assessment of its implications.
-Additionally, SEMF has only been evaluated on MCAR data and does not address missing at random cases (MAR), which require further investigation for real-world applicability.
-Finally, applying the same hyper-parameters across all datasets without specific tuning for incomplete data likely contributes to the observed decline in accuracy and robustness.
 Future work presents several intriguing avenues for exploration.
@@
 These advancements can enhance the broader appeal of end-to-end approaches like SEMF in the ML community.
-Appendix A SEMF algorithm
-Algorithm 1 SEMF Training: two input sources where $x_{1}$ can be missing
-1:$y,x_{1},x_{2},R$
-2:$\theta,\phi_{1},\phi_{2},\xi$
-3:Initialize $\theta,\phi_{1},\phi_{2},\xi$
-4:Initialize $D_{y},D_{z_{1}},D_{z_{2}},D_{x},D_{i}$ to $\emptyset$
-5:Split $I=\{1,\ldots,N\}$ into $L$ batches $\{b_{1},\ldots,b_{L}\}$
-6:for $\ell=1$, …, $L$ do
-7:   for all $i$ in $b_{\ell}$ do
-8:      if $x_{1,i}$ is absent then
-9:         for $r=1$, …, $R$ do
-10:            Simulate $[j_{i,r},x_{1,i,r}]\sim p_{\xi}(\cdot|x_{2,i})$
-11:            Simulate $z_{1,i,r}\sim p_{\phi_{1}}(\cdot|x_{1,i,r})$
-12:            Simulate $z_{2,i,r}\sim p_{\phi_{2}}(\cdot|x_{2,i})$
-13:            Set $z_{i,r}=[z_{1,i,r},z_{2,i,r}]$
-14:         end for
-15:      else
-16:         for $r=1$, …, $R$ do
-17:            Simulate $z_{1,i,r}\sim p_{\phi_{1}}(\cdot|x_{1,i})$
-18:            Simulate $z_{2,i,r}\sim p_{\phi_{2}}(\cdot|x_{2,i})$
-19:            Set $z_{i,r}=[z_{1,i,r},z_{2,i,r}]$
-20:         end for
-21:      end if
-22:      for $r=1$, …, $R$ do
-23:         Compute
-$w_{i,r}=\frac{p_{\theta}(y_{i}|z_{i,r})}{\sum_{t=1}^{R}p_{\theta}(y_{i}|z_{i,t})}$
-24:         Update $D_{y}\leftarrow D_{y}\cup[y_{i}|z_{i,r}|w_{i,r}]$
-25:         Update $D_{z_{2}}\leftarrow D_{z_{2}}\cup[z_{2,i,r}|x_{2,i}|w_{i,r}]$
-26:         if $x_{1,i}$ is absent then
-27:            Update $D_{z_{1}}\leftarrow D_{z_{1}}\cup[z_{1,i,r}|x_{1,i,r}|w_{i,r}]$
-28:            Update $D_{x}\leftarrow D_{x}\cup[j_{i,r}|x_{2,i}|w_{i,r}]$
-29:         else
-30:            Update $D_{z_{1}}\leftarrow D_{z_{1}}\cup[z_{1,i,r}|x_{1,i}|w_{i,r}]$
-31:         end if
-32:      end for
-33:      Update $\theta\leftarrow Q_{y}(\theta,D_{y})$
-34:      Update $\phi_{1}\leftarrow Q_{1}(\phi_{1},D_{z_{1}})$
-35:      Update $\phi_{2}\leftarrow Q_{2}(\phi_{2},D_{z_{2}})$
-36:      Update $\xi\leftarrow Q_{x}(\xi,D_{x})$
-37:   end for
-38:end for
-39:Check convergence; Go to step 4 if not
-Algorithm 2 SEMF Inference
-1:$\theta^{*},\phi_{1}^{*},\phi_{2}^{*},\xi^{*},x_{1},x_{2},R$
-2:$z_{i,r}$
-3:for $r=1,\ldots,R$ do
-4:   if $x_{1,i}$ is absent then
-5:      Simulate $[j_{i,r},x_{1,i,r}]\sim p_{\xi^{*}}(\cdot|x_{2,i})$
-6:      Simulate $z_{1,i,r}\sim p_{\phi_{1}^{*}}(\cdot|x_{1,i,r})$
-7:      Simulate $z_{2,i,r}\sim p_{\phi_{2}^{*}}(\cdot|x_{2,i})$
-8:      Set $z_{i,r}=[z_{1,i,r},z_{2,i,r}]$
-9:   else
-10:      Simulate $z_{1,i,r}\sim p_{\phi_{1}^{*}}(\cdot|x_{1,i})$
-11:      Simulate $z_{2,i,r}\sim p_{\phi_{2}^{*}}(\cdot|x_{2,i})$
-12:      Set $z_{i,r}=[z_{1,i,r},z_{2,i,r}]$
-13:   end if
-14:end for
+Appendix A Optimal set of hyper-parameters
+Remark.
+The datasets mentioned here have been explained in Table 6 and Appendix B.
+The hyper-parameter tuning for SEMF is implemented and monitored using Weights & Biases [45].
+A random search is done in the hyper-parameter space for a maximum of 500 iterations on all 11 datasets, focusing on tuning the models only on the the non-conformalized performance.
+Key hyper-parameters are varied across a predefined set to balance accuracy and computational efficiency.
+The following grid is used for hyper-parameter tuning: the number of importance sampling operations $R\in\{5,10,25,50,100\}$ (100 is omitted for MultiMLPS), nodes per latent dimension $m_{k}\in\{1,5,10,20,30\}$, and standard deviations $\sigma_{k}\in\{0.001,0.01,0.1,1.0\}$.
+For some datasets, the $\sigma_{k}$ parameter is set to train_residual_models.
+This indicates that a separate residual model is trained to predict the latent-variable scale, which can be particularly suitable in heteroscedastic scenarios where noise levels vary across instances.
+Early stopping steps (patience) are set to five or ten, and $R_{\text{infer}}$, the $R$ for inference as $R_{\text{infer}}$, is explored at [30, 50, 70].
+For california_housing and cpu_activity, the $R_{\text{infer}}$ value is set to 30, while for all the other datasets, it is set to 50.
+We do this to ensure efficient computation, speed, and memory usage (especially for the GPU).
+Finally, the option to run the models in parallel must be consistently enabled.
+Table 5 shows the optimal set of the selected hyper-parameters.
+Dataset R $m_{k}$ $\sigma_{k}$ Patience $R_{\mathrm{infer}}$
+MultiXGBs
+space_ga 10 30 train_residual_models 5 50
+cpu_activity 5 30 1 5 30
+naval_propulsion_plant 5 30 0.01 5 50
+miami_housing 5 10 train_residual_models 5 50
+kin8nm 5 30 train_residual_models 10 50
+concrete_compressive_strength 25 30 train_residual_models 5 50
+cars 50 10 train_residual_models 10 50
+energy_efficiency 5 1 0.01 10 50
+california_housing 5 10 0.1 5 30
+airfoil_self_noise 25 1 train_residual_models 10 50
+QSAR_fish_toxicity 50 30 1 5 50
+MultiETs
+space_ga 10 30 train_residual_models 10 50
+cpu_activity 5 30 1 5 30
+naval_propulsion_plant 5 30 0.01 5 50
+miami_housing 10 10 train_residual_models 5 50
+kin8nm 5 30 train_residual_models 10 50
+concrete_compressive_strength 25 30 train_residual_models 10 50
+cars 100 5 train_residual_models 10 50
+energy_efficiency 5 1 0.01 10 50
+california_housing 5 10 0.1 5 30
+airfoil_self_noise 25 1 train_residual_models 10 50
+QSAR_fish_toxicity 50 30 train_residual_models 10 50
+MultiMLPs
+space_ga 25 10 train_residual_models 10 50
+cpu_activity 5 20 0.001 5 30
+naval_propulsion_plant 5 20 0.001 5 50
+miami_housing 5 20 train_residual_models 5 50
+kin8nm 5 20 train_residual_models 5 50
+concrete_compressive_strength 5 30 train_residual_models 10 50
+cars 5 30 train_residual_models 5 50
+energy_efficiency 50 30 0.1 10 50
+california_housing 5 20 0.01 5 30
+airfoil_self_noise 25 10 train_residual_models 10 50
+QSAR_fish_toxicity 50 30 train_residual_models 5 50
+Table 5: Hyper-parameters for MultiXGBs, MultiETs, and MultiMLPs.
+MultiXGBs and MultiMLPs benefit from early stopping to reduce computation time.
+Similarly, the baseline models for these instances use the same hyper-parameters for early stopping.
+Further, the number of epochs in the case of MultiMLPs is set as 1000, except for energy_efficiency and QSAR_fish_toxicity, where this is changed to 5000.
+Any model-specific hyperparameter we did not specify in this paper remains at the implementation’s default value (e.g., the number of leaves in XGBoost from [41]).
+Along with the supplementary code, we provide three additional CSV files: one for the results and hyperparameters of all 165 runs and the other two for the optimal hyperparameters of SEMF models, both raw (directly from SEMF) and conformalized.
+For training MultiXGBs and MultiETs, the computations are performed in parallel using CPU cores (Intel® Core i9-13900KF).
+Due to this, MultiETs are not fully deterministic and may exhibit slight variations between runs.
+For MultiMLPs, they are done on a GPU (NVIDIA® GeForce RTX 4090) which should give deterministic results across runs.
+All the computations are done on a machine with 32 GB of memory.
+The code provides further details on hardware and reproducibility.
 Appendix B Datasets for tabular benchmark
@@
 This leads us to the final list of 11 datasets listed in Table 6.
-Table 4: Summary of benchmark tabular datasets retained from [45]
 Dataset Name N Samples N Features OpenML Data ID Y [Min:Max] Source
@@
 QSAR_fish_toxicity 908 7 44970 [0.053:9.612] [57]
-Appendix C Optimal set of hyper-parameters
-The hyper-parameter tuning for SEMF is implemented and monitored using Weights & Biases [66].
-A random search is done in the hyper-parameter space for a maximum of 500 iterations on all 11 datasets, focusing on tuning the models only on the complete datasets.
-Key hyper-parameters are varied across a predefined set to balance accuracy and computational efficiency.
-The following grid is used for hyper-parameter tuning: the number of importance sampling operations $R\in\{5,10,25,50,100\}$ (100 is omitted for MultiMLPS), nodes per latent dimension $m_{k}\in\{1,5,10,20,30\}$, and standard deviations $\sigma_{m_{k}}\in\{0.001,0.01,0.1,1.0\}$.
-Early stopping steps (patience) are set to five or ten, and $R_{\text{infer}}$ is explored at [30, 50, 70].
-The option to run the models in parallel must be consistently enabled.
-Table 5 shows the optimal set of hyper-parameters.
-This table includes common hyper-parameters for complete and 50% datasets and another part showing $\xi_{\mathrm{nodes}}$, which is tuned manually and only relevant to the missing data.
-Table 5: Hyper-parameters for MultiXGBs, MultiETs, and MultiMLPs used for both complete and missing data.
-Dataset Complete and Missing Missing
-R $m_{k}$ $\sigma_{k}$ Patience $R_{\mathrm{infer}}$ $\xi_{\mathrm{nodes}}$
-MultiXGBs
-space_ga 10 30 1.0 5 70 100
-cpu_activity 5 30 1.0 5 70 50
-naval_propulsion_plant 5 30 0.01 5 50 100
-miami_housing 5 10 0.1 5 50 50
-kin8nm 5 30 1.0 10 70 100
-concrete_compressive_strength 25 30 1.0 5 70 100
-cars 50 10 1.0 10 70 100
-energy_efficiency 5 1 0.01 10 70 100
-california_housing 5 10 0.1 10 50 100
-airfoil_self_noise 25 1 0.01 10 70 100
-QSAR_fish_toxicity 50 30 1.0 5 70 100
-MultiETs
-space_ga 10 30 1.0 10 70 100
-cpu_activity 5 30 1.0 10 70 50
-naval_propulsion_plant 5 30 0.01 10 50 100
-miami_housing 10 10 0.1 10 50 50
-kin8nm 5 30 1.0 10 70 100
-concrete_compressive_strength 25 30 1.0 10 70 100
-cars 100 5 0.1 10 100 100
-energy_efficiency 5 1 0.01 10 70 100
-california_housing 5 10 0.1 5 50 100
-airfoil_self_noise 25 1 0.01 10 70 100
-QSAR_fish_toxicity 50 30 1.0 10 70 100
-MultiMLPs
-space_ga 25 10 0.001 10 50 100
-cpu_activity 5 20 0.001 5 50 50
-naval_propulsion_plant 5 20 0.001 5 50 100
-miami_housing 5 20 0.01 5 50 50
-kin8nm 5 20 0.001 5 50 100
-concrete_compressive_strength 5 30 0.001 10 50 100
-cars 5 30 0.1 5 50 100
-energy_efficiency 50 30 0.1 10 50 100
-california_housing 5 20 0.01 5 50 100
-airfoil_self_noise 25 10 0.01 10 50 100
-QSAR_fish_toxicity 50 30 1.0 10 70 100
-MultiXGBs and MultiMLPs benefit from early stopping to reduce computation time in complete and incomplete cases.
-Similarly, the baseline models for these instances use the same hyper-parameters for early stopping.
-Further, the number of epochs in the case of MultiMLPs is set as 1000, except for energy_efficiency and QSAR_fish_toxicity, where this is changed to 5000.
-Any model-specific hyperparameter we did not specify in this paper remains at the implementation’s default value (e.g., the number of leaves in XGBoost from [50]).
-Along with the supplementary code, we provide three additional CSV files: one for the results and hyperparameters of all 330 runs and the other two for the optimal hyperparameters of SEMF models, both raw (directly from SEMF) and conformalized.
-Additional conditions are applied only for experiments with missing data.
-Datasets california_housing, cpu_activity, miami_housing, and naval_propulsion_plant—have a patience of five to expedite the training process.
-Additionally, for california_housing and cpu_activity, the $R_{\text{infer}}$ value is set to 30, while for all the other datasets, it is set to 50.
-We do this to ensure efficient computation, speed, and memory usage (especially for the GPU).
-For training MultiXGBs and MultiETs, the computations are performed in parallel using CPU cores (Intel® Core i9-13900KF).
-For MultiMLPs, they are done on a GPU (NVIDIA® GeForce RTX 4090).
-The GPU is also consistently used for the missing data simulator and training $p_{\xi}$.
-All the computations are done on a machine with 32 GB of memory.
-The code provides further details on hardware and reproducibility.
-Appendix D Metrics for prediction intervals
-D.1 Common metrics
-The most common metrics for evaluating prediction intervals [52, 53] are:
+Table 6: Summary of benchmark tabular datasets retained from [44]
+Appendix C Metrics for prediction intervals
+C.1 Metrics’ definitions
+The most common metrics for evaluating prediction intervals [58, 59, 60, 61] are:
   • Prediction Interval Coverage Probability (PICP): This metric assesses the proportion of times the true value of the target variable falls within the constructed prediction intervals.
@@
     where $\max(y)$ and $\min(y)$ are the maximum and minimum values of the target variable, respectively. The interpretation remains the same as MPIW.
-D.2 Impact of relative metrics for modeling
+  • Continuous Ranked Probability Score (CRPS): We can approximate a full predictive cumulative distribution function by assuming that the forecast is uniformly distributed between its lower and upper bounds.
+Let $L$ and $U$ denote the lower and upper bounds (corresponding to the $\alpha/2$ and $21-\alpha/$ 2 quantiles, respectively).
+Then, for an observation $y$, we define
+    $\text{CRPS}(y,L,U)=\begin{cases}L-y+\frac{U-L}{3},&y\leq L,\\ \frac{(y-L)^{3}+(U-y)^{3}}{3\,(U-L)^{2}},&y\in[L,U],\\ y-U+\frac{U-L}{3},&y\geq U.\end{cases}$ (23)
+    This closed-form expression yields a lower CRPS when the forecast is both sharper and better calibrated.
+  • Quantile (pinball) Loss: For a given quantile level $\tau\in(0,1)$ and prediction $q$, the pinball loss is defined as:
+    $\ell_{\tau}(y,q)=(y-q)\left(\tau-\mathds{1}\{y\leq q\}\right),$ (24)
+    where $\mathds{1}\{y\leq q\}$ equals 1 if $y\leq q$ and 0 otherwise. This loss penalizes underestimation and overestimation asymmetrically and is used to tune the quantile predictions.
+C.2 Impact of relative metrics for modeling
 As our primary focus is on interval prediction, configurations demonstrating the most significant improvements in $\Delta$CWR and $\Delta$PICP are prioritized when selecting the optimal hyper-parameters.
@@
 In instances where no configuration meets the initial improvement criteria for both metrics, we relax the requirement for positive $\Delta$PICP to accept values greater than -5% and subsequently -10%, allowing us to consider configurations where SEMF significantly improves CWR, even if the PICP improvement is less marked but remains within an acceptable range for drawing comparisons.
-Appendix E Full conformalized results
+Appendix D Full synthetic experiments results
+Relative Metrics Absolute Metrics
+Seed $\Delta$CWR $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball $\Delta$PICP CWR NMPIW PICP
+Normal
+0 17% 15% 6% 11% 0% 1.78 0.54 0.97
+10 11% 11% 3% 5% -1% 1.60 0.60 0.96
+20 19% 18% 8% 14% -2% 1.89 0.50 0.95
+30 10% 10% 5% 10% -1% 1.99 0.48 0.96
+40 35% 28% 10% 17% -3% 2.49 0.37 0.93
+Uniform
+0 -2% -5% -1% 0% 2% 2.36 0.42 0.98
+10 34% 26% 12% 26% -1% 2.12 0.47 0.99
+20 9% 8% 5% 12% 1% 2.11 0.47 0.99
+30 28% 26% 9% 19% -5% 3.59 0.25 0.91
+40 58% 39% 23% 32% -3% 3.61 0.26 0.95
+Lognormal
+0 23% 19% 15% 9% 0% 2.08 0.46 0.96
+10 59% 39% 18% 7% -3% 2.07 0.46 0.95
+20 8% 9% 8% 2% -1% 1.96 0.48 0.95
+30 9% 8% 9% 5% 0% 1.59 0.62 0.98
+40 38% 28% 15% 14% 0% 2.30 0.42 0.97
+Gumbel
+0 4% 1% 7% 16% 4% 1.61 0.60 0.98
+10 17% 18% 7% 4% -3% 1.78 0.53 0.94
+20 20% 17% 14% 16% -1% 2.01 0.48 0.96
+30 -5% -9% 1% 0% 4% 1.75 0.55 0.96
+40 15% 14% 10% 5% -1% 2.02 0.47 0.95
+Table 7: Test results for MultiXGBs, over 5 seeds, with cosine $f(x)$ using 2 predictors, organized by distribution and seed. Relative metrics are in bold when positive.
+Relative Metrics Absolute Metrics
+Seed $\Delta$CWR $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball $\Delta$PICP CWR NMPIW PICP
+Normal
+0 8% 8% 4% 0% -1% 1.98 0.48 0.96
+10 17% 18% 6% 12% -3% 1.66 0.57 0.94
+20 1% 0% 0% 2% 1% 1.81 0.54 0.97
+30 15% 14% 8% 16% -1% 1.99 0.49 0.97
+40 32% 27% 8% 9% -4% 2.46 0.38 0.94
+Uniform
+0 35% 28% 10% 21% -3% 2.80 0.35 0.97
+10 3% -1% 3% 8% 4% 2.45 0.41 0.99
+20 7% 5% 4% 11% 1% 2.58 0.37 0.96
+30 9% 8% 6% 14% 0% 3.42 0.28 0.97
+40 29% 25% 7% 12% -3% 3.63 0.26 0.96
+Lognormal
+0 -14% -18% 8% -10% 1% 1.85 0.52 0.96
+10 18% 17% 10% -2% -2% 1.87 0.51 0.95
+20 12% 12% 13% 0% -1% 2.03 0.47 0.96
+30 -6% -8% 5% -4% 2% 1.74 0.56 0.98
+40 14% 13% 11% 2% -2% 2.10 0.46 0.97
+Gumbel
+0 -4% -7% 9% 0% 2% 1.66 0.59 0.97
+10 8% 7% 8% -2% 0% 1.72 0.56 0.97
+20 -2% -2% 7% 5% 0% 1.90 0.50 0.96
+30 2% 4% 6% -2% -2% 1.85 0.51 0.95
+40 14% 14% 11% 5% -2% 1.96 0.50 0.97
+Table 8: Test results for MultiETs, over 5 seeds, with cosine $f(x)$ using 2 predictors, organized by distribution and seed. Relative metrics are in bold when positive.
+Relative Metrics Absolute Metrics
+Seed $\Delta$CWR $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball $\Delta$PICP CWR NMPIW PICP
+Normal
+0 0% 1% -2% 0% -1% 1.98 0.48 0.95
+10 9% 8% 4% 5% 0% 1.71 0.56 0.96
+20 -2% -1% -1% -3% -1% 2.04 0.46 0.94
+30 -1% -2% 1% 3% 0% 2.20 0.44 0.97
+40 7% 7% 1% 5% 0% 2.48 0.39 0.96
+Uniform
+0 2% 2% 0% 8% 0% 2.98 0.33 0.98
+10 6% 2% 5% 14% 3% 2.80 0.35 0.98
+20 3% 7% 0% -5% -4% 3.02 0.31 0.94
+30 -2% -2% -1% -4% 0% 3.85 0.25 0.96
+40 13% 10% 8% 23% 1% 3.67 0.27 0.99
+Lognormal
+0 2% 3% 1% 4% 0% 1.71 0.57 0.98
+10 18% 16% 7% 8% 0% 1.99 0.48 0.95
+20 -4% -5% -2% 0% 1% 1.86 0.52 0.98
+30 15% 13% 7% 9% -1% 1.74 0.56 0.98
+40 4% 5% 1% 2% -1% 2.10 0.46 0.97
+Gumbel
+0 -3% -1% -2% -4% -1% 1.54 0.64 0.98
+10 8% 8% 1% 4% -1% 1.88 0.51 0.95
+20 8% 8% 3% 4% 0% 2.14 0.44 0.94
+30 13% 14% 2% -4% -2% 1.87 0.51 0.95
+40 -3% -5% -2% -2% 2% 1.76 0.56 0.98
+Table 9: Test results for MultiMLPs, over 5 seeds, with cosine $f(x)$ using 2 predictors, organized by distribution and seed. Relative metrics are in bold when positive.
+Relative Metrics Absolute Metrics
+Seed $\Delta$CWR $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball $\Delta$PICP CWR NMPIW PICP
+Normal
+0 35% 26% 28% 16% 0% 5.80 0.17 0.96
+10 65% 41% 34% 35% -3% 4.80 0.20 0.96
+20 40% 29% 31% 16% 0% 4.41 0.21 0.94
+30 30% 23% 26% 23% 0% 5.32 0.18 0.95
+40 28% 22% 27% 28% 0% 5.83 0.16 0.95
+Uniform
+0 80% 44% 42% 30% 0% 10.12 0.09 0.94
+10 54% 34% 47% 32% 1% 5.78 0.17 0.98
+20 62% 38% 40% 31% 0% 5.62 0.17 0.96
+30 62% 39% 39% 16% -1% 7.28 0.13 0.95
+40 68% 41% 39% 18% -1% 7.93 0.12 0.97
+Lognormal
+0 2% 1% 20% 0% 1% 4.37 0.22 0.96
+10 -2% -2% 14% -4% 1% 2.92 0.34 0.99
+20 7% 2% 27% 6% 5% 3.81 0.25 0.94
+30 -13% -16% 15% 3% 1% 3.28 0.29 0.96
+40 7% 6% 23% 16% 0% 3.42 0.29 0.98
+Gumbel
+0 34% 27% 30% 20% -2% 5.03 0.19 0.96
+10 49% 35% 24% 21% -3% 3.70 0.26 0.95
+20 0% 0% 9% 3% 1% 2.81 0.35 0.99
+30 11% 7% 17% 14% 4% 4.49 0.21 0.95
+40 18% 15% 21% 20% 1% 4.37 0.22 0.97
+Table 10: Test results for MultiXGBs, over 5 seeds, with quadratic $f(x)$ using 2 predictors, organized by distribution and seed. Relative metrics are in bold when positive.
+Relative Metrics Absolute Metrics
+Seed $\Delta$CWR $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball $\Delta$PICP CWR NMPIW PICP
+Normal
+0 7% 7% 0% -4% -1% 5.77 0.16 0.92
+10 35% 26% 21% 23% 0% 4.66 0.21 0.98
+20 29% 24% 15% 14% -2% 4.69 0.20 0.93
+30 26% 21% 15% 13% -1% 4.91 0.19 0.95
+40 21% 16% 18% 12% 1% 4.79 0.20 0.98
+Uniform
+0 28% 23% 16% -4% -1% 8.81 0.11 0.94
+10 25% 20% 24% 21% 1% 4.95 0.20 0.99
+20 47% 34% 25% 7% -3% 4.51 0.21 0.96
+30 30% 22% 22% 15% 2% 5.85 0.17 0.98
+40 71% 42% 33% 21% 0% 7.34 0.13 0.97
+Lognormal
+0 -9% -13% 0% -6% 3% 4.34 0.22 0.96
+10 -1% -2% 13% 12% 1% 3.36 0.29 0.98
+20 33% 27% 16% 10% -2% 4.09 0.23 0.95
+30 8% 8% 11% 9% 0% 3.29 0.30 0.97
+40 10% 8% 15% 9% 1% 3.78 0.26 0.97
+Gumbel
+0 5% 4% 10% 7% 1% 4.67 0.21 0.98
+10 42% 30% 27% 28% -1% 4.55 0.21 0.94
+20 33% 25% 16% 15% -1% 3.44 0.28 0.97
+30 -3% -4% 5% 0% 1% 3.86 0.25 0.97
+40 68% 42% 30% 24% -2% 3.97 0.24 0.97
+Table 11: Test results for MultiETs, over 5 seeds, with quadratic $f(x)$ using 2 predictors, organized by distribution and seed. Relative metrics are in bold when positive.
+Relative Metrics Absolute Metrics
+Seed $\Delta$CWR $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball $\Delta$PICP CWR NMPIW PICP
+Normal
+0 -14% -20% 3% 6% 3% 6.63 0.15 0.97
+10 24% 21% 13% 19% -2% 6.07 0.16 0.96
+20 0% -4% 2% 6% 3% 5.81 0.16 0.93
+30 -1% -5% 7% 17% 4% 6.13 0.16 0.95
+40 15% 14% 9% 20% 0% 7.47 0.13 0.95
+Uniform
+0 42% 26% 25% 36% 6% 15.86 0.06 0.97
+10 27% 20% 14% 30% 2% 9.42 0.11 1.00
+20 66% 41% 27% 29% -2% 10.02 0.10 0.95
+30 34% 26% 13% 9% 0% 12.50 0.08 0.96
+40 15% 13% 16% 40% -1% 13.66 0.07 0.97
+Lognormal
+0 1% 2% 4% 0% -1% 5.41 0.18 0.97
+10 3% 2% 8% 0% 1% 5.23 0.18 0.96
+20 10% 11% 3% -10% -2% 4.79 0.20 0.95
+30 -12% -15% 4% 0% 1% 3.86 0.25 0.97
+40 4% 3% 2% 4% 0% 4.88 0.20 0.97
+Gumbel
+0 38% 30% 19% 18% -3% 6.90 0.14 0.96
+10 32% 25% 23% 23% -1% 4.56 0.21 0.96
+20 -6% -8% -3% -5% 2% 4.10 0.23 0.96
+30 13% 12% 3% 8% -1% 5.40 0.18 0.96
+40 -3% -5% 10% 10% 1% 5.14 0.19 0.96
+Table 12: Test results for MultiMLPs, over 5 seeds, with quadratic $f(x)$ using 2 predictors, organized by distribution and seed. Relative metrics are in bold when positive.
+Appendix E Full conformalized results on benchmark datasets
+Relative Metrics Absolute Metrics
+Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball CWR NMPIW PICP
+space_ga 9% 0% 9% 12% 10% 6.28 0.15 0.95
+space_ga 5% 2% 2% 11% 13% 2.95 0.32 0.95
+space_ga 4% 1% 2% 10% 9% 3.23 0.30 0.97
+space_ga 21% -2% 19% 11% 9% 3.37 0.28 0.96
+space_ga 16% -2% 15% 11% 6% 6.74 0.14 0.94
+cpu_activity 7% 0% 7% 17% 10% 9.84 0.10 0.96
+cpu_activity -3% 1% -4% 10% 0% 8.69 0.11 0.96
+cpu_activity 11% -1% 11% 14% 10% 10.20 0.09 0.94
+cpu_activity 11% 1% 10% 26% 36% 10.14 0.09 0.95
+cpu_activity 1% 1% 0% 13% 18% 9.99 0.10 0.95
+naval_propulsion_plant 200% -3% 68% 73% 43% 9.19 0.10 0.92
+naval_propulsion_plant 114% 0% 53% 66% 31% 7.21 0.13 0.95
+naval_propulsion_plant 159% 1% 61% 74% 50% 7.97 0.12 0.96
+naval_propulsion_plant 187% 0% 65% 74% 50% 8.73 0.11 0.95
+naval_propulsion_plant 156% 2% 60% 71% 50% 8.24 0.12 0.96
+miami_housing 9% -1% 9% 14% -5% 9.08 0.10 0.95
+miami_housing -4% 0% -4% 8% -12% 8.28 0.12 0.96
+miami_housing 2% 0% 3% 9% -19% 8.49 0.11 0.94
+miami_housing 12% 0% 10% 17% -5% 9.23 0.10 0.95
+miami_housing 8% -1% 9% 11% -21% 8.70 0.11 0.95
+kin8nm 16% -1% 15% 22% 3% 2.29 0.41 0.94
+kin8nm 19% -1% 16% 25% 11% 2.31 0.41 0.94
+kin8nm 16% -1% 14% 20% 6% 2.27 0.42 0.95
+kin8nm 19% -1% 17% 22% 8% 2.29 0.41 0.95
+kin8nm 19% -1% 17% 24% 11% 2.40 0.40 0.95
+concrete_compressive_strength 31% -1% 24% 35% 25% 2.95 0.33 0.96
+concrete_compressive_strength 73% -5% 45% 34% 14% 3.60 0.26 0.92
+concrete_compressive_strength 27% -3% 24% 29% 6% 3.14 0.29 0.92
+concrete_compressive_strength 47% -2% 34% 33% 9% 3.14 0.30 0.94
+concrete_compressive_strength 24% -1% 20% 27% 4% 3.25 0.29 0.93
+cars 32% -2% 25% 25% 6% 5.78 0.17 0.95
+cars 10% -1% 10% 20% 18% 5.02 0.19 0.95
+cars 27% 0% 22% 26% 17% 5.08 0.19 0.95
+cars 67% -3% 42% 42% 32% 6.15 0.15 0.94
+cars 22% -2% 20% 23% 15% 4.63 0.21 0.97
+energy_efficiency 114% 1% 52% 66% 43% 15.29 0.06 0.96
+energy_efficiency 112% 0% 53% 67% 50% 14.18 0.07 0.95
+energy_efficiency 109% -2% 53% 64% 44% 10.78 0.09 0.96
+energy_efficiency 230% 8% 68% 81% 73% 17.66 0.05 0.96
+energy_efficiency 72% 5% 39% 59% 58% 11.84 0.08 0.95
+california_housing 3% 0% 3% 23% -3% 2.09 0.46 0.96
+california_housing 4% -1% 4% 18% -10% 2.27 0.42 0.94
+california_housing 2% 0% 1% 21% -7% 2.29 0.41 0.95
+california_housing 0% 0% 0% 15% -9% 2.27 0.41 0.94
+california_housing -1% 0% -1% 15% -10% 2.28 0.42 0.95
+airfoil_self_noise -5% 0% -5% 9% -19% 2.60 0.37 0.96
+airfoil_self_noise 6% 0% 6% 20% 4% 2.56 0.38 0.98
+airfoil_self_noise -32% 0% -47% -36% -43% 1.57 0.63 0.99
+airfoil_self_noise -4% 0% -5% 8% -10% 2.26 0.43 0.97
+airfoil_self_noise -8% -1% -7% 9% -15% 2.43 0.39 0.95
+QSAR_fish_toxicity 2% 0% 1% 8% 10% 1.80 0.54 0.97
+QSAR_fish_toxicity 11% -1% 10% 8% 12% 1.63 0.59 0.96
+QSAR_fish_toxicity 4% 2% 3% 9% 8% 1.80 0.54 0.97
+QSAR_fish_toxicity 26% 0% 20% 17% 18% 1.55 0.64 0.99
+QSAR_fish_toxicity 53% -1% 35% 24% 30% 1.73 0.57 0.98
 Table 13: Test results for MultiXGBs with complete data at 95% quantiles for seeds 0, 10, 20, 30, 40, with rows ordered by seed (ascending). Performance over the baseline is highlighted in bold.
-Interval Predictions Point Predictions
-Relative Absolute Relative Absolute
-Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP MPIW NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
-space_ga 2% -2% 4% 0.95 2.40 0.26 -10% -13% 0.60
-space_ga 10% -1% 10% 0.95 2.34 0.25 -10% -10% 0.62
-space_ga 8% -2% 9% 0.93 2.24 0.24 -9% -8% 0.60
-space_ga 6% -2% 7% 0.94 2.38 0.26 -8% -10% 0.59
-space_ga 7% -1% 7% 0.95 2.55 0.28 -8% -10% 0.59
-cpu_activity 7% 1% 5% 0.95 0.49 0.09 22% 2% 0.98
-cpu_activity 20% 1% 17% 0.93 0.45 0.09 20% -5% 0.98
-cpu_activity 20% 0% 17% 0.93 0.45 0.09 21% 0% 0.98
-cpu_activity 21% 1% 16% 0.94 0.45 0.09 23% -2% 0.98
-cpu_activity 10% 2% 7% 0.95 0.47 0.09 20% -3% 0.98
-naval_propulsion_plant 163% -1% 62% 0.95 0.36 0.11 -15% -12% 0.99
-naval_propulsion_plant 190% 0% 66% 0.95 0.37 0.11 -20% -14% 0.99
-naval_propulsion_plant 151% 0% 60% 0.96 0.40 0.12 -12% -9% 0.99
-naval_propulsion_plant 185% -1% 65% 0.94 0.35 0.10 -9% -15% 0.99
-naval_propulsion_plant 169% -1% 63% 0.96 0.37 0.11 -15% -11% 0.99
-miami_housing -2% 0% -2% 0.95 0.99 0.12 4% 1% 0.91
-miami_housing -10% 0% -10% 0.95 1.03 0.13 7% 7% 0.90
-miami_housing -6% 1% -8% 0.95 0.99 0.12 9% 5% 0.91
-miami_housing -7% 0% -8% 0.95 0.99 0.12 3% -1% 0.91
-miami_housing -10% 0% -12% 0.95 1.05 0.13 -4% 0% 0.90
-kin8nm 6% 0% 6% 0.94 2.27 0.45 -21% -24% 0.62
-kin8nm 4% 1% 3% 0.95 2.39 0.47 -19% -21% 0.64
-kin8nm 9% 1% 8% 0.94 2.20 0.43 -17% -20% 0.64
-kin8nm 9% 1% 7% 0.95 2.36 0.46 -21% -22% 0.63
-kin8nm 3% 1% 2% 0.94 2.33 0.46 -20% -23% 0.62
-concrete_compressive_strength 40% -3% 31% 0.93 1.42 0.29 -16% -22% 0.86
-concrete_compressive_strength 47% -6% 36% 0.91 1.48 0.30 -29% -52% 0.84
-concrete_compressive_strength 59% -2% 39% 0.96 1.60 0.32 -21% -34% 0.86
-concrete_compressive_strength 35% -3% 28% 0.95 1.48 0.30 -28% -39% 0.83
-concrete_compressive_strength 24% 0% 19% 0.97 1.59 0.32 -34% -52% 0.86
-cars 56% -4% 38% 0.91 0.70 0.12 -7% -4% 0.95
-cars 36% -5% 30% 0.90 0.86 0.14 -4% -4% 0.95
-cars 30% -3% 25% 0.91 0.76 0.13 2% 5% 0.96
-cars 65% -1% 40% 0.93 0.75 0.13 -5% -1% 0.95
-cars 16% -4% 17% 0.91 0.87 0.15 -2% 0% 0.95
-energy_efficiency 165% -5% 64% 0.92 0.20 0.06 -50% -29% 1.00
-energy_efficiency 288% 0% 74% 0.91 0.14 0.04 -3% -13% 1.00
-energy_efficiency 217% -1% 69% 0.96 0.18 0.05 -21% -31% 1.00
-energy_efficiency 253% -3% 73% 0.92 0.16 0.04 17% 16% 1.00
-energy_efficiency 185% -10% 68% 0.91 0.23 0.07 -17% -22% 1.00
-california_housing -5% 0% -6% 0.95 1.82 0.43 -2% -3% 0.81
-california_housing 4% 0% 4% 0.95 1.72 0.41 -1% -2% 0.81
-california_housing -4% 1% -5% 0.95 1.76 0.42 0% -1% 0.82
-california_housing 1% 0% 1% 0.95 1.74 0.41 0% 0% 0.81
-california_housing -3% 1% -4% 0.95 1.79 0.42 -1% -1% 0.82
-airfoil_self_noise 15% -4% 16% 0.95 1.73 0.37 -82% -93% 0.86
-airfoil_self_noise 45% 1% 30% 0.97 1.42 0.30 -18% -18% 0.93
-airfoil_self_noise 5% 1% 4% 0.98 1.86 0.40 -73% -76% 0.85
-airfoil_self_noise -6% 0% -7% 0.98 2.15 0.46 -109% -138% 0.78
-airfoil_self_noise 46% -2% 33% 0.95 1.37 0.29 -41% -43% 0.89
-QSAR_fish_toxicity 21% -2% 19% 0.88 2.12 0.33 3% 3% 0.53
-QSAR_fish_toxicity 39% -9% 35% 0.85 1.97 0.31 8% 7% 0.58
-QSAR_fish_toxicity 43% -3% 32% 0.89 2.21 0.34 4% 3% 0.57
-QSAR_fish_toxicity 38% -3% 30% 0.89 2.11 0.33 -2% -1% 0.54
-QSAR_fish_toxicity 29% -6% 28% 0.85 2.11 0.33 0% -6% 0.52
+Relative Metrics Absolute Metrics
+Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball CWR NMPIW PICP
+space_ga 4% -2% 6% 6% -9% 5.86 0.16 0.95
+space_ga -2% 0% -2% 7% -3% 2.66 0.36 0.96
+space_ga 10% -1% 10% 7% -5% 3.19 0.30 0.97
+space_ga -2% 0% -2% 7% -5% 2.99 0.33 0.97
+space_ga 15% -3% 16% 11% -9% 7.22 0.13 0.91
+cpu_activity 5% -5% 10% 2% -22% 8.69 0.11 0.94
+cpu_activity 1% -3% 4% 7% -11% 8.05 0.12 0.96
+cpu_activity 3% -5% 8% 5% -11% 8.26 0.11 0.94
+cpu_activity -1% -4% 2% -2% -22% 8.14 0.12 0.95
+cpu_activity 3% -4% 7% 2% -22% 8.11 0.12 0.95
+naval_propulsion_plant 154% -6% 63% 69% 40% 3.55 0.27 0.94
+naval_propulsion_plant 153% 0% 60% 68% 46% 3.95 0.24 0.95
+naval_propulsion_plant 160% 0% 61% 72% 52% 3.86 0.25 0.96
+naval_propulsion_plant 136% 0% 58% 66% 41% 3.67 0.26 0.95
+naval_propulsion_plant 118% -4% 56% 64% 41% 3.23 0.29 0.95
+miami_housing -13% 1% -16% -4% -75% 6.25 0.15 0.96
+miami_housing -6% 0% -6% 0% -67% 6.71 0.14 0.95
+miami_housing -5% 1% -6% 0% -67% 6.47 0.15 0.95
+miami_housing -5% 0% -4% -1% -75% 6.75 0.14 0.95
+miami_housing -15% 1% -19% -7% -81% 6.03 0.16 0.95
+kin8nm 8% 0% 7% 15% 0% 2.14 0.44 0.95
+kin8nm 7% 0% 7% 13% -3% 2.11 0.45 0.95
+kin8nm 7% 0% 7% 14% 0% 2.10 0.46 0.96
+kin8nm 5% 1% 4% 12% 0% 2.08 0.46 0.95
+kin8nm 10% 1% 8% 17% 0% 2.22 0.43 0.95
+concrete_compressive_strength 26% -3% 23% 12% 0% 3.27 0.29 0.94
+concrete_compressive_strength 5% 0% 5% 8% 0% 2.97 0.33 0.97
+concrete_compressive_strength 13% 3% 9% 12% 3% 3.39 0.28 0.93
+concrete_compressive_strength -14% 2% -20% 4% -12% 2.73 0.35 0.95
+concrete_compressive_strength 3% 0% 3% 9% -4% 2.87 0.34 0.97
+cars 12% -1% 12% 3% 0% 4.95 0.19 0.96
+cars 2% -2% 4% -1% 0% 4.97 0.19 0.96
+cars 29% -3% 24% 15% 15% 3.64 0.27 0.97
+cars -25% 4% -40% -8% -5% 6.19 0.15 0.92
+cars 54% -2% 36% 20% 28% 4.70 0.21 0.97
+energy_efficiency 26% -3% 23% 12% 0% 15.20 0.06 0.96
+energy_efficiency 16% -3% 17% 6% 0% 16.14 0.06 0.94
+energy_efficiency 24% 0% 20% 9% 25% 14.43 0.07 0.99
+energy_efficiency 8% 0% 8% 4% 0% 20.70 0.04 0.91
+energy_efficiency 1% -1% 1% -15% -80% 13.19 0.07 0.96
+california_housing -$-1$1% 16% -26% 1.55 0.62 0.96
+california_housing -1% -2% 1% 20% -24% 1.68 0.57 0.95
+california_housing 2% -2% 4% 24% -17% 1.70 0.56 0.95
+california_housing 4% -3% 7% 20% -25% 1.76 0.53 0.93
+california_housing 9% -3% 11% 25% -17% 1.83 0.52 0.94
+airfoil_self_noise -21% -1% -24% -26% -40% 2.15 0.45 0.97
+airfoil_self_noise 22% 0% 18% 19% 18% 2.91 0.34 0.98
+airfoil_self_noise -20% -2% -22% -21% -36% 2.17 0.44 0.96
+airfoil_self_noise -39% 1% -66% -52% -61% 1.52 0.64 0.98
+airfoil_self_noise -6% -1% -5% -3% -12% 2.49 0.39 0.97
+QSAR_fish_toxicity -13% 0% -14% -4% -10% 1.69 0.57 0.97
+QSAR_fish_toxicity -20% 0% -26% -16% -19% 1.36 0.72 0.98
+QSAR_fish_toxicity 17% -4% 18% 6% -2% 2.00 0.47 0.94
+QSAR_fish_toxicity 17% -2% 16% 4% 2% 1.87 0.52 0.97
+QSAR_fish_toxicity -18% 0% -22% -7% -14% 1.60 0.61 0.97
 Table 14: Test results for MultiETs with complete data at 95% quantiles for seeds 0, 10, 20, 30, 40, with rows ordered by seed (ascending). Performance over the baseline is highlighted in bold.
-Interval Predictions Point Predictions
-Relative Absolute Relative Absolute
-Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP MPIW NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
-space_ga -8% 2% -11% 0.95 2.59 0.28 -13% -15% 0.55
-space_ga -3% 0% -3% 0.95 2.77 0.30 -14% -17% 0.55
-space_ga -2% 1% -4% 0.96 2.61 0.28 -13% -13% 0.57
-space_ga -9% 1% -11% 0.96 2.81 0.30 -15% -18% 0.54
-space_ga -7% 0% -8% 0.96 2.79 0.30 -19% -24% 0.51
-cpu_activity 11% -4% 13% 0.95 0.57 0.11 -13% -20% 0.98
-cpu_activity 14% -5% 16% 0.94 0.54 0.10 -13% -18% 0.98
-cpu_activity 9% -4% 11% 0.95 0.58 0.11 -13% -18% 0.98
-cpu_activity 10% -4% 13% 0.94 0.57 0.11 -15% -22% 0.97
-cpu_activity 10% -4% 13% 0.94 0.57 0.11 -16% -22% 0.97
-naval_propulsion_plant 131% 1% 56% 0.96 0.96 0.28 -320% -418% 0.96
-naval_propulsion_plant 144% 1% 59% 0.96 0.88 0.26 -286% -369% 0.97
-naval_propulsion_plant 174% 0% 63% 0.96 0.81 0.24 -257% -348% 0.97
-naval_propulsion_plant 127% 0% 56% 0.96 0.94 0.28 -324% -426% 0.96
-naval_propulsion_plant 108% 1% 52% 0.96 1.03 0.30 -391% -469% 0.95
-miami_housing -7% 0% -7% 0.95 1.22 0.15 -11% -20% 0.90
-miami_housing -10% -1% -10% 0.95 1.24 0.15 -6% -17% 0.90
-miami_housing -8% -1% -8% 0.95 1.25 0.15 -13% -23% 0.89
-miami_housing -9% -1% -10% 0.95 1.22 0.15 -9% -19% 0.90
-miami_housing -9% 0% -9% 0.94 1.24 0.15 -10% -22% 0.89
-kin8nm -10% 1% -12% 0.95 2.75 0.54 -36% -40% 0.46
-kin8nm -8% 0% -9% 0.94 2.66 0.52 -33% -36% 0.49
-kin8nm -11% 3% -15% 0.95 2.76 0.54 -32% -35% 0.50
-kin8nm -9% 2% -12% 0.94 2.67 0.52 -36% -40% 0.48
-kin8nm -10% 1% -13% 0.95 2.73 0.53 -33% -36% 0.49
-concrete_compressive_strength -18% -3% -18% 0.89 1.67 0.34 -78% -110% 0.74
-concrete_compressive_strength -5% -3% -3% 0.89 1.51 0.30 -71% -97% 0.76
-concrete_compressive_strength -6% -6% 0% 0.87 1.42 0.29 -63% -87% 0.77
-concrete_compressive_strength -6% -3% -4% 0.90 1.47 0.30 -63% -91% 0.77
-concrete_compressive_strength -6% 1% -7% 0.92 1.59 0.32 -61% -88% 0.77
-cars -35% 5% -60% 0.95 1.01 0.17 6% 3% 0.95
-cars -25% 1% -34% 0.91 0.83 0.14 8% 4% 0.95
-cars -19% 0% -24% 0.92 0.80 0.14 4% -2% 0.95
-cars -24% -3% -28% 0.90 0.86 0.15 2% 0% 0.95
-cars -21% -3% -22% 0.91 0.91 0.15 -3% 1% 0.94
-energy_efficiency 16% -8% 21% 0.90 0.18 0.05 0% 0% 1.00
-energy_efficiency 6% -2% 6% 0.97 0.21 0.06 4% 1% 1.00
-energy_efficiency 8% -4% 11% 0.95 0.21 0.06 5% 2% 1.00
-energy_efficiency 3% -3% 5% 0.96 0.21 0.06 1% 0% 1.00
-energy_efficiency 17% -4% 18% 0.95 0.19 0.06 5% 0% 1.00
-california_housing -2% -2% 1% 0.95 2.38 0.57 -16% -22% 0.70
-california_housing 3% -2% 5% 0.95 2.31 0.55 -13% -19% 0.72
-california_housing 2% -2% 4% 0.95 2.29 0.55 -15% -20% 0.71
-california_housing 0% -3% 3% 0.95 2.32 0.55 -16% -22% 0.71
-california_housing 1% -3% 4% 0.95 2.32 0.55 -14% -20% 0.72
-airfoil_self_noise -25% -5% -26% 0.94 2.17 0.46 -166% -184% 0.74
-airfoil_self_noise 0% -1% 1% 0.98 1.78 0.38 -59% -80% 0.89
-airfoil_self_noise -15% -1% -16% 0.98 1.98 0.42 -97% -106% 0.84
-airfoil_self_noise -28% -5% -32% 0.93 2.20 0.47 -169% -213% 0.67
-airfoil_self_noise -14% -3% -13% 0.96 1.93 0.41 -98% -125% 0.84
-QSAR_fish_toxicity -16% -2% -17% 0.88 2.45 0.38 -6% -10% 0.53
-QSAR_fish_toxicity -10% -2% -9% 0.89 2.43 0.38 -7% -12% 0.52
-QSAR_fish_toxicity 3% -6% 9% 0.88 2.38 0.37 -8% -12% 0.50
-QSAR_fish_toxicity -12% -4% -9% 0.88 2.41 0.38 -5% -10% 0.55
-QSAR_fish_toxicity 8% -9% 16% 0.84 2.03 0.32 -6% -13% 0.53
+Relative Metrics Absolute Metrics
+Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball CWR NMPIW PICP
+space_ga 5% -3% 8% 0% -8% 7.12 0.13 0.92
+space_ga -6% -2% -4% -9% -13% 3.01 0.32 0.95
+space_ga 3% 0% 3% 2% 3% 3.60 0.27 0.97
+space_ga 25% -2% 21% 13% 8% 4.46 0.21 0.93
+space_ga 0% 0% -1% 3% 8% 7.29 0.13 0.94
+cpu_activity -16% -1% -18% -14% -10% 8.51 0.11 0.95
+cpu_activity -9% 0% -11% -9% -11% 8.44 0.11 0.96
+cpu_activity -14% 0% -16% -11% -11% 8.43 0.11 0.96
+cpu_activity -16% 1% -20% -1% 17% 8.73 0.11 0.95
+cpu_activity -14% 0% -16% -8% 0% 8.75 0.11 0.95
+naval_propulsion_plant 84% -1% 46% 46% 33% 15.95 0.06 0.94
+naval_propulsion_plant 26% 0% 20% 27% 0% 12.02 0.08 0.95
+naval_propulsion_plant -2% -1% -1% -3% -25% 13.52 0.07 0.94
+naval_propulsion_plant 8% 1% 7% 14% -25% 11.94 0.08 0.96
+naval_propulsion_plant 1% 0% 1% 2% -20% 10.46 0.09 0.95
+miami_housing -19% -1% -23% -13% -35% 9.00 0.11 0.95
+miami_housing -35% 0% -56% -23% -38% 7.03 0.14 0.96
+miami_housing -13% 0% -15% -9% -33% 9.98 0.09 0.94
+miami_housing -20% 0% -25% -9% -11% 9.56 0.10 0.94
+miami_housing -17% 0% -21% -9% -12% 9.71 0.10 0.95
+kin8nm 17% 0% 14% 14% 15% 4.50 0.21 0.96
+kin8nm 4% 1% 3% 5% 6% 4.86 0.20 0.95
+kin8nm 18% -2% 16% 11% 11% 4.97 0.19 0.93
+kin8nm 2% 0% 2% 5% 6% 4.59 0.21 0.95
+kin8nm 3% 0% 3% 5% 0% 4.80 0.20 0.93
+concrete_compressive_strength 54% -4% 37% 26% 21% 3.33 0.29 0.96
+concrete_compressive_strength -5% -2% -2% -2% -19% 3.15 0.30 0.94
+concrete_compressive_strength 13% -3% 14% 8% -12% 3.51 0.26 0.92
+concrete_compressive_strength 73% -2% 43% 35% 31% 3.47 0.28 0.97
+concrete_compressive_strength 19% 2% 14% 17% 15% 3.67 0.26 0.95
+cars -11% -2% -11% -13% -29% 5.13 0.18 0.94
+cars -7% -1% -7% -9% -13% 5.21 0.19 0.97
+cars -18% 0% -22% -17% -19% 4.00 0.24 0.98
+cars 21% -5% 22% 5% -21% 7.75 0.12 0.92
+cars -44% 7% -89% -23% -14% 4.89 0.20 0.99
+energy_efficiency 102% 5% 48% 56% 50% 19.46 0.05 0.96
+energy_efficiency 74% -1% 42% 40% 40% 19.62 0.05 0.95
+energy_efficiency 65% -3% 41% 36% 25% 18.91 0.05 0.96
+energy_efficiency 46% 3% 30% 33% 0% 26.91 0.03 0.90
+energy_efficiency 149% -6% 63% 42% 0% 24.65 0.04 0.92
+california_housing -10% 0% -11% 0% -3% 1.99 0.48 0.96
+california_housing -14% 0% -16% -5% -13% 2.12 0.45 0.94
+california_housing -11% 0% -12% -1% -9% 2.12 0.45 0.95
+california_housing -7% -1% -6% 0% -9% 2.22 0.42 0.94
+california_housing -12% 0% -14% -2% -9% 2.20 0.43 0.95
+airfoil_self_noise 60% -1% 38% 33% 30% 5.45 0.18 0.97
+airfoil_self_noise 55% -1% 36% 37% 27% 4.83 0.20 0.97
+airfoil_self_noise 63% 0% 39% 43% 32% 4.98 0.20 0.98
+airfoil_self_noise 17% 1% 14% 20% 23% 4.84 0.20 0.95
+airfoil_self_noise 16% -2% 15% 17% 5% 4.48 0.21 0.95
+QSAR_fish_toxicity 21% -3% 19% 4% 6% 1.85 0.52 0.96
+QSAR_fish_toxicity 2% 1% 1% 1% 2% 1.57 0.62 0.98
+QSAR_fish_toxicity -2% 0% -2% -3% 2% 1.67 0.58 0.97
+QSAR_fish_toxicity 36% -3% 28% 13% 15% 2.04 0.47 0.96
+QSAR_fish_toxicity -15% 1% -19% -9% -11% 1.43 0.69 0.98
 Table 15: Test results for MultiMLPs with complete data at 95% quantiles for seeds 0, 10, 20, 30, 40, with rows ordered by seed (ascending). Performance over the baseline is highlighted in bold.
-Interval Predictions Point Predictions
-Relative Absolute Relative Absolute
-Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP MPIW NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
-space_ga 5% -2% 7% 0.95 2.11 0.23 0% -1% 0.74
-space_ga 3% -1% 3% 0.96 2.13 0.23 -1% -1% 0.75
-space_ga 6% -1% 7% 0.94 2.04 0.22 -3% -3% 0.74
-space_ga 5% -2% 6% 0.95 2.15 0.23 0% 0% 0.76
-space_ga 10% 0% 9% 0.95 2.06 0.22 1% 0% 0.76
-cpu_activity -18% 1% -22% 0.96 0.58 0.11 4% 2% 0.98
-cpu_activity -3% 0% -2% 0.95 0.51 0.10 8% 7% 0.98
-cpu_activity -9% 0% -10% 0.95 0.53 0.10 6% 5% 0.98
-cpu_activity -6% 0% -7% 0.95 0.54 0.10 6% 5% 0.98
-cpu_activity -2% 0% -2% 0.95 0.53 0.10 8% 5% 0.98
-naval_propulsion_plant 32% 1% 24% 0.95 0.23 0.07 -44% -48% 1.00
-naval_propulsion_plant 0% 1% -2% 0.96 0.31 0.09 -9% -11% 1.00
-naval_propulsion_plant 7% 0% 5% 0.95 0.24 0.07 -35% -13% 1.00
-naval_propulsion_plant -8% 1% -11% 0.96 0.28 0.08 -86% -63% 1.00
-naval_propulsion_plant -13% 2% -16% 0.96 0.34 0.10 -51% -45% 0.99
-miami_housing -38% 0% -60% 0.95 1.18 0.15 -6% 1% 0.91
-miami_housing -37% 0% -59% 0.95 1.16 0.14 -2% 7% 0.91
-miami_housing -43% 0% -74% 0.95 1.25 0.15 -4% 6% 0.91
-miami_housing -38% 1% -62% 0.95 1.15 0.14 -2% 3% 0.92
-miami_housing -33% -1% -47% 0.95 1.10 0.14 -1% 4% 0.91
-kin8nm 10% 1% 9% 0.95 1.03 0.20 4% 5% 0.93
-kin8nm 2% 1% 1% 0.96 1.07 0.21 10% 10% 0.94
-kin8nm 4% 1% 3% 0.95 1.06 0.21 10% 6% 0.94
-kin8nm 10% -3% 12% 0.93 0.99 0.19 6% 5% 0.93
-kin8nm 15% -1% 13% 0.94 1.03 0.20 7% 7% 0.93
-concrete_compressive_strength 7% -3% 9% 0.91 1.24 0.25 18% 19% 0.92
-concrete_compressive_strength 18% 1% 14% 0.97 1.39 0.28 9% 12% 0.91
-concrete_compressive_strength 2% 1% 1% 0.97 1.63 0.33 9% 15% 0.91
-concrete_compressive_strength 7% -5% 11% 0.94 1.62 0.33 4% 10% 0.90
-concrete_compressive_strength 19% -3% 18% 0.93 1.26 0.26 18% 20% 0.91
-cars 5% -9% 13% 0.88 0.66 0.11 -5% -2% 0.95
-cars 0% -3% 2% 0.93 0.80 0.14 -6% -2% 0.95
-cars -9% 0% -10% 0.93 0.80 0.14 1% 1% 0.96
-cars -8% -3% -4% 0.93 0.83 0.14 -2% -1% 0.95
-cars 25% -1% 21% 0.94 0.79 0.13 -4% 5% 0.95
-energy_efficiency 76% -3% 46% 0.94 0.16 0.04 31% 31% 1.00
-energy_efficiency 46% 7% 25% 0.99 0.19 0.06 35% 32% 1.00
-energy_efficiency 34% -1% 27% 0.95 0.17 0.05 28% 32% 1.00
-energy_efficiency 29% -3% 25% 0.95 0.16 0.05 31% 32% 1.00
-energy_efficiency 79% -3% 45% 0.96 0.17 0.05 37% 38% 1.00
-california_housing -13% 1% -16% 0.95 1.88 0.45 8% 11% 0.82
-california_housing -8% 0% -9% 0.95 1.78 0.42 7% 8% 0.81
-california_housing -16% 1% -20% 0.95 1.83 0.44 7% 9% 0.82
-california_housing -$-1$2% 0.94 1.82 0.43 7% 8% 0.81
-california_housing -12% 0% -13% 0.94 1.76 0.42 7% 9% 0.82
-airfoil_self_noise 98% -2% 51% 0.97 0.82 0.18 21% 19% 0.97
-airfoil_self_noise 33% 0% 24% 0.98 0.90 0.19 17% 12% 0.97
-airfoil_self_noise 36% -3% 28% 0.95 0.79 0.17 11% 9% 0.97
-airfoil_self_noise 83% -4% 47% 0.95 0.76 0.16 22% 22% 0.97
-airfoil_self_noise 93% -1% 49% 0.98 0.88 0.19 18% 18% 0.97
-QSAR_fish_toxicity -11% 2% -16% 0.91 2.53 0.39 5% 11% 0.51
-QSAR_fish_toxicity 3% 2% 1% 0.91 2.38 0.37 0% -1% 0.57
-QSAR_fish_toxicity 7% -1% 7% 0.91 2.31 0.36 6% 11% 0.54
-QSAR_fish_toxicity 18% 2% 14% 0.90 2.10 0.33 -4% 6% 0.52
-QSAR_fish_toxicity 9% -1% 9% 0.83 1.84 0.29 10% 9% 0.61
-Table 9: Test results for MultiXGBs with 50% missing data at 95% quantiles for seeds 0, 10, 20, 30, 40, with rows ordered by seed (ascending). Performance over the baseline is highlighted in bold.
-Interval Predictions Point Predictions
-Relative Absolute Relative Absolute
-Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP MPIW NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
-space_ga 4% -4% 7% 0.93 2.82 0.31 -11% -13% 0.35
-space_ga 2% -1% 3% 0.94 2.97 0.32 -6% -6% 0.45
-space_ga -4% -3% -2% 0.94 3.06 0.33 -12% -11% 0.44
-space_ga -1% -1% -1% 0.95 2.97 0.32 -1% -6% 0.46
-space_ga 2% 0% 2% 0.95 2.77 0.30 -5% -11% 0.54
-cpu_activity -9% 1% -11% 0.95 0.89 0.17 -7% -23% 0.94
-cpu_activity -26% 1% -44% 0.93 1.02 0.19 -1% -26% 0.81
-cpu_activity -15% -2% -14% 0.89 0.80 0.15 -30% -30% 0.73
-cpu_activity -1% -1% -3% 0.95 0.86 0.16 36% 5% 0.88
-cpu_activity -19% 2% -26% 0.98 1.12 0.21 -2% -11% 0.81
-naval_propulsion_plant -8% 0% -11% 0.98 2.93 0.86 -45% -90% 0.73
-naval_propulsion_plant 31% -2% 24% 0.93 1.50 0.44 5% -6% 0.84
-naval_propulsion_plant 18% -2% 17% 0.96 2.06 0.61 -30% -42% 0.83
-naval_propulsion_plant 53% -5% 37% 0.87 1.27 0.37 6% 6% 0.76
-naval_propulsion_plant -5% -3% -6% 0.93 2.22 0.65 -7% -11% 0.70
-miami_housing -10% -2% -11% 0.94 1.45 0.18 -7% -12% 0.80
-miami_housing -38% 0% -63% 0.93 1.86 0.23 -35% -35% 0.54
-miami_housing -26% -1% -33% 0.95 1.74 0.22 -10% -19% 0.76
-miami_housing 1% -3% 2% 0.95 1.35 0.17 -14% -20% 0.79
-miami_housing -17% 1% -22% 0.94 1.64 0.20 -4% -4% 0.72
-kin8nm -6% 0% -8% 0.97 3.36 0.66 0% 1% 0.40
-kin8nm -15% 1% -22% 0.97 3.46 0.68 -6% -8% 0.39
-kin8nm -8% 3% -12% 0.95 3.00 0.59 1% 2% 0.41
-kin8nm -9% 2% -14% 0.96 3.11 0.61 -5% -7% 0.44
-kin8nm -8% 1% -9% 0.95 3.23 0.63 -5% -6% 0.37
-concrete_compressive_strength -8% -3% -8% 0.95 2.81 0.57 -19% -22% 0.48
-concrete_compressive_strength 7% -5% 11% 0.93 2.79 0.56 -37% -44% 0.57
-concrete_compressive_strength 14% -3% 13% 0.97 2.69 0.55 -4% -9% 0.64
-concrete_compressive_strength 4% -1% 3% 0.95 2.83 0.57 -15% -24% 0.50
-concrete_compressive_strength -11% 0% -14% 0.97 3.01 0.61 -4% -3% 0.69
-cars 31% -3% 24% 0.93 1.91 0.32 -9% -16% 0.75
-cars 51% -7% 38% 0.93 1.84 0.31 -37% -25% 0.81
-cars -16% -11% -17% 0.89 2.17 0.37 -68% -80% 0.55
-cars -13% -3% -21% 0.94 2.18 0.37 -8% -14% 0.76
-cars -5% -12% 5% 0.83 2.00 0.34 -28% -39% 0.40
-energy_efficiency -11% 0% -13% 1.00 1.15 0.33 -15% -24% 0.97
-energy_efficiency 20% -10% 25% 0.89 0.84 0.24 -705% -520% 0.76
-energy_efficiency 42% -1% 23% 0.95 0.78 0.23 -37% -39% 0.94
-energy_efficiency 61% -4% 40% 0.94 0.70 0.20 -16% -44% 0.96
-energy_efficiency 20% -3% 16% 0.97 1.12 0.33 -155% -137% 0.93
-california_housing -14% 0% -17% 0.96 2.67 0.64 -3% -3% 0.65
-california_housing 0% -2% 0% 0.96 2.41 0.57 -11% -13% 0.73
-california_housing -7% -1% -8% 0.96 2.48 0.59 0% 1% 0.73
-california_housing 2% -3% 3% 0.92 2.28 0.54 -4% -7% 0.61
-california_housing -12% -1% -16% 0.97 2.63 0.63 -1% -3% 0.73
-airfoil_self_noise -30% -2% -43% 0.97 3.95 0.84 -94% -110% 0.50
-airfoil_self_noise -18% -1% -22% 0.91 2.98 0.64 -19% -29% 0.38
-airfoil_self_noise -23% -4% -25% 0.91 3.49 0.75 -43% -48% 0.19
-airfoil_self_noise -31% 1% -49% 0.97 3.61 0.77 -14% -8% 0.52
-airfoil_self_noise -17% -1% -21% 0.97 3.81 0.81 -14% -22% 0.43
-QSAR_fish_toxicity -8% 3% -14% 0.98 4.22 0.66 4% 4% 0.43
-QSAR_fish_toxicity -6% 1% -7% 0.97 3.93 0.61 -9% -3% 0.35
-QSAR_fish_toxicity 8% -5% 10% 0.88 2.59 0.40 -4% -5% 0.38
-QSAR_fish_toxicity -5% -5% -5% 0.88 2.83 0.44 -7% -3% 0.35
-QSAR_fish_toxicity 12% -11% 15% 0.85 2.58 0.40 -5% -3% 0.30
-Table 10: Test results for MultiETs with 50% missing data at 95% quantiles for seeds 0, 10, 20, 30, 40, with rows ordered by seed (ascending). Performance over the baseline is highlighted in bold.
-Interval Predictions Point Predictions
-Relative Absolute Relative Absolute
-Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP MPIW NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
-space_ga -7% -2% -6% 0.94 3.08 0.33 -18% -19% 0.29
-space_ga -12% 3% -18% 0.97 3.41 0.37 -7% -7% 0.40
-space_ga -13% 0% -16% 0.95 3.05 0.33 -13% -11% 0.42
-space_ga -13% 0% -18% 0.96 3.21 0.35 -8% -8% 0.41
-space_ga -10% 0% -11% 0.96 2.91 0.32 -15% -18% 0.46
-cpu_activity -18% -3% -23% 0.95 1.03 0.19 -40% -56% 0.92
-cpu_activity -12% -6% -11% 0.91 1.04 0.20 -15% -44% 0.81
-cpu_activity -5% -5% -2% 0.92 1.00 0.19 -43% -59% 0.75
-cpu_activity -13% -2% -14% 0.96 0.97 0.18 2% -19% 0.87
-cpu_activity -30% -1% -42% 0.96 1.31 0.25 -10% -43% 0.76
-naval_propulsion_plant -19% 0% -23% 0.98 3.17 0.94 -63% -158% 0.71
-naval_propulsion_plant 11% -5% 13% 0.94 2.24 0.66 -31% -78% 0.73
-naval_propulsion_plant 14% -2% 14% 0.97 2.26 0.67 -61% -147% 0.80
-naval_propulsion_plant 24% -9% 25% 0.89 1.99 0.59 -13% -42% 0.66
-naval_propulsion_plant -10% -2% -11% 0.95 2.75 0.81 -19% -52% 0.64
-miami_housing -15% -2% -17% 0.94 1.86 0.23 -11% -25% 0.79
-miami_housing -44% 5% -87% 0.96 2.39 0.30 -43% -64% 0.47
-miami_housing -41% 0% -69% 0.96 2.30 0.29 -27% -42% 0.72
-miami_housing -35% -2% -52% 0.96 2.18 0.27 -21% -33% 0.79
-miami_housing -31% 1% -48% 0.95 2.24 0.28 -22% -30% 0.64
-kin8nm -15% -1% -19% 0.96 3.55 0.69 -8% -8% 0.29
-kin8nm -14% 2% -19% 0.96 3.43 0.67 -10% -14% 0.32
-kin8nm -11% 0% -13% 0.94 3.20 0.63 -9% -12% 0.32
-kin8nm -14% 0% -18% 0.95 3.30 0.65 -14% -18% 0.37
-kin8nm -14% 2% -19% 0.96 3.42 0.67 -13% -16% 0.30
-concrete_compressive_strength -20% 4% -35% 0.96 3.01 0.61 -30% -43% 0.41
-concrete_compressive_strength -16% -8% -10% 0.92 3.08 0.62 -72% -94% 0.47
-concrete_compressive_strength -18% -3% -24% 0.96 2.99 0.60 -21% -33% 0.56
-concrete_compressive_strength -25% -4% -34% 0.94 2.86 0.58 -26% -36% 0.41
-concrete_compressive_strength -22% 2% -31% 0.96 3.05 0.62 -35% -50% 0.53
-cars -24% -6% -27% 0.92 1.99 0.34 -22% -23% 0.73
-cars -49% 4% -103% 0.97 2.70 0.46 -20% -17% 0.80
-cars -14% -5% -12% 0.93 2.02 0.34 -59% -56% 0.60
-cars -36% -6% -54% 0.93 2.44 0.41 -37% -37% 0.66
-cars -37% -6% -68% 0.89 2.12 0.36 -50% -25% 0.34
-energy_efficiency 2% -3% 5% 0.97 0.62 0.18 -18% -32% 0.98
-energy_efficiency -42% -6% -62% 0.94 0.85 0.25 -159% -104% 0.90
-energy_efficiency -4% -6% 2% 0.93 0.69 0.20 -8% -20% 0.95
-energy_efficiency -29% -4% -41% 0.94 0.79 0.23 -39% -82% 0.95
-energy_efficiency -75% 0% -316% 1.00 2.61 0.76 -122% -136% 0.92
-california_housing -2% -2% 0% 0.96 2.88 0.69 -11% -16% 0.60
-california_housing -9% -2% -8% 0.96 2.90 0.69 -19% -23% 0.64
-california_housing -5% -2% -4% 0.97 3.02 0.72 -11% -13% 0.64
-california_housing -1% -3% 1% 0.93 2.77 0.66 -15% -18% 0.50
-california_housing -10% -1% -10% 0.97 3.07 0.73 -13% -17% 0.64
-airfoil_self_noise -33% -3% -46% 0.97 3.62 0.77 -163% -218% 0.34
-airfoil_self_noise -36% -2% -56% 0.93 3.42 0.73 -27% -36% 0.42
-airfoil_self_noise -31% -5% -40% 0.92 3.42 0.73 -50% -83% 0.17
-airfoil_self_noise -32% -2% -45% 0.97 3.91 0.83 -45% -57% 0.43
-airfoil_self_noise -22% -2% -30% 0.94 3.29 0.70 -27% -51% 0.36
-QSAR_fish_toxicity 2% 1% 0% 0.98 4.12 0.64 4% 2% 0.51
-QSAR_fish_toxicity -10% -4% -7% 0.93 3.43 0.53 -12% -12% 0.40
-QSAR_fish_toxicity -7% 1% -11% 0.93 2.87 0.45 -6% -8% 0.45
-QSAR_fish_toxicity -17% -7% -16% 0.86 2.99 0.47 -23% -26% 0.25
-QSAR_fish_toxicity 0% -3% -1% 0.90 2.63 0.41 -4% -6% 0.34
-Table 11: Test results for MultiMLPs with 50% missing data at 95% quantiles for seeds 0, 10, 20, 30, 40, with rows ordered by seed (ascending). Performance over the baseline is highlighted in bold.
-Interval Predictions Point Predictions
-Relative Absolute Relative Absolute
-Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP MPIW NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
-space_ga -25% -3% -32% 0.95 3.45 0.37 -41% -40% 0.15
-space_ga -22% -3% -27% 0.96 3.58 0.39 -28% -25% 0.35
-space_ga -45% -2% -82% 0.95 4.52 0.49 -85% -61% -0.21
-space_ga -13% -2% -13% 0.95 3.12 0.34 -21% -20% 0.37
-space_ga -20% -6% -18% 0.92 3.07 0.33 -25% -26% 0.38
-cpu_activity -22% -2% -32% 0.95 1.14 0.21 14% 1% 0.92
-cpu_activity -44% 2% -80% 0.95 1.40 0.27 -30% -34% 0.72
-cpu_activity -49% 1% -100% 0.94 1.44 0.27 -37% -49% 0.54
-cpu_activity -45% 2% -89% 0.96 1.23 0.23 -20% -17% 0.78
-cpu_activity -38% -1% -61% 0.96 1.29 0.24 -3% -16% 0.78
-naval_propulsion_plant -36% -4% -52% 0.95 5.40 1.59 -215% -562% -0.85
-naval_propulsion_plant -44% -2% -78% 0.95 4.14 1.22 -99% -94% 0.16
-naval_propulsion_plant -64% -1% -175% 0.98 4.66 1.37 -141% -163% 0.50
-naval_propulsion_plant -35% 2% -67% 0.92 3.53 1.04 -24% -27% 0.27
-naval_propulsion_plant -38% -2% -70% 0.94 4.75 1.40 -62% -68% -0.21
-miami_housing -32% -1% -46% 0.94 1.48 0.18 -21% -20% 0.74
-miami_housing -46% 1% -88% 0.96 1.98 0.25 -41% -48% 0.53
-miami_housing -25% -2% -31% 0.95 1.69 0.21 -18% -19% 0.75
-miami_housing -34% -2% -49% 0.96 1.82 0.23 -23% -22% 0.77
-miami_housing -42% -2% -72% 0.91 1.91 0.24 -24% -32% 0.57
-kin8nm -14% 1% -17% 0.98 3.68 0.72 -3% 0% 0.43
-kin8nm -18% 0% -22% 0.97 3.67 0.72 -20% -22% 0.36
-kin8nm -17% 1% -22% 0.96 3.56 0.70 -12% -10% 0.39
-kin8nm -24% 0% -32% 0.96 3.46 0.68 -26% -24% 0.36
-kin8nm -20% 0% -25% 0.95 3.51 0.69 -20% -19% 0.31
-concrete_compressive_strength -10% 2% -13% 0.95 2.80 0.57 -9% -6% 0.51
-concrete_compressive_strength 9% -3% 8% 0.95 2.68 0.54 -15% -13% 0.59
-concrete_compressive_strength -1% -4% 3% 0.95 2.86 0.58 -20% -23% 0.52
-concrete_compressive_strength 16% -5% 16% 0.91 2.49 0.50 -10% -10% 0.46
-concrete_compressive_strength 13% -3% 11% 0.96 2.40 0.48 -5% -9% 0.60
-cars -23% -3% -33% 0.95 1.93 0.33 -3% -5% 0.75
-cars -28% -2% -43% 0.97 1.94 0.33 -36% -38% 0.75
-cars -27% -3% -39% 0.94 2.06 0.35 -34% -33% 0.70
-cars -41% 6% -85% 0.97 2.70 0.46 -32% -37% 0.63
-cars 8% -4% 10% 0.90 1.68 0.28 -32% -15% 0.48
-energy_efficiency -12% 0% -14% 0.99 1.44 0.42 -8% 1% 0.97
-energy_efficiency -4% -2% -2% 0.97 0.69 0.20 14% 18% 0.98
-energy_efficiency -24% -3% -37% 0.94 1.23 0.36 8% 2% 0.92
-energy_efficiency 10% -3% -1% 0.96 1.07 0.31 -8% 6% 0.96
-energy_efficiency -7% -10% 0% 0.89 0.84 0.24 -1% -32% 0.93
-california_housing -28% 1% -42% 0.95 3.00 0.71 -8% -9% 0.55
-california_housing -20% -2% -25% 0.97 2.85 0.68 -1% 0% 0.68
-california_housing -30% 0% -42% 0.96 2.90 0.69 -7% -3% 0.63
-california_housing -16% -1% -19% 0.93 2.70 0.64 -4% -3% 0.54
-california_housing -28% 0% -40% 0.96 2.83 0.67 -2% -1% 0.65
-airfoil_self_noise -14% 0% -16% 1.00 3.46 0.74 -20% -11% 0.84
-airfoil_self_noise 2% -6% 7% 0.89 2.60 0.56 -22% -13% 0.31
-airfoil_self_noise -12% -8% -9% 0.88 2.56 0.55 -23% -20% 0.44
-airfoil_self_noise -13% -5% -12% 0.93 2.94 0.63 -9% -1% 0.62
-airfoil_self_noise -5% 1% -8% 0.94 2.77 0.59 -3% 1% 0.54
-QSAR_fish_toxicity -16% 1% -20% 0.99 4.78 0.74 -8% -9% 0.44
-QSAR_fish_toxicity -7% -5% -7% 0.92 3.31 0.51 -21% -28% 0.28
-QSAR_fish_toxicity -5% -2% -8% 0.95 3.17 0.49 -6% -8% 0.39
-QSAR_fish_toxicity -8% -5% -8% 0.86 2.74 0.43 -21% -17% 0.28
-QSAR_fish_toxicity -17% 2% -28% 0.90 3.03 0.47 -10% -9% 0.32
-Appendix F Aggregated results for non-conformalized predictions
-Table 12: Test results for all models with complete-raw data at 95% quantiles aggregated over five seeds.
-For each metric, the mean and standard deviation of the performance across the seeds are separated by $\pm$.
-Performance over the baseline is highlighted in bold.
-Interval Predictions Point Predictions
-Relative Absolute Relative Absolute
-Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
-MultiXGBs
-space_ga -1%$\pm$56%$\pm$2-7%$\pm$60.89$\pm$0.01 0.20$\pm$0.01 -9%$\pm$1-10%$\pm$20.60$\pm$0.01
-cpu_activity 25%$\pm$69%$\pm$112%$\pm$40.89$\pm$0.01 0.07$\pm$0.00 21%$\pm$1-1%$\pm$20.98$\pm$0.00
-naval_propulsion_plant 156%$\pm$138%$\pm$158%$\pm$20.96$\pm$0.00 $0$14%$\pm$4-12%$\pm$20.99$\pm$0.00
-miami_housing 61%$\pm$5-7%$\pm$243%$\pm$30.82$\pm$0.01 0.06$\pm$0.00 4%$\pm$43%$\pm$30.91$\pm$0.01
-kin8nm 10%$\pm$35%$\pm$24%$\pm$40.90$\pm$0.01 0.38$\pm$0.01 -20%$\pm$1-22%$\pm$10.63$\pm$0.01
-concrete_compressive_strength 13%$\pm$426%$\pm$6-11%$\pm$70.91$\pm$0.01 $0$26%$\pm$6-40%$\pm$110.85$\pm$0.01
-cars 18%$\pm$815%$\pm$42%$\pm$80.89$\pm$0.02 0.12$\pm$0.01 -3%$\pm$3-1%$\pm$30.95$\pm$0.00
-energy_efficiency 247%$\pm$8922%$\pm$660%$\pm$170.88$\pm$0.02 0.05$\pm$0.01 -15%$\pm$22-16%$\pm$171.00$\pm$0.00
-california_housing 31%$\pm$50%$\pm$123%$\pm$40.88$\pm$0.01 $0$1%$\pm$1-1%$\pm$10.81$\pm$0.00
-airfoil_self_noise 61%$\pm$414%$\pm$931%$\pm$160.81$\pm$0.07 0.19$\pm$0.04 -64%$\pm$32-73%$\pm$410.86$\pm$0.05
-QSAR_fish_toxicity 10%$\pm$311%$\pm$3-1%$\pm$30.76$\pm$0.01 0.23$\pm$0.00 3%$\pm$31%$\pm$40.55$\pm$0.02
-MultiETs
-space_ga 9%$\pm$3-2%$\pm$19%$\pm$40.91$\pm$0.01 0.22$\pm$0.01 -16%$\pm$2-18%$\pm$30.54$\pm$0.02
-cpu_activity 27%$\pm$3-8%$\pm$027%$\pm$20.90$\pm$0.00 $0$14%$\pm$1-20%$\pm$20.98$\pm$0.00
-naval_propulsion_plant 219%$\pm$32-9%$\pm$071%$\pm$30.91$\pm$0.00 0.19$\pm$0.02 -320%$\pm$41-409%$\pm$360.96$\pm$0.01
-miami_housing 69%$\pm$3-13%$\pm$048%$\pm$10.86$\pm$0.00 $0$10%$\pm$2-20%$\pm$10.90$\pm$0.00
-kin8nm 10%$\pm$2-9%$\pm$017%$\pm$20.88$\pm$0.01 0.44$\pm$0.01 -35%$\pm$2-38%$\pm$20.48$\pm$0.01
-concrete_compressive_strength 9%$\pm$9-12%$\pm$219%$\pm$80.81$\pm$0.01 0.24$\pm$0.02 -67%$\pm$7-94%$\pm$90.76$\pm$0.01
-cars 11%$\pm$616%$\pm$2-5%$\pm$70.83$\pm$0.02 0.09$\pm$0.01 3%$\pm$31%$\pm$20.95$\pm$0.00
-energy_efficiency -4%$\pm$50%$\pm$7-5%$\pm$140.65$\pm$0.03 0.02$\pm$0.00 3%$\pm$30%$\pm$11.00$\pm$0.00
-california_housing 31%$\pm$2-7%$\pm$029%$\pm$10.90$\pm$0.00 0.41$\pm$0.01 -15%$\pm$1-21%$\pm$10.71$\pm$0.01
-airfoil_self_noise 18%$\pm$25-15%$\pm$425%$\pm$170.83$\pm$0.04 0.26$\pm$0.06 -117%$\pm$45-141%$\pm$510.80$\pm$0.08
-QSAR_fish_toxicity 12%$\pm$2-9%$\pm$119%$\pm$20.82$\pm$0.01 $0$6%$\pm$1-11%$\pm$10.53$\pm$0.01
-MultiMLPs
-space_ga 8%$\pm$10%$\pm$37%$\pm$30.81$\pm$0.01 0.14$\pm$0.00 0%$\pm$1-1%$\pm$10.75$\pm$0.01
-cpu_activity 8%$\pm$3-9%$\pm$415%$\pm$60.72$\pm$0.02 0.05$\pm$0.00 7%$\pm$25%$\pm$20.98$\pm$0.00
-naval_propulsion_plant 21%$\pm$180%$\pm$316%$\pm$130.92$\pm$0.01 0.07$\pm$0.01 -45%$\pm$25-36%$\pm$201.00$\pm$0.00
-miami_housing 9%$\pm$42%$\pm$36%$\pm$60.81$\pm$0.02 $0$3%$\pm$24%$\pm$20.91$\pm$0.00
-kin8nm 2%$\pm$78%$\pm$5-6%$\pm$100.81$\pm$0.02 0.13$\pm$0.01 7%$\pm$27%$\pm$20.93$\pm$0.00
-concrete_compressive_strength 98%$\pm$13-23%$\pm$761%$\pm$50.56$\pm$0.04 0.06$\pm$0.00 12%$\pm$515%$\pm$40.91$\pm$0.01
-cars 0%$\pm$146%$\pm$12-9%$\pm$300.76$\pm$0.08 0.07$\pm$0.01 -3%$\pm$30%$\pm$30.95$\pm$0.00
-energy_efficiency 82%$\pm$1825%$\pm$2331%$\pm$110.62$\pm$0.04 0.02$\pm$0.00 32%$\pm$333%$\pm$31.00$\pm$0.00
-california_housing -14%$\pm$28%$\pm$1-26%$\pm$30.89$\pm$0.01 0.31$\pm$0.01 7%$\pm$09%$\pm$10.82$\pm$0.00
-airfoil_self_noise 74%$\pm$410%$\pm$639%$\pm$150.75$\pm$0.02 0.08$\pm$0.00 18%$\pm$416%$\pm$50.97$\pm$0.00
-QSAR_fish_toxicity -6%$\pm$411%$\pm$3-19%$\pm$50.76$\pm$0.03 0.23$\pm$0.01 4%$\pm$57%$\pm$40.55$\pm$0.04
-Table 13: Test results for all models with missing-raw data at 95% quantiles aggregated over five seeds.
-For each metric, the mean and standard deviation of the performance across the seeds are separated by $\pm$.
-Performance over the baseline is highlighted in bold.
-Interval Predictions Point Predictions
-Relative Absolute Relative Absolute
-Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
-MultiXGBs
-space_ga -5%$\pm$64%$\pm$2-12%$\pm$70.86$\pm$0.02 0.22$\pm$0.01 -7%$\pm$4-9%$\pm$30.45$\pm$0.06
-cpu_activity -5%$\pm$1912%$\pm$8-37%$\pm$460.89$\pm$0.06 0.14$\pm$0.04 -1%$\pm$21-17%$\pm$130.83$\pm$0.07
-naval_propulsion_plant 105%$\pm$29-16%$\pm$656%$\pm$60.70$\pm$0.06 0.18$\pm$0.02 -14%$\pm$20-29%$\pm$350.77$\pm$0.05
-miami_housing 16%$\pm$13-5%$\pm$510%$\pm$160.82$\pm$0.06 0.10$\pm$0.02 -14%$\pm$11-18%$\pm$100.72$\pm$0.10
-kin8nm -5%$\pm$23%$\pm$2-14%$\pm$50.88$\pm$0.01 0.48$\pm$0.01 -3%$\pm$3-4%$\pm$40.40$\pm$0.02
-concrete_compressive_strength 2%$\pm$38%$\pm$4-10%$\pm$30.77$\pm$0.03 0.30$\pm$0.01 -16%$\pm$12-20%$\pm$140.58$\pm$0.08
-cars -15%$\pm$1426%$\pm$9-63%$\pm$260.85$\pm$0.04 0.27$\pm$0.03 -30%$\pm$22-35%$\pm$240.66$\pm$0.15
-energy_efficiency 10%$\pm$2611%$\pm$6-24%$\pm$430.92$\pm$0.02 0.22$\pm$0.05 -186%$\pm$265-153%$\pm$1880.91$\pm$0.08
-california_housing 26%$\pm$7-2%$\pm$321%$\pm$60.88$\pm$0.03 0.40$\pm$0.03 -4%$\pm$4-5%$\pm$50.69$\pm$0.05
-airfoil_self_noise -2%$\pm$18-3%$\pm$4-4%$\pm$170.73$\pm$0.04 0.34$\pm$0.06 -37%$\pm$31-43%$\pm$360.41$\pm$0.12
-QSAR_fish_toxicity -2%$\pm$30%$\pm$5-6%$\pm$70.73$\pm$0.03 0.26$\pm$0.02 -4%$\pm$4-2%$\pm$30.36$\pm$0.04
-MultiETs
-space_ga 4%$\pm$3-5%$\pm$17%$\pm$40.88$\pm$0.02 0.24$\pm$0.01 -12%$\pm$4-14%$\pm$40.40$\pm$0.05
-cpu_activity -16%$\pm$10-5%$\pm$1-20%$\pm$180.93$\pm$0.01 0.20$\pm$0.03 -30%$\pm$23-54%$\pm$200.79$\pm$0.10
-naval_propulsion_plant 58%$\pm$5-17%$\pm$446%$\pm$30.82$\pm$0.05 0.41$\pm$0.03 -36%$\pm$21-93%$\pm$480.71$\pm$0.05
-miami_housing 20%$\pm$14-10%$\pm$421%$\pm$130.87$\pm$0.05 0.14$\pm$0.03 -27%$\pm$11-37%$\pm$110.67$\pm$0.11
-kin8nm 5%$\pm$2-8%$\pm$212%$\pm$40.87$\pm$0.02 0.50$\pm$0.02 -11%$\pm$2-13%$\pm$30.32$\pm$0.03
-concrete_compressive_strength 1%$\pm$5-14%$\pm$313%$\pm$20.80$\pm$0.02 0.37$\pm$0.01 -35%$\pm$21-50%$\pm$250.49$\pm$0.06
-cars -23%$\pm$6-8%$\pm$5-35%$\pm$140.83$\pm$0.04 0.24$\pm$0.02 -33%$\pm$18-31%$\pm$180.66$\pm$0.09
-energy_efficiency -15%$\pm$160%$\pm$9-36%$\pm$420.85$\pm$0.05 0.15$\pm$0.01 -73%$\pm$74-84%$\pm$640.94$\pm$0.04
-california_housing 21%$\pm$7-7%$\pm$222%$\pm$50.91$\pm$0.02 0.52$\pm$0.03 -14%$\pm$3-17%$\pm$30.61$\pm$0.05
-airfoil_self_noise -7%$\pm$18-20%$\pm$28%$\pm$160.76$\pm$0.03 0.42$\pm$0.08 -65%$\pm$54-93%$\pm$700.33$\pm$0.11
-QSAR_fish_toxicity 10%$\pm$7-9%$\pm$316%$\pm$30.82$\pm$0.03 0.33$\pm$0.03 -8%$\pm$8-10%$\pm$90.39$\pm$0.09
-MultiMLPs
-space_ga -11%$\pm$5-15%$\pm$21%$\pm$60.68$\pm$0.03 0.16$\pm$0.01 -40%$\pm$23-34%$\pm$150.21$\pm$0.23
-cpu_activity -50%$\pm$1622%$\pm$14-193%$\pm$1180.86$\pm$0.10 0.18$\pm$0.07 -15%$\pm$19-23%$\pm$170.75$\pm$0.12
-naval_propulsion_plant -52%$\pm$19-14%$\pm$24-102%$\pm$560.63$\pm$0.18 0.34$\pm$0.07 -108%$\pm$66-183%$\pm$195-0.03$\pm$0.47
-miami_housing -33%$\pm$1513%$\pm$10-92%$\pm$700.85$\pm$0.06 0.12$\pm$0.04 -25%$\pm$8-28%$\pm$110.67$\pm$0.10
-kin8nm -11%$\pm$5-5%$\pm$2-8%$\pm$60.77$\pm$0.01 0.35$\pm$0.02 -16%$\pm$8-15%$\pm$90.37$\pm$0.04
-concrete_compressive_strength -15%$\pm$13-13%$\pm$28-10%$\pm$480.60$\pm$0.22 0.22$\pm$0.11 -12%$\pm$5-12%$\pm$60.54$\pm$0.05
-cars -51%$\pm$443%$\pm$7-211%$\pm$340.93$\pm$0.02 0.29$\pm$0.03 -28%$\pm$12-26%$\pm$130.66$\pm$0.10
-energy_efficiency -13%$\pm$189%$\pm$15-30%$\pm$250.66$\pm$0.13 0.09$\pm$0.04 1%$\pm$9-1%$\pm$160.95$\pm$0.03
-california_housing -17%$\pm$64%$\pm$3-28%$\pm$140.88$\pm$0.03 0.44$\pm$0.04 -5%$\pm$2-3%$\pm$30.61$\pm$0.06
-airfoil_self_noise 2%$\pm$16-2%$\pm$5-1%$\pm$210.75$\pm$0.07 0.25$\pm$0.08 -15%$\pm$8-9%$\pm$80.55$\pm$0.18
-QSAR_fish_toxicity -17%$\pm$56%$\pm$6-30%$\pm$110.76$\pm$0.04 0.28$\pm$0.03 -13%$\pm$7-14%$\pm$80.34$\pm$0.06
-Appendix G Example of learning with non-normality
-To illustrate how SEMF adapts to non-normal outcomes, we provide an example from the naval_propulsion_plant dataset [57].
-Figure 2 shows the distribution of the ground-truth $y$ variable which in this case is gt_compressor_decay_state_coefficient.
-The values are uniformly distributed, and we only standardize the values without changing the shape of the distribution.
-Figure 2: Distribution of the standardized outcome ($y$) variable for the naval_propulsion_plant dataset which shows that $y$ is uniformly distributed prior to any training.
-After training our SEMF model under the normality assumption with the ideal hyper-parameters (and a seed of 0), sampling from a normal distribution for the $z$ dimension, we infer on some randomly sampled test instances that provide us with the prediction intervals in Figure 3.
-The ‘SEMF intervals’ can be compared with XGBoost quantile regression, constituting our baseline.
-This figure shows that SEMF’s predicted intervals are better than the baseline.
-This plot alone does not tell us much about the predicted output distribution.
-Therefore, we provide Figure 4.
-The last plot shows that for a handful of the instances, the predicted values can take any shape and are not necessarily normal.
-Figure 3: Predicted intervals and true values on 100 randomly selected test samples for SEMF MultiXGBs and XGBoost quantile regression with complete data at 95% quantiles (according to Eq 13).
-Figure 4: Violin plot for the first ten instances of the test set where both the ground truth and 50 values inferred by SEMF (according to Eq 12) are generated for each instance.
-We have added jitter here, so the points do not perfectly align along the x-axis.